### v1

# Compression fronts from fast radio bursts

Abstract
When a fast radio burst (FRB) expands from its source through a surrounding tenuous plasma, it strongly heats and compresses the plasma at radii up to $\sim 10^{14}$ cm. The likely central engines of FRBs are magnetars, and their ambient plasma at radii $r\gg 10^{10}$ cm is a magnetized $e^{\pm}$ wind. We formulate equations describing the FRB-plasma interaction, solve them numerically, and provide an approximate analytical picture of the interaction. We find the following: (1) FRBs emitted at $r<r_{\rm stoch}\sim 10^{12}$ cm induce fast stochastic heating and strong compression of the wind, sweeping it like a broom. The outcome of this interaction is determined by the energy losses of the radio wave. We evaluate the parameter space where the FRB survives its interaction with the wind. (2) At radii $r>r_{\rm stoch}$, the FRB induces regular heating to the Lorentz factor $\sim a_{0}$, where $a_{0}\propto r^{-1}$ is the wave strength parameter. At $r>r_{\star}\sim 10^{13}\,$cm, the FRB drives a quasisteady compression wave in the wind, with compression factor $C_{\star}\approx 1+a_{0}^{2}$. Both characteristic radii $r_{\rm stoch}$ and $r_{\star}$ scale with FRB luminosity $L$ as $L^{1/3}$. FRBs avoid damping if they are released into the wind medium outside $r_{\rm damp}\sim 10^{11}$ cm.
Magnetars (992); Plasma astrophysics (1261); Radio bursts (1339); X-ray transient sources (1852); Magnetohydrodynamics (1964); Radiative processes (2055)
1 Introduction
Observations of fast radio bursts (FRBs) demonstrate the existence of ultrastrong electromagnetic waves in the Universe. They have characteristic frequencies $\nu=\omega/2\pi\sim 1$ GHz and durations $T\sim 1$ ms (Petroff et al., 2019), and their dimensionless strength parameter is [^1] Throughout the paper we use notation $Q_{n}$ for a quantity $Q$ normalized to $10^{n}$ in CGS units, e.g. $\nu_{9}=\nu/10^{9}{\rm Hz}$.
$a\equiv\frac{e\langle E_{\rm wave}^{2}\rangle^{1/2}}{mc\,\omega}\approx 16\,\nu_{9}^{-1}L_{42}^{1/2}r_{12}^{-1},$ (1)
where $L=cr^{2}\langle E_{\rm wave}^{2}\rangle$ is the isotropic equivalent of wave luminosity; $\langle E_{\rm wave}^{2}\rangle^{1/2}$ is the root-mean-square of the wave electric field when the wave expands to a radius $r$.
FRBs have to propagate through the ambient plasma surrounding the source. The likely central engines of FRBs are neutron stars, which are surrounded by electron-positron plasma. The plasma is strongly magnetized:
$\sigma\equiv\frac{B^{2}}{4\pi\rho c^{2}}\gg 1,$ (2)
where the magnetic field $B$ and density $\rho$ are measured in the plasma rest frame. Near the neutron star, the plasma is locked in the closed magnetosphere, rotating with a period $P\sim 1-10$ s. Outside the light cylinder $R_{\rm LC}=cP/2\pi\approx 5\times 10^{9}(P/1{\rm\,s})$ cm, a relativistic wind flows from the magnetosphere, powered by the star’s rotation. Both the magnetosphere and the wind have $\sigma\gg 1$ (see Cerutti & Beloborodov (2017) for a review).
The plasma around neutron stars is tenuous, however it can have a strong effect on FRB propagation. The radio wave with $a\gg 1$ accelerates plasma particles to high energies, and this process can damp the wave, in particular if the energized plasma experiences radiative losses. Strong radiative damping has been demonstrated for FRBs emitted near a magnetar and propagating through its equilibrium outer magnetosphere at radii $r\sim 10^{8}-10^{9}$ cm (Beloborodov, 2021, 2024).
[…]
The present paper revisits the FRB-wind interaction. The problem is formulated in section 2 and approached in two steps:
First, we use test-particle simulations to demonstrate two types of heating, regular and stochastic, of a plasma exposed to (and compressed by) an FRB (section 3). When the plasma magnetic field is increased by compression, regular oscillations of particles in the radio wave are changed to stochastic resonant heating. This change boosts the heating and makes it irreversible.
Then, we examine how heating by the FRB results in the bulk acceleration and compression of the plasma. Heating determines acceleration because energy $\Delta{\cal E}$ received from the radio wave always comes with momentum $\Delta{\cal E}/c$ (the wave loses momentum $\Delta{\cal E}/c$). Plasma is accelerated and compressed on timescales much longer than the Larmor period and the wave oscillation period. [^2] In simplest cases, the effect may be thought of as a ponderomotive force exerted by the wave. In the presence of strong radiative losses it is more appropriate to view it as a scattering process (B22). In any case, it is described from first principles by solving for the particle motion in the wave and finding $\Delta{\cal E}/c$. Therefore, it behaves as a fluid governed by magnetohydrodynamics (MHD). In section 4, we formulate MHD equations for the plasma bulk motion inside a propagating radio wave packet. These equations determine plasma compression by FRBs.
Before considering FRBs expanding through a radial wind, we examine the compression front in a plane geometry with a uniform background. First, we show that the problem has a simple steady-state solution (section 5). Its compression factor $C_{\star}$ is directly related to the particle energy gained from the radio wave and becomes huge for waves with strength parameter $a\gg 1$. Remarkably, $C_{\star}$ is independent of the plasma magnetization $\sigma$. This fact casts some doubt on the relevance of the steady-state solution; in particular, it clearly becomes irrelevant in the vacuum limit ($\sigma\rightarrow\infty$) where the radio wave should have no compression effect on a background magnetic field. A complete picture is revealed only by the full time-dependent solution (section 6). It shows that the MHD flow pattern in the propagating wave packet is gradually attracted to the steady state, and this relaxation takes a long time $t_{\star}$ proportional to $\sigma$. For propagation times $t<t_{\star}$, the steady state with large $C_{\star}$ is not reached; we evaluate the actual compression factor $C\ll C_{\star}$ in this situation.
These results allow one to investigate the FRB interaction with a magnetar wind in a broad range of radii. We describe the wind in section 7, and then investigate the compression front induced by the FRB in the wind. We identify two zones, with regular and stochastic heating, and find the transition radius $r_{\rm stoch}$ between them. Compression fronts from FRBs emitted in each zone are described in sections 8 and 9. FRB damping by its interaction with the wind is investigated in section 10. Our conclusions are summarized in section 11.
2 Problem formulation
2.1 The role of magnetic field
For an unmagnetized plasma, particle acceleration in a strong electromagnetic wave is a textbook problem (Landau & Lifshitz, 1975) well known in the context of laser experiments (Bulanov et al., 2015). It was also discussed in the context of early pulsar models (Gunn & Ostriker, 1971). The wave accelerates initially static particles to Lorentz factors $\gamma=1+a^{2}/2$, and their average velocity is aligned with the wave propagation direction. This average motion gives the plasma the bulk Lorentz factor $\Gamma\approx a\gg 1$.
The problem qualitatively changes when the background plasma is magnetized (B22). Now, in addition to the oscillating electromagnetic field of the wave, there is a background magnetic field $\boldsymbol{B}_{\rm bg}$, which arrests the particle motion across $\boldsymbol{B}_{\rm bg}$ and changes the frame where the particles have zero average momentum (the “fluid frame”). In the media surrounding FRB sources, $B_{\rm bg}$ is well below the radio wave amplitude $E_{0}$ at all radii $r>R_{\rm LC}$ considered in this paper. At the same time, $B_{\rm bg}$ far exceeds a minimum value sufficient to “magnetize” the plasma: the particle Larmor period is much smaller than the duration of the wave packet $T\sim 1$ ms. Besides the wave strength parameter $a$, the problem of wave-plasma interaction has two more dimensionless parameters: the plasma magnetization $\sigma$ and the ratio of gyrofrequency $\omega_{B}=eB_{\rm bg}/mc$ to the wave frequency $\omega$.
The ratio $\omega_{B}/\omega$ plays a particularly important role. B22 showed that waves with $\omega>\omega_{B}$ drive a regular particle motion in the fluid frame: fast oscillations with wave frequency $\omega$ and Lorentz factor $\gamma\approx a$, superposed with (subdominant) slow gyration in the background magnetic field. If $1<\omega_{B}/\omega<a$, gyration is accompanied by a resonant wave-particle interaction each Larmor period, leading to quick stochastic particle acceleration to $\gamma\gg a$. Oscillations in the wave then become energetically subdominant compared to the Larmor rotation with the stochastically pumped $\gamma$. The stochastic heating gives a well-defined mean expectation for the particle energy, which has been found analytically and confirmed numerically.
In addition to heating the plasma, the radio wave also deposits momentum, pushing the plasma along the wave direction. The effective inertia of a plasma with $\sigma\gg 1$ is increased by its magnetic field, which reduces its bulk acceleration. The bulk acceleration leads to compression of the plasma and its magnetic field (Beloborodov, 2021).
Dynamics of a plasma exposed to the radio wave is easiest to investigate in the plane geometry, assuming a uniform magnetized background ahead of the wave. Details of the radio wave are not important, however one needs to choose a concrete shape of the wave packet when simulating the plasma dynamics numerically.
2.2 Wave packet
As an example, consider a nearly harmonic plane electromagnetic wave with amplitude $E_{0}$, frequency $\omega$, and duration $T\gg 2\pi/\omega$. Its strength parameter (defined in Equation (1)) is
$a=\frac{a_{0}}{\sqrt{2}},\qquad a_{0}=\frac{eE_{0}}{mc\omega}.$ (3)
The wave speed is weakly affected by the presence of a tenuous plasma, so it propagates with speed of light $c$. It will be convenient to use the coordinate
$\xi\equiv t-\frac{z}{c},$ (4)
where the $z$-axis is chosen along the wave propagation direction. The radio wave packet is static in the $\xi$-coordinate and occupies the region $0<\xi<T$. The wave amplitude is zero at the leading edge of the packet $\xi=0$, rises on a timescale $\xi_{\rm rise}<T$, and then drops back to zero at $\xi=T$. In numerical examples we will use packets with the envelope
$a(\xi)=a_{\rm max}\sin^{2}(\pi\xi/T),\qquad 0<\xi<T.$ (5)
The cold unperturbed medium ahead of the packet (at $\xi<0$) will be called “upstream.” The upstream medium is described by a uniform mass density $\rho_{\rm u}$ and a uniform magnetic field $\boldsymbol{B}_{\rm u}$. We will focus on the problem with $\boldsymbol{B}_{\rm u}$ oriented perpendicular to the wave propagation direction (as expected in winds, which carry magnetic fields transverse to the radial direction).
Two possible polarizations of the radio wave can be considered: the X-mode ($\boldsymbol{E}_{\rm wave}\perp\boldsymbol{B}_{\rm u}$) and the O-mode ($\boldsymbol{E}_{\rm wave}\parallel\boldsymbol{B}_{\rm u}$). We will use below the results of B22 for test-particle motion in the X-mode waves; similar results within numerical coefficients ${\cal O}(1)$ were recently published for the O-mode (Huang et al., 2024). The X-mode has the oscillating magnetic field $\boldsymbol{B}_{\rm wave}$ parallel to the background field $\boldsymbol{B}_{\rm u}$.
2.3 Plasma drift through the wave packet
The plasma behaves as an MHD fluid on timescales longer than the Larmor period. The fluid motion in the plane geometry with $\boldsymbol{B}_{\rm u}$ perpendicular to the $z$-axis is fully described by the drift velocity parallel to the $z$-axis,
$\boldsymbol{\beta}_{\rm D}=\frac{\boldsymbol{v}_{\rm D}}{c}=\frac{\overline{\boldsymbol{E}}\times\overline{\boldsymbol{B}}}{{\overline{B}}^{2}}.$ (6)
Here $\overline{\boldsymbol{E}}$ and $\overline{\boldsymbol{B}}$ are the MHD fields; the bar signifies averaging over a timescale longer than the Larmor time (and longer than the wave oscillation period). The average fields can also be denoted as $\boldsymbol{E}_{\rm bg}$ and $\boldsymbol{B}_{\rm bg}$ because they serve as the effective background fields in the problem of particle motion in the wave, as described below.
A natural choice for a reference frame in the wave-plasma interaction problem is such that the upstream plasma is at rest. When discussing wave propagation in a wind, we will also view the problem in the lab frame where the upstream moves with a large speed $\beta_{\rm u}>0$ along the wave propagation direction. In addition, frames with opposite motion, where the upstream plasma has $\beta_{\rm u}<0$, may be useful in numerical simulations. In all these frames, $\boldsymbol{v}_{\rm D}$ is parallel to the $z$-axis, both outside and inside the wave packet. The fluid four-velocity $u_{\rm D}^{\mu}$ then has two non-zero components:
$u_{\rm D}^{t}=\gamma_{\rm D}=(1-\beta_{\rm D}^{2})^{-1/2},\qquad u_{\rm D}^{z}=\gamma_{\rm D}\beta_{\rm D}.$ (7)
Dynamical equations for $u_{\rm D}^{\mu}$ and final results will be formulated in a Lorentz-invariant way, so they will hold for any frame boosted along $z$ or, equivalently, any $\beta_{\rm u}$.
Plasma with any speed $v_{\rm D}=dz/dt$ changes its $\xi$-coordinate with rate
$\frac{d\xi}{dt}=1-\beta_{\rm D}>0.$ (8)
Thus, the MHD fluid always moves toward increasing $\xi$ through the wave packet. When viewed in the rest frame of the upstream medium, the initially static fluid enters the wave packet at $\xi=0$ and develops speed $\beta_{\rm D}\neq 0$ at $\xi>0$. We wish to find the fluid motion in the packet, $\beta_{\rm D}(\xi)$, and the fluid compression factor $C(\xi)=\rho(\xi)/\rho_{\rm u}$.
3 Plasma heating by the wave
[…]
(1) Waves with frequency $\tilde{\omega}$ exceeding the gyrofrequency $\tilde{\omega}_{B}=e\tilde{B}_{\rm bg}/mc$ drive regular particle oscillations with
$\tilde{\gamma}=\sqrt{1+a^{2}}\approx a\gg 1\qquad(\tilde{\omega}_{B}<\tilde{\omega}).$ (9)
Thus, the local $\tilde{\gamma}$ in the packet simply tracks the local strength parameter of the wave $a$.
(2) In the other heating regime, $1<\tilde{\omega}_{B}/\tilde{\omega}<a$, the resonant pumping mechanism is activated, and the wave drives stochastic particle acceleration. Then, the average $\tilde{\gamma}$ inside a fluid element quickly exceeds $a$ and keeps growing as the fluid moves through the wave packet (until radiative losses limit the growth or fluid exits behind the packet). The mean expectation for the particle Lorentz factor $\tilde{\gamma}$ grows according to the equation (B22,Beloborodov 2024)
$\frac{d\tilde{\gamma}}{d\tilde{t}}=\chi\frac{a_{0}^{2}\tilde{\omega}_{B}^{1/3}\tilde{\omega}^{2/3}}{\tilde{\gamma}^{4/3}}\qquad(1<\tilde{\omega}_{B}/\tilde{\omega}<a_{0}),$ (10)
where $a_{0}^{2}=2a^{2}$, $\chi\sim 1$ is a numerical coefficient, and all quantities are measured in the fluid rest frame $\tilde{\cal K}$.
Calculations in B22 were performed assuming that the background magnetic field $\boldsymbol{B}_{\rm bg}$ is static, i.e. there is no background electric field $\boldsymbol{E}_{\rm bg}$ and hence no average particle drift: $\boldsymbol{v}_{\rm D}/c=\boldsymbol{E}_{\rm bg}\times\boldsymbol{B}_{\rm bg}/B_{\rm bg}^{2}=0$. This corresponds to using the frame $\tilde{\cal K}$ where the magnetized fluid is at rest — the energized particles have no net momentum as they gyrate in a static magnetic field. Note that the wave fields $\boldsymbol{E}_{\rm wave}$ and $\boldsymbol{B}_{\rm wave}$ oscillate around zero and do not contribute to the MHD fields $\overline{\boldsymbol{E}}$ and $\overline{\boldsymbol{B}}$ (averaged over a timescale longer than the particle Larmor period and the wave period). So, the MHD fields serve as the background fields in the test-particle calculations:
$\overline{\boldsymbol{E}}=\overline{\boldsymbol{E}_{\rm bg}+\boldsymbol{E}_{\rm wave}}=\boldsymbol{E}_{\rm bg},\quad\overline{\boldsymbol{B}}=\overline{\boldsymbol{B}_{\rm bg}+\boldsymbol{B}_{\rm wave}}=\boldsymbol{B}_{\rm bg}.$ (11)
In general, $\overline{\boldsymbol{E}}\neq 0$ which corresponds to a moving MHD fluid $v_{\rm D}\neq 0$. If $v_{\rm D}$ varies (the plasma is accelerated in the wave packet), no global inertial frame exists where the fluid would be static, and one has to consider the heating effect of the wave in a moving MHD fluid. Then, the test particle motion should be examined with $E_{\rm bg}\neq 0$.
Remarkably, the results found in B22 at $E_{\rm bg}=0$ also hold in the accelerating background, as long as the particle motion is measured in the local drift frame $\tilde{\cal K}$ where $\tilde{\boldsymbol{E}}_{\rm bg}=0$. This fact is illustrated below by direct calculations of particle motion in a wave packet with a non-static background field. We will prescribe smoothly varying fields $\boldsymbol{B}_{\rm bg}(\xi)$ and $\boldsymbol{E}_{\rm bg}(\xi)$ (varying on a scale much greater than the particle Larmor radius) and repeat the calculation of B22 for test particles.
In the simulations, we will use the frame where the upstream plasma is at rest, $\beta_{\rm D}=\beta_{\rm u}=0$, and prescribe the background field of the form
$B_{\rm bg}=\frac{\kappa^{2}+1}{2}B_{\rm bg}^{\rm u},\qquad E_{\rm bg}=\frac{\kappa^{2}-1}{2}B_{\rm bg}^{\rm u},$ (12)
$\kappa\equiv\gamma_{\rm D}(1+\beta_{\rm D})=\frac{1}{\gamma_{\rm D}(1-\beta_{\rm D})}.$ (13)
The upstream has $\kappa=\kappa_{\rm u}=1$, $E_{\rm bg}=0$, and $B_{\rm bg}=B_{\rm bg}^{\rm u}$. Fluid motion inside the wave packet is specified by a chosen function $\kappa(\xi)$. For the present purposes, Equation (12) may be viewed as an arbitrary prescription to test heating of fluid with a non-uniform motion $\beta_{\rm D}(\xi)$. As explained in next sections, any steady pattern of fluid motion in the wave packet satisfies Equation (12). It also satisfies $\tilde{\rho}/\kappa=const$ (uniform across the packet), where $\tilde{\rho}=\rho/\gamma_{\rm D}$ is the fluid proper density. Thus, the adopted prescription corresponds to a steady flow pattern with the proper compression factor $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}=\kappa$. The magnetic field $\tilde{B}_{\rm bg}$ is frozen in the fluid and compressed by the same factor $\kappa$ while $B_{\rm bg}=\gamma_{\rm D}\tilde{B}_{\rm bg}$ obeys Equation (12).
Our simulation will track a test particle initially located in the static upstream region and exposed to the propagating wave packet. The particle motion in the wave packet is found by solving the dynamical equation,
$mc\,\frac{d\boldsymbol{u}}{dt}=e\left(\boldsymbol{E}+\boldsymbol{\beta}\times\boldsymbol{B}\right),$ (14)
where vector $\boldsymbol{u}=\gamma\boldsymbol{\beta}$ is the spatial part of the particle four-velocity; radiative losses are neglected here. The fields $\boldsymbol{E}(\xi)$ and $\boldsymbol{B}(\xi)$ include the background and wave fields: $\boldsymbol{E}=\boldsymbol{E}_{\rm bg}+\boldsymbol{E}_{\rm wave}$ and $\boldsymbol{B}=\boldsymbol{B}_{\rm bg}+\boldsymbol{B}_{\rm wave}$. [^4] Notation here differs from B22 where $\boldsymbol{E}$ and $\boldsymbol{B}$ denoted the wave fields and the dynamical equation assumed $\boldsymbol{E}_{\rm bg}=0$.
3.1 Regular heating ($\tilde{\omega}_{B}<\tilde{\omega}$)
Let us prescribe $B_{\rm bg}(\xi)$ and $E_{\rm bg}(\xi)$ by choosing $\kappa^{2}$ of the form
$\kappa^{2}=1+\zeta a^{2},$ (15)
where $\zeta$ is a constant coefficient. The prescribed $\kappa(\xi)$ implies the fastest drift of the fluid in the middle of the packet where $a(\xi)$ peaks (Equation 5); the fluid is static ahead and behind the wave packet where $a=0$. The case of $\zeta=0$ corresponds to a static magnetic background throughout the packet ($E_{\rm bg}=0$); in this case, the particle motion is similar to that found in B22. When $\zeta\neq 0$, the presence of $E_{\rm bg}\neq 0$ changes the particle motion by introducing an additional electric force.
After solving for the oscillating particle motion in the wave packet, we calculate its Lorentz factor averaged over the wave oscillation period. Figure 1 shows the obtained Lorentz factor as a function of position $\xi$ in the packet, in three sample models. The models have the same radio wave (packet with amplitude $a_{\max}=4$) and different $\zeta=0$, 0.4, and 1. The results demonstrate that the average particle Lorentz factor $\tilde{\gamma}(\xi)$ measured in the local fluid frame is the same in all three models, independent of the fluid motion in the packet, and follows the solution found with a static background (Equation 9). In the presented numerical models, we chose $\omega_{B}/\omega=0.05$, where $\omega_{B}\equiv\tilde{\omega}_{B}^{\rm u}$ is the gyro-frequency in the upstream plasma. Then, the condition $\tilde{\omega}_{B}<\tilde{\omega}$ is satisfied throughout the packet, so the heating everywhere acts in the regular regime according to Equation (9).
Figure 1: Lorentz factor $\gamma$ of a test particle moving through the wave packet (averaged over the wave oscillation period). The upstream plasma is at rest ($\gamma=1$) and has the gyro-frequency $\omega_{B}=0.05\omega$. The plasma particle moves through the wave packet in the coordinate $\xi=t-z/c$ from right to left. The simulated packet consists of $10^{3}$ oscillations with an envelope described by Equation (5) with $a_{\max}=4$, and the background field is described by Equations (12) and (15). Different colors correspond to models with $\zeta=0$ (static background, $E_{\rm bg}=0$; black), $\zeta=0.4$ (blue), and $\zeta=1$ (green). For each model, dashed curve shows $\gamma$ measured in the fixed frame (where the upstream plasma is at rest), and solid curve shows $\tilde{\gamma}$ measured in the local fluid rest frame. All solid curves overlap, demonstrating that $\tilde{\gamma}(\xi)$ is independent of the local fluid velocity $\beta_{\rm D}=E_{\rm bg}/B_{\rm bg}$ and follows the solution $\tilde{\gamma}^{2}=1+a^{2}(\xi)$ (red dotted curve).
The test-particle motion changes if $\omega_{B}/\omega$ is increased so that $\tilde{\omega}_{B}/\tilde{\omega}$ exceeds unity somewhere inside the packet, triggering stochastic acceleration. Compression of $\tilde{B}_{\rm bg}$ by the factor of $\kappa$ implies $\tilde{\omega}_{B}=\kappa\omega_{B}$, and the Doppler-shifted wave frequency in the fluid frame is $\tilde{\omega}=\omega/\kappa$. This gives
$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}=\kappa^{2}\,\frac{\omega_{B}}{\omega}.$ (16)
[…]
3.2 Stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$)
First, let us formulate the analytical prediction for stochastic heating by solving Equation (10) for $\tilde{\gamma}(\xi)$. The fluid proper time $\tilde{t}$ is related to time $t$ in the lab frame by $d\tilde{t}=dt/\gamma_{\rm D}$. During time $dt$, the fluid element moves in $\xi$ by $d\xi=(1-\beta_{\rm D})dt$, so
$d\tilde{t}=\frac{d\xi}{\gamma_{\rm D}(1-\beta_{\rm D})}=\kappa\,d\xi.$ (17)
Using also $\tilde{\omega}=\omega/\kappa$ and $\tilde{\omega}_{B}=\kappa\omega_{B}$ (where $\omega_{B}\equiv\tilde{\omega}_{B}^{\rm u}$ is the upstream gyrofrequency), we rewrite Equation (10) as
$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}=\chi a_{0}^{2}\kappa^{2/3}\left(\omega_{B}\omega^{2}\right)^{1/3}.$ (18)
For any chosen $\kappa(\xi)$, Equation (18) can be integrated for $\tilde{\gamma}(\xi)$, and this analytical prediction can be compared with simulations of test particle motion. Unlike regular heating (where $\tilde{\gamma}$ tracks the rise and drop of $a$), stochastic heating gives a monotonically growing mean expectation $\tilde{\gamma}(\xi)$: $d\tilde{\gamma}/d\xi\geq 0$.
Our simulation setup is described above; in particular, $B_{\rm bg}$ and $E_{\rm bg}$ are given by Equation (12). We arbitrarily choose $\kappa(\xi)$ of the form, [^5] Plasma in packets with monotonically growing $\tilde{\gamma}(\xi)$ is expected to have a monotonically growing $\kappa(\xi)$. Therefore, in sample models of stochastic heating we choose $\kappa(\xi)$ that satisfy $d\kappa/d\xi\geq 0$.
$\kappa=\left\{\begin{array}[]{cc}1+(\kappa_{0}-1)\sin^{2}(\pi\xi/T)&\qquad\xi<T/2\\ \kappa_{0}&\qquad\xi>T/2\end{array}\right.$ (21)
so $\kappa(\xi)$ grows until it reaches $\kappa_{0}$ in the middle of the wave packet and stays constant in the second half of the packet. We have calculated particle motion in packets with $\kappa_{0}=1$ (no fluid motion, $\beta_{\rm D}=0$) and $\kappa_{0}=3$.
Figure 2: Average Lorentz factor $\tilde{\gamma}$ (measured in the local fluid rest frame) of particles crossing the wave packet in the regime of stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$). The average $\tilde{\gamma}$ is calculated by tracking $4\times 10^{3}$ test particles. The upstream plasma is at rest ($\gamma_{\rm D}=1$) and has a small temperature $kT=10^{-2}mc^{2}$; its gyro-frequency is $\omega_{B}=2\omega$. The packet $0<\xi<T$ consists of $N=4\times 10^{3}$ oscillations; its envelope $a_{0}=\sqrt{2}a$ is described by Equation (5) with $a_{0}^{\max}=30$ (blue curve). Background field is described by Equations (12) and (21). As one can see in the figure, the plasma experiences heating as it moves through the wave ($\tilde{\gamma}$ grows with $\xi$). Black solid curves show the simulation results for $\kappa_{0}=1$ (static background field, no bulk motion of the plasma) and $\kappa_{0}=3$ ($\beta_{\rm D}^{\max}=0.8$). The results approximately agree with the analytical prediction (Equation (18), red dotted curves) with $\chi=0.66$ for $\kappa_{0}=1$ and $\chi=0.67$ for $\kappa_{0}=3$.
The stochastic nature of heating requires calculations of many realizations of particle motion in the packet. We draw 4000 particles from an initial Maxwellian distribution with a small temperature $kT=10^{-2}mc^{2}$, seeding small initial perturbations for the stochastic behavior. The particle motion is followed for each of the 4000 realizations, and the result is averaged to find the mean expectation for $\tilde{\gamma}(\xi)$.
The results are shown in Figure 2 for $\omega_{B}/\omega=2$ and $a_{0}=30$. In both shown cases, $\kappa_{0}=1$ and $\kappa_{0}=3$, the found $\tilde{\gamma}(\xi)$ is in good agreement with the analytical prediction given by Equation (18) with $\chi\approx 2/3$.
In summary, the results of B22 for regular and stochastic heating hold in the local frame $\tilde{\cal K}$ regardless of fluid motion, as long as the fluid speed varies on timescales much longer than the Larmor period of particles. The Lorentz factor $\tilde{\gamma}\gg 1$ implies a strong heating effect, giving the fluid a relativistic effective temperature. The energized particles gyrate in the plane perpendicular to $\boldsymbol{B}_{\rm bg}$, so the plasma behaves as a two-dimensional relativistic gas. Its internal energy density $U_{\rm p}$ and pressure $P_{\rm p}$ (measured in frame $\tilde{\cal K}$) are
$U_{\rm p}=2P_{\rm p}=(\tilde{\gamma}-1)\tilde{\rho}c^{2},\qquad\tilde{\rho}=\frac{\rho}{\gamma_{\rm D}}.$ (22)
The heated particles gyrate with the Larmor frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}\ll\tilde{\omega}_{B}$. It is below the wave frequency $\tilde{\omega}$ in both heating regimes, $\tilde{\omega}_{B}<\tilde{\omega}$ and $\tilde{\omega}_{B}>\tilde{\omega}$ (B22).
4 MHD equations for fluid motion
In the test-particle numerical experiments described above, we prescribed an arbitrary pattern of the fluid drift speed $\beta_{\rm D}(\xi)$. Our goal now is to find the actual pattern of fluid motion in the radio wave packet. It can be quantified by the profiles of two quantities: $\kappa$ (related to $\beta_{\rm D}$) and $\tilde{C}=\tilde{\rho}/\tilde{\rho}_{\rm u}$ (the proper compression factor). They are equal, $\tilde{C}=\kappa$, when a steady pattern of fluid motion is established in the propagating wave packet, i.e. when $\kappa$ and $\tilde{\rho}$ are functions of $\xi$ only (see below). The full problem, however, involves evolving profiles. Therefore, we will formulate general MHD equations for $\kappa(t,\xi)$ and $\tilde{\rho}(t,\xi)$, allowing time dependence. We first state the equations in the usual variables $t,z$ and later switch to the more convenient coordinates $t,\xi$.
4.1 Conservation laws
Fluid motion in the wave packet obeys conservation of mass, energy, and momentum. They are formulated using mass four-flux $\tilde{\rho}u_{\rm D}^{\nu}$ and stress-energy tensor $T^{\mu\nu}$. Only the $t$ and $z$ components are relevant in the plane-parallel problem.
Mass conservation for an MHD flow (assuming negligible $e^{\pm}$ creation/annihilation) reads
$\partial_{\mu}(\tilde{\rho}u_{\rm D}^{\mu})=\partial_{t}\rho+\partial_{z}(\rho v_{\rm D})=0.$ (23)
Contributions to $T^{\mu\nu}$ are made by the plasma, the MHD electromagnetic field, and the radio wave:
$T^{\mu\nu}=T^{\mu\nu}_{\rm p}+T^{\mu\nu}_{\rm f}+T^{\mu\nu}_{\rm w}.$ (24)
The relevant components of $T^{\mu\nu}$ are $T^{tt}$, $T^{tz}$, and $T^{zz}$. In particular, the wave stress-energy tensor is
$T^{zz}_{\rm w}=T^{tz}_{\rm w}=T^{tt}_{\rm w}=U=\frac{\langle E_{\rm wave}^{2}\rangle}{4\pi}=\frac{m\,\omega^{2}a^{2}}{4\pi r_{e}},$ (25)
where $r_{e}=e^{2}/mc^{2}$ is the classical electron radius.
The plasma stress-energy tensor is
$T^{\mu\nu}_{\rm p}=(\tilde{\rho}c^{2}+U_{\rm p}+P_{\rm p})u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P_{\rm p},$ (26)
where $\eta^{\mu\nu}$ is the Minkowski metric. The plasma pressure $P_{\rm p}$ and internal energy $U_{\rm p}$ (Equation 22) are controlled by the wave strength $a(\xi)$. In particular, $\tilde{\gamma}=\sqrt{1+a^{2}}$ in the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}$).
Relevant components of the stress-energy tensor of the MHD electromagnetic field are
$T^{tt}_{\rm f}=T^{zz}_{\rm f}=\frac{B^{2}+E^{2}}{8\pi},\qquad T^{tz}_{\rm f}=\frac{EB}{4\pi},$ (27)
Hereafter, we drop the bar in $\overline{E}$ and $\overline{B}$ — in all remaining sections, $\boldsymbol{E}$ and $\boldsymbol{B}$ will have the meaning of MHD fields averaged over the Larmor timescale. $T^{\mu\nu}_{\rm f}$ may be re-written in a form similar to ideal fluid:
$T^{\mu\nu}_{\rm f}=(U_{\rm f}+P_{\rm f})u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P_{\rm f}\qquad(\mu,\nu=t,z),$ (28)
where
$U_{\rm f}=P_{\rm f}=\frac{\tilde{B}^{2}}{8\pi},\qquad\tilde{B}=\frac{B}{\gamma_{\rm D}}.$ (29)
Note that the MHD electric field vanishes in the fluid frame (there is no plasma drift in frame $\tilde{\cal K}$ by definition). The magnetic field in the fluid frame $\tilde{B}$ controls the energy density and effective pressure $P_{\rm f}=U_{\rm f}$. Magnetic flux freezing in each fluid element implies
$\frac{\tilde{B}}{B_{\rm u}}=\frac{\tilde{\rho}}{\rho_{\rm u}}.$ (30)
The total MHD stress-energy tensor may be summarized as follows
$T^{\mu\nu}_{\rm MHD}=T^{\mu\nu}_{\rm p}+T^{\mu\nu}_{\rm f}=h\tilde{\rho}c^{2}u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P,$ (31)
where
$h=1+w+\sigma,\qquad P=\tilde{\rho}c^{2}\left(\frac{w}{3}+\frac{\sigma}{2}\right),$ (32)
$w\equiv\frac{U_{\rm p}+P_{\rm p}}{\tilde{\rho}c^{2}}=\frac{3}{2}(\tilde{\gamma}-1),\quad\sigma\equiv\frac{\tilde{B}^{2}}{4\pi\tilde{\rho}c^{2}}=\frac{\tilde{\rho}}{\tilde{\rho}_{\rm u}}\,\sigma_{\rm u}.$ (33)
Note that the stress-energy tensor of the MHD fluid depends on the strength parameter $a$ of the wave it interacts with. In particular, in the regular heating regime, Equation (9) implies that the dimensionless enthalpy $w$ tracks the local strength of the wave $a(\xi)$:
$w\approx\frac{3}{2}a\qquad\left({\rm when~{}}\,\tilde{\omega}_{B}<\tilde{\omega}~{}{\rm and}~{}a\gg 1\right).$ (34)
Interaction of the wave with the MHD fluid obeys energy and momentum conservation,
$\partial_{\nu}\left(T^{\mu\nu}_{\rm MHD}+T^{\mu\nu}_{\rm w}\right)=-\tilde{n}\dot{{\cal E}}_{e}u_{\rm D}^{\mu}.$ (35)
The term $\tilde{n}\dot{{\cal E}}_{e}u^{\mu}$ is the radiative loss rate (per unit volume) of the plasma particles, which oscillate and gyrate in the wave; $\tilde{n}=\tilde{\rho}/m$ is the particle number density and ˙$\dot{{\cal E}}_{e}$ is the emitted power per particle, which is a Lorentz-invariant quantity.
The power ˙$\dot{{\cal E}}_{e}$ can be calculated in the fluid frame $\tilde{\cal K}$. When $\tilde{\omega}_{B}<\tilde{\omega}$, it is given by
$\dot{\gamma}_{e}\equiv\frac{\dot{{\cal E}}_{e}}{mc^{2}}\approx\frac{2}{3}\frac{r_{e}}{c}\tilde{\omega}^{2}a^{2}\tilde{\gamma}^{2}.$ (36)
A similar expression (with the numerical coefficient $4/3$ instead of $2/3$) holds in the regime of stochastic heating, $\tilde{\omega}_{B}>\tilde{\omega}$ (B22). The strength parameter of the radio wave $a$ is invariant, and its frequency $\omega$ and electric field $E_{\rm wave}$ transform to the fluid frame as
$\frac{\tilde{\omega}}{\omega}=\frac{\tilde{E}_{\rm wave}}{E_{\rm wave}}=\kappa^{-1},$ (37)
where $\kappa$ is defined in Equation (13). Equation (36) may also be written in terms of the wave energy density $U$,
$\dot{\gamma}_{e}\approx\frac{\sigma_{\rm T}\tilde{U}}{mc}\,\tilde{\gamma}^{2}=\frac{\sigma_{\rm T}U}{mc\,\kappa^{2}}\,\tilde{\gamma}^{2},$ (38)
where $\sigma_{\rm T}=(8\pi/3)r_{e}^{2}$ is the Thomson cross section.
4.2 MHD equations in coordinates $t,\xi$
The conservation laws can also be stated using fluxes through the surfaces of $\xi=const$, which move with speed $c$. This will considerably simplify the analysis of wave-fluid interaction.
The fluid speed in the $\xi$-coordinate is $d\xi/dt=1-\beta_{\rm D}$, so the mass flux through a surface of $\xi=const$ is
$F_{\rm m}=\rho(c-v_{\rm D})=\frac{c\tilde{\rho}}{\kappa}.$ (39)
Mass conservation in the $t,\xi$-coordinates reads
$\left.\frac{\partial\rho}{\partial t}\right|_{\xi}+\left.\frac{\partial F_{\rm m}}{\partial\xi}\right|_{t}=0.$ (40)
Similarly, the flux of four-momentum ${\cal P}^{\mu}$ through a surface $\xi=const$ is $T^{\mu t}-T^{\mu z}$. Conservation of four-momentum (Equation 35) in the $t,\xi$-coordinates becomes [^6] This equation can also be derived formally by transforming tensor $T^{\mu\nu}$ to new coordinates: $(t,z)\rightarrow(t,\xi)$. The resulting $T^{t\xi}$ and $T^{z\xi}$ are the energy and momentum fluxes crossing $\xi=const$.
$\left.\frac{\partial T^{\mu t}}{\partial t}\right|_{\xi}+\frac{\partial}{\partial\xi}\left(T^{\mu t}-T^{\mu z}\right)_{t}=-\tilde{n}\dot{{\cal E}}_{e}u_{\rm D}^{\mu}.$ (41)
Note that the wave is static in $\xi$, so it creates no flux of energy or momentum across a surface of $\xi=const$:
$T^{\mu t}_{\rm w}-T^{\mu z}_{\rm w}=U-U=0\qquad(\mu=t,z).$ (42)
Only the MHD fluid contributes to the fluxes. Thus, the fluxes of energy and momentum through a surface of $\xi=const$ are
$\frac{F_{\rm en}}{c}$ $=T^{tt}_{\rm MHD}-T^{tz}_{\rm MHD}=\frac{c^{2}\rho h}{\kappa}-P,$ (43)
$F_{\rm mom}$ $=T^{zt}_{\rm MHD}-T^{zz}_{\rm MHD}=\frac{c^{2}\rho h}{\kappa}\beta_{\rm D}-P.$ (44)
4.3 Conservation of ${\cal Q}_{\pm}$
Instead of energy ${\cal E}$ and momentum ${\cal P}$ it will be helpful to use quantities
${\cal Q}_{\pm}={\cal E}\pm c{\cal P}.$ (45)
The fluxes of ${\cal Q}_{\pm}$ through a surface of $\xi=const$ are
$F_{-}$ $=F_{\rm en}-cF_{\rm mom}=\frac{c^{3}\tilde{\rho}h}{\kappa^{2}},$ (46)
$F_{+}$ $=F_{\rm en}+cF_{\rm mom}=c^{3}\tilde{\rho}h-2cP.$ (47)
Conservation of ${\cal Q}_{\pm}$ is given by the sum/difference of equations describing energy and momentum conservation (Equation (41) with $\mu=t$ and $z$):
$\left.\frac{\partial}{\partial t}\left(T^{tt}\pm T^{tz}\right)\right|_{\xi}+\frac{\partial}{\partial\xi}\left(\frac{F_{\pm}}{c}\right)_{t}=-\tilde{n}\dot{{\cal E}}_{e}(u_{\rm D}^{t}\pm u_{\rm D}^{z}).$ (48)
Note that the stress-energy tensor of the radio wave disappears from the conservation law for ${\cal Q}_{-}$. This has a simple physical reason: the energy and momentum of the wave satisfy ${\cal E}_{\rm w}=c{\cal P}_{\rm w}$, so the wave carries ${\cal Q}_{-}=0$. There is no exchange of ${\cal Q}_{-}$ in the wave-fluid interaction: whenever the radio wave loses energy $\Delta{\cal E}$ to the MHD fluid, it also loses momentum $\Delta{\cal P}=\Delta{\cal E}/c$, so
$\Delta{\cal Q}_{-}=\Delta{\cal E}-c\,\Delta{\cal P}=0.$ (49)
This observation significantly simplifies the problem. As long as the shape of the propagating radio wave $a(\xi)$ is approximately steady (i.e. the wave is not damped), the fluid dynamics in the packet is fully described by two equations that express conservation of mass and ${\cal Q}_{-}$:
$\partial_{t}\rho$ $=-\partial_{\xi}\left(\frac{\tilde{\rho}}{\kappa}\right),$ (50)
$\partial_{t}\left(\frac{\rho h}{\kappa}-\frac{P}{c^{2}}\right)$ $=-\partial_{\xi}\left(\frac{\tilde{\rho}h}{\kappa^{2}}\right)-\frac{\tilde{\rho}}{\kappa}\dot{\gamma}_{e}.$ (51)
The radio wave does not contribute to these conservation laws, because it carries no mass or ${\cal Q}_{-}$. The wave strength parameter $a(\xi)$ still enters the ${\cal Q}_{-}$ equation through $h=1+w+\sigma$ (since $a$ controls the fluid enthalpy $w$) and through $\dot{\gamma}_{e}\propto a^{2}$. As long as $a(\xi)$ is steady, it can be treated as a given function in the wave-fluid interaction problem, so Equations (50) and (51) form a closed set for two unknowns $\rho(t,\xi)$ and $\kappa(t,\xi)$. For a given radio wave packet, the equations can be solved in the domain $0<\xi<T$ with given initial conditions $\rho(0,\xi)$ and $\kappa(0,\xi)$, and boundary conditions $\rho(t,0)=\rho_{\rm u}$ and $\kappa(t,0)=\kappa_{\rm u}$.
Equations (50) and (51) allow one to focus on the fluid dynamics in the wave packet without considering the evolution of the radio wave itself, assuming that the evolution of $a(\xi)$ is negligible on timescales of interest. The consistency of this assumption can be checked later using the remaining independent equation that states conservation of ${\cal Q}_{+}$ and governs the evolution of $a(t,\xi)$. A simpler consistency check can be done by comparing the energy exchanged in the wave-plasma interaction with the wave energy (Section 10).
Note also that Equations (50) and (51) are Lorentz-invariant: they have the same form in any frame boosted along $z$ [^7] Lorentz transformation of initial conditions is not, however, straightforward, since hypersurfaces of simultaneity ($t=const$) are different in different frames. They can be solved e.g. in the frame where the upstream plasma is at rest: $\kappa_{\rm u}=1$ and $\rho_{\rm u}=\tilde{\rho}_{\rm u}$. The problem can also be solved in other frames where $\kappa_{\rm u}\neq 1$ and the upstream plasma has speed $\beta_{\rm u}=(\kappa_{\rm u}^{2}-1)/(\kappa_{\rm u}^{2}+1)$, Lorentz factor $\gamma_{\rm u}=(\kappa_{\rm u}^{2}+1)/2\kappa_{\rm u}$, and proper density $\tilde{\rho}_{\rm u}=\rho_{\rm u}/\gamma_{\rm u}$.
5 Steady compression front
The pattern of fluid motion inside the propagating wave packet will be called below the “front” for brevity. Its structure is described by the profiles of $\rho(t,\xi)$ and $\kappa(\xi,t)$, which are governed by Equations (50) and (51). In general, the front structure can evolve in time, as demonstrated in the next section. Here, we first show that there is a unique steady flow pattern, where all quantities depend on $\xi$ only, with no evolution with $t$. As verified in section 6, the steady state is an attractor for the front evolution. It is reached under two conditions: (1) the front propagation time is long enough for the flow pattern to relax to the steady state, and (2) the propagation time is short enough to neglect damping of the wave, so $a(\xi)$ can be treated as constant in time.
The steady front solution is found by setting $\left.\partial_{t}\right|_{\xi}=0$ and $\partial_{\xi}=d/d\xi$ in the MHD equations. Then, Equations (50) and (51) give
$\frac{d}{d\xi}\left(\frac{\tilde{\rho}}{\kappa}\right)$ $=0\quad\Rightarrow\quad\tilde{\rho}=\frac{\kappa}{\kappa_{\rm u}}\tilde{\rho}_{\rm u},$ (52)
$\frac{d}{d\xi}\left(\frac{h}{\kappa}\right)$ $=-\dot{\gamma}_{e}.$ (53)
The corresponding lab-frame density $\rho=\gamma_{\rm D}\tilde{\rho}$ can be expressed using
$\gamma_{\rm D}=\frac{\kappa^{2}+1}{2\kappa},$ (54)
so the solution $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}$ gives
$\frac{\rho}{\rho_{\rm u}}=\frac{1-\beta_{\rm u}}{1-\beta}=\frac{\kappa^{2}+1}{\kappa_{\rm u}^{2}+1}.$ (55)
The solution $\tilde{\rho}/\rho_{\rm u}=\kappa/\kappa_{\rm u}$ also implies $\sigma=\sigma_{\rm u}\kappa/\kappa_{\rm u}$ and $d/d\xi(h/\kappa)=d/d\xi[(1+w)/\kappa]$. Thus, Equation (53) becomes
$\frac{d}{d\xi}\left(\frac{1+w}{\kappa}\right)=-\dot{\gamma}_{e}=-\frac{\sigma_{\rm T}U}{mc}\,\frac{\tilde{\gamma}^{2}}{\kappa^{2}}.$ (56)
Note that the fluid magnetization $\sigma$ has dropped out. Remarkably, the fluid motion and compression in a steady front are independent of $\sigma_{\rm u}$. In particular, in the case of negligible radiative losses, we find
$\kappa=(1+w)\kappa_{\rm u}\qquad({\rm when}~{}\,\dot{\gamma}_{e}=0),$ (57)
where we used the cold upstream condition $w_{\rm u}=0$.
One can also solve for $\kappa(\xi)$ without neglecting radiative losses. In particular, consider the simpler regime of regular heating ($\tilde{\omega}_{B}<\tilde{\omega}_{B}$). We are interested in waves with $a\gg 1$, which heat the plasma to $w\approx(3/2)\tilde{\gamma}\gg 1$. Equation (56) can be integrated for $(1+w)/\kappa$, and we find
$\kappa=(1+w)\kappa_{\rm u}(1+q),$ (58)
where
[…]
The importance of radiative losses is controlled by the (Lorentz-invariant) dimensionless parameter $q$, which is determined by the column energy density of the wave packet, $\Sigma$. Equation (57) holds when $q\ll 1$.
6 Relaxation to the steady state
A simple argument shows that the timescale to approach the steady state is long if the plasma magnetization $\sigma_{\rm u}$ is large. Consider the vacuum limit of $\rho_{\rm u}\rightarrow 0$ (which corresponds to $\sigma_{\rm u}\rightarrow\infty$). Then, the background field must remain unchanged from its upstream state: in a vacuum, the electromagnetic wave goes through the background field without interaction, so no $E_{\rm bg}$ is generated. Thus, at $\sigma_{\rm u}\rightarrow\infty$ the magnetized plasma must remain static and will not be compressed by the radio wave packet. The steady-state front with large $C$ still provides a consistent solution, but relaxing to it takes an infinite time, so it becomes irrelevant at $\sigma_{\rm u}\rightarrow\infty$.
A similar conclusion is reached by considering energy. Compression of the magnetically dominated plasma requires deposition of energy $\propto\sigma_{\rm u}$, much greater than the energy of plasma particles that interact with the wave. The background field serves as a ballast coupled to the plasma. It increases the effective mass per particle from $m$ to $\sigma_{\rm u}m$, which makes it heavy and impossible to accelerate in the limit of $\sigma_{\rm u}\rightarrow\infty$. An analogy can be drawn with collision of a photon of energy ${\cal E}_{\rm ph}$ with a particle of mass $\sigma m$. Using the frame where the particle is initially at rest, one finds that the particle gains speed $\beta\approx{\cal E}_{\rm ph}/\sigma mc^{2}\ll 1$ if $\sigma\gg{\cal E}_{\rm ph}/mc^{2}$.
It is worth looking in more detail at a similar collision problem formulated for the wave-plasma system.
6.1 Wave-plasma collision
Suppose a radio wave packet of energy ${\cal E}$ passes through an initially static, cold slab of plasma with magnetization $\sigma_{\rm u}$. Some energy $\Delta{\cal E}$ and momentum $\Delta{\cal E}/c$ are lost by the radio wave and received by the slab. As a result, the slab is accelerated to speed $\beta_{\rm D}$ and compressed by the factor $1-\beta_{\rm D}$ (cf. Equation (55)). Compression increases the slab magnetic field from initial $B_{\rm u}$ to $B=B_{\rm u}/(1-\beta_{\rm D})$.
In the simplified collision picture, the slab is treated as a single body characterized by its volume $V$, particle number $N$, inertial mass of the plasma $M=Nm(1+w)$ (increased from $M_{\rm u}=Nm$ by deposited heat), magnetic field $B$, and electric field $E=\beta_{\rm D}B$. Energy conservation gives
$M_{\rm u}c^{2}+V_{\rm u}\frac{B_{\rm u}^{2}}{8\pi}+\Delta{\cal E}=\gamma_{\rm D}Mc^{2}+V\frac{B^{2}+E^{2}}{8\pi}+{\cal E}_{\rm rad},$ (60)
where we included possible radiative losses of the plasma ${\cal E}_{\rm rad}$. The radiated momentum ${\cal P}^{\prime}_{\rm rad}=0$ in the fluid rest frame implies ${\cal P}_{\rm rad}=\beta_{\rm D}{\cal E}_{\rm rad}/c$ in the lab frame. Momentum conservation equation (multiplied by $c$) is
$\Delta{\cal E}=\gamma_{\rm D}\beta_{\rm D}Mc^{2}+V\frac{EB}{4\pi}+\beta_{\rm D}{\cal E}_{\rm rad}.$ (61)
Subtracting the momentum equation from the energy equation one can exclude $\Delta{\cal E}$:
$M_{\rm u}c^{2}+V_{\rm u}\frac{B_{\rm u}^{2}}{8\pi}=\frac{Mc^{2}}{\kappa}+V\frac{(B-E)^{2}}{8\pi}+(1-\beta_{\rm D}){\cal E}_{\rm rad}.$ (62)
We now substitute $E=\beta_{\rm D}B$, $B=B_{\rm u}/(1-\beta_{\rm D})$, $VB=V_{\rm u}B_{\rm u}$, and $M=(1+w)M_{\rm u}$, and obtain
$1=\frac{1+w}{\kappa}-\frac{\beta_{\rm D}\sigma_{\rm u}}{2}+(1-\beta_{\rm D})\epsilon_{\rm rad},$ (63)
where
$\sigma_{\rm u}\equiv\frac{V_{\rm u}B_{\rm u}^{2}}{4\pi M_{\rm u}c^{2}},\qquad\epsilon_{\rm rad}\equiv\frac{{\cal E}_{\rm rad}}{M_{\rm u}c^{2}}.$ (64)
Radiative losses can be described by the effective scattering cross section $\sigma_{\rm sc}$ per electron, with $\sigma_{\rm sc}\sim w^{2}\sigma_{\rm T}$ within a factor of $\sim 2$ (B22). It implies
$\epsilon_{\rm rad}\sim\frac{\sigma_{\rm T}w^{2}\Sigma}{mc^{2}},$ (65)
where $\Sigma$ is the column energy density of the radio wave packet.
Using the identity $1-\beta_{\rm D}=2/(\kappa^{2}+1)$ one can rewrite Equation (63) as a cubic equation for $\kappa$. Its solution simplifies in two limits: $\kappa\gg 1$ ($\beta_{\rm D}\approx 1$) and $\beta_{\rm D}\ll 1$ ($\kappa\approx 1$). Assuming $\sigma_{\rm u}\gg 1$, we find
$\kappa-1\approx\left\{\begin{array}[]{lr}\vspace*{1mm}\displaystyle{\frac{2}{\sigma_{\rm u}}(w+\epsilon_{\rm rad})}&{\rm when}\;\displaystyle{w+\epsilon_{\rm rad}\ll\sigma_{\rm u}}\\ \displaystyle{\frac{w}{\sigma_{\rm u}}+\sqrt{\frac{\epsilon_{\rm rad}}{\sigma_{\rm u}}}}&{\rm when}\;\displaystyle{w+\epsilon_{\rm rad}\gg\sigma_{\rm u}}\end{array}\right.$ (66)
It also gives the proper compression factor $\tilde{C}\approx\kappa$.
Once $\kappa$ is known, we can also find the energy $\Delta{\cal E}$ taken from the radio wave packet. Equation (61) gives
$\frac{\Delta{\cal E}}{M_{\rm u}c^{2}}=\beta_{\rm D}\left[\gamma_{\rm D}(1+w+\sigma_{\rm u}\kappa)+\epsilon_{\rm rad}\right],$ (67)
and using Equation (66) we find
$\frac{\Delta{\cal E}}{M_{\rm u}c^{2}}\approx\left\{\begin{array}[]{lr}\vspace*{1mm}\displaystyle{2(w+\epsilon_{\rm rad})},&\displaystyle{w+\epsilon_{\rm rad}\ll\sigma_{\rm u}}\\ \displaystyle{\frac{w^{2}}{\sigma_{\rm u}}+\frac{3}{2}w\sqrt{\frac{\epsilon_{\rm rad}}{\sigma_{\rm u}}}+\frac{3}{2}\epsilon_{\rm rad}},&\displaystyle{w+\epsilon_{\rm rad}\gg\sigma_{\rm u}}\end{array}\right.$ (68)
The upstream plasma is composed of many slabs. One after another, they fill the propagating packet and exit behind it, and one might conclude that Equation (66) approximately describes $\kappa$ in a steady propagating front. In fact, Equation (66) is very far from the correct MHD steady-state solution given in section 5. When viewed in the upstream rest frame, a non-radiative steady front should have $\kappa=1+w$ regardless of $\sigma_{\rm u}$ (Equation 57). In contrast, the collision model described above yields $\kappa\approx 1+(2w/\sigma_{\rm u})\approx 1$ when $\sigma_{\rm u}\gg w$.
The simplified collision model (Equation 66) works well at the beginning of packet propagation through the plasma; however, it fails to describe the gradual relaxation of the compression front to the correct steady state after many interactions with many plasma slabs. When $\sigma_{\rm u}\gg 1$, this relaxation takes a long time and occurs due to a slow growth of magnetic pressure in the front, as demonstrated below. The grown magnetic pressure drops out from the final steady pattern of the MHD flow because it occurs with a zero gradient of $\sigma/\kappa$ (section 5). This behavior is not captured by the simplified collision model, which treats each plasma slab passing through the packet as a single body and does not consider any gradients. By contrast, in full MHD, the gradients across the front play an essential role: they endow the propagating front with a “memory” of its earlier interactions with the fluid that has been disposed behind the front. The accumulating memory enables the front relaxation to the steady state.
Below we calculate this relaxation by solving the full time-dependent MHD problem described by Equations (50) and (51). We solve it first in the non-radiative regime (neglecting $\dot{\gamma}_{e}$ in Equation (51)) and then in the radiative regime. In the remainder of this section, we assume regular heating ($\tilde{\omega}_{B}>\tilde{\omega}$). A similar relaxation trend holds with stochastic heating ($\tilde{\omega}_{B}<\tilde{\omega}$); however, this regime gives a huge enthalpy $w$ and relaxation takes a very long time, so steady fronts with stochastic heating are not expected in FRBs (section 9).
6.2 Evolution of a non-radiative compression front
Before solving numerically the time-depending MHD equations, we make a heuristic estimate of the timescale $t_{\star}$ for the front relaxation toward the steady state.
First, note the time it takes a fluid element to cross the entire wave packet $0<\xi<T$:
[…]
In particular, in a steady state, $\kappa=\kappa_{\star}\approx 1+(3/2)a$. Then, for a packet with $a_{\max}\gg 1$, one finds
[…]
Hereafter, star indicates the steady state.
Now consider the momentum per particle carried by the MHD fluid inside the wave packet,
${\cal P}_{\rm MHD}=\frac{T^{tz}_{\rm MHD}}{nc}=h\gamma_{\rm D}\beta_{\rm D}mc.$ (71)
It grows when the fluid receives momentum from the wave. Recall that the wave interacts with the plasma particles only, and the plasma shares the received momentum with the background field (through gyration). Therefore, the budget for the wave-fluid momentum exchange (per particle crossing the packet) is set by $T^{tz}_{\rm p}$ rather than $T^{tz}_{\rm MHD}$:
$\Delta{\cal P}=\frac{T^{tz}_{\rm p}}{nc}=(1+w)\gamma_{\rm D}\beta_{\rm D}mc.$ (72)
When the fluid flow in the packet is far from the steady state, $\kappa\ll\kappa_{\star}$, the momentum flux in $\xi$ is non-uniform, $\partial_{\xi}F_{\rm mom}\neq 0$, and momentum comparable to $\Delta{\cal P}$ is retained inside the propagating front $0<\xi<T$, supporting $(\partial_{t}\kappa)_{\xi}>0$ and $(\partial_{t}{\cal P}_{\rm MHD})_{\xi}>0$.
Thus, as long as $\kappa\ll\kappa_{\star}$, one may expect each particle crossing the wave packet to deposit $\sim\Delta{\cal P}$ on the timescale $t_{\rm cross}$ toward future fluid momentum in the propagating front. Then, the timescale to grow ${\cal P}_{\rm MHD}$ is
$t_{{\cal P}}\sim\frac{{\cal P}_{\rm MHD}}{\Delta{\cal P}}t_{\rm cross}\sim\frac{h}{1+w}t_{\rm cross},$ (73)
where $h=1+w+\sigma$. Let us now specialize to the frame where the upstream fluid is at rest ($\kappa_{\rm u}=1$), and the steady state has $\kappa=\kappa_{\star}=1+w$ and $\sigma=\kappa_{\star}\sigma_{\rm u}=(1+w)\sigma_{\rm u}$. In this frame, the expected timescale to approach the steady state is
$t_{\star}\sim\frac{1+w+\kappa_{\star}\sigma_{\rm u}}{1+w}\,t_{\rm cross}^{\star}=(1+\sigma_{\rm u})t_{\rm cross}^{\star}\quad(\kappa_{\rm u}=1).$ (74)
Figure 3: Top: relaxation of the non-radiative compression front $C(\xi)$ toward the steady-state $C_{\star}(\xi)$ (red curve, Equation (76)) from an initial state with a uniform static plasma, $C=1$ and $\kappa=1$. Black curves show snapshots of the front, with time $t$ indicated next to each curve (in units of packet duration $T$). The wave packet (Equation 5) has $a_{\max}=(4/3)\sqrt{7}$, equivalent to $w_{\max}=4$. The upstream magnetization is $\sigma_{\rm u}=10$. Bottom: evolution of $\kappa$ from the initial $\kappa=1$ toward $\kappa_{\star}$ in the same simulation.
This simple estimate agrees with our numerical solutions for the front evolution. We find the evolution by solving Equations (50) and (51) for $\kappa(t,\xi)$ and $\rho(t,\xi)$ in a given wave packet $a(\xi)$ propagating in a background plasma with a given upstream magnetization $\sigma_{\rm u}$. The fluid enthalpy inside the packet, $w\equiv(U_{\rm p}+P_{\rm p})/\tilde{\rho}c^{2}$, is determined by Equations (9) and (22):
$w(\xi)=\frac{3}{2}\left[\sqrt{a^{2}(\xi)+1}-1\right].$ (75)
In sample models, we use the packet described by Equation (5) with $a_{\max}=(4/3)\sqrt{7}$ which corresponds to $w_{\max}=4$. As an initial state at $t=0$, we take a static plasma, $\kappa(\xi)=1$ and $\rho(\xi)=\rho_{\rm u}$, and follow the evolution of the flow pattern inside the propagating wave packet until it relaxes to the steady-state solution. In the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), the steady-state solution is
$\kappa_{\star}=1+w,\quad\rho_{\star}=\frac{\kappa_{\star}^{2}+1}{2}\rho_{\rm u}\qquad(\kappa_{\rm u}=1).$ (76)
The model with $\sigma_{\rm u}=10$ is shown in Figure 3. At the beginning of the simulation we observe that $\kappa(\xi)$ starts to grow from unity and develops a peaked profile — the MHD fluid is accelerated in the first half of the packet and decelerated in the second half, before exiting behind the packet. Over time, $\kappa(\xi)$ approaches $\kappa_{\star}(\xi)$, which is an attractor for the flow pattern. In particular, the maximum of the $\kappa$-profile, $\kappa_{\max}$, eventually approaches the expected $\kappa_{\max}^{\star}=1+w_{\max}=5$. The maximum compression factor $C_{\max}=\rho_{\max}/\rho$ approaches
$C_{\max}^{\star}=\frac{1}{2}\left[(1+w_{\max})^{2}+1\right].$ (77)
The evolution of compression peak $C_{\max}(t)$ provides a simple way to quantify the front relaxation toward the steady-state. The evolution of $C_{\max}$ is shown in Figure 4 for $\sigma_{\rm u}=3$, 10, 20. One can see that the timescale $t_{\star}$ for approaching the steady-state $C_{\max}^{\star}=13$ increases with $\sigma_{\rm u}$ in agreement with Equation (74). The slow evolution occurs through a sequence of quasi-steady states with a nearly uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)\approx const$. It implies a correlated growth of $\kappa$ and $C$: the relation $C\approx(\kappa^{2}+1)/2$ is sustained during the evolution (Figure 5).
Figure 4: Evolution of the compression peak $C_{\max}$ from the initial $C_{\max}=1$ to the steady-state $C_{\max}^{\star}=13$ (indicated by the horizontal dashed line) for the non-radiative fronts with $\sigma_{\rm u}=3$, 10, and 20.
Figure 5: Relation between $\kappa=\gamma_{\rm D}(1+\beta_{\rm D})$ and $C=\rho/\rho_{\rm u}$ (here measured at the peak inside the compression front) during relaxation toward the final steady state. Black curve shows the simulation result (the model with $\sigma_{\rm u}=20$), and red dashed curve shows the relation $C=(\kappa^{2}+1)/2$, equivalent to $\tilde{\rho}/\kappa=\rho_{\rm u}$ (a uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)$). Fluid density in the front closely tracks the steady-state solution $\tilde{\rho}=\kappa\rho_{\rm u}$ defined for an instantaneous profile of $\kappa(\xi)$, which slowly grows during relaxation. The correlated growth of $\kappa$ and $C$ stops when they reach $\kappa_{\max}=5$ and $C_{\max}=13$. Then, the fluxes $F_{\rm m}$ and $F_{-}$ both become exactly uniform across the front, and a true steady state is achieved.
We conclude that at any large but finite $\sigma_{\rm u}$, the wave packet approaches the steady state on a long timescale $t_{\star}\sim\sigma_{\rm u}t_{\rm cross}$ (assuming negligible damping of the radio wave during this time). The conclusion might seem counter-intuitive: one might expect that the front has no memory on timescales longer than $t_{\rm cross}$ since the plasma interacting with the wave packet is continually refreshed on the timescale $t_{\rm cross}$. Nevertheless, the conservation laws require the propagating wave to retain some memory of its previous interactions with the plasma. Given enough time, the packet gradually builds up the strong compression $\tilde{\rho}/\rho_{\rm u}=\kappa\approx w$ — the pattern of the MHD flow in the wave is “groomed” toward the steady state. Remarkably, this gradual evolution leads to a strong magnetic field in the packet, so that the plasma magnetization approaches $\sigma\approx w\sigma_{\rm u}$. Thus, tremendous magnetic pressure develops in the propagating front while it finds a self-consistent steady pattern. At the same time, the final pattern of fluid motion established in the wave packet is the same as it would be in the absence of magnetic pressure, as the background magnetic field drops out from the steady-state solution $\kappa_{\star}(\xi)$.
6.3 Evolution of a radiative compression front
In the regime of $q\gg 1$, the fluid element moving through the wave packet experiences strong radiative losses. The radio wave sustains the fluid enthalpy $w\approx(3/2)a$ by continually depositing energy to offset the losses (B22). Thus, the received power per particle (measured in the local fluid frame) equals the emitted power ˙$\dot{{\cal E}}_{e}$ (Lorentz invariant). The wave-fluid interaction is basically a scattering process with the cross section $\sigma_{\rm sc}=\dot{{\cal E}}_{e}/F$, where $F=cU$ is the energy flux of the radio wave.
The scattering is accompanied by a radiative pressure force $\dot{\cal P}$, which is Lorentz-invariant and easiest to find in the local fluid frame where $\dot{\cal P}=\dot{{\cal E}}_{e}/c$, so
$\dot{\cal P}=\frac{\dot{{\cal E}}_{e}}{c}=\sigma_{\rm T}\tilde{U}\,\tilde{\gamma}^{2}=\frac{\sigma_{\rm T}U}{\kappa^{2}}\,\tilde{\gamma}^{2}.$ (78)
A fluid element crossing the front gains from scattering the following net momentum per particle:
[…]
$\sim\left(1+\frac{1}{\kappa^{2}}\right)\sigma_{\rm T}Ua^{2}T,$ (79)
with $a\sim a_{\max}$. This sets the budget for the fluid momentum exchange with the wave per crossing time $t_{\rm cross}$. The gained momentum is shared with the magnetic field frozen in the fluid, increasing ${\cal P}_{\rm MHD}$.
Similarly to the non-radiative case, one can make a heuristic timescale estimate for the growth of ${\cal P}_{\rm MHD}$ in the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$):
$t_{\cal P}\sim\frac{{\cal P}_{\rm MHD}}{\Delta{\cal P}_{\rm sc}}\,t_{\rm cross}\sim\frac{h\kappa}{qw^{2}}\,t_{\rm cross},$ (80)
where $q$ is evaluated using the entire column energy of the wave packet $\Sigma(T)$ (Equation 59). Equation (80) gives the timescale for approaching the steady state, $t_{\star}=t_{\cal P}(\kappa_{\star})$ if we substitute $\kappa\sim\kappa_{\star}\approx qw$ (Equation 58) and $h_{\star}=1+w+\kappa_{\star}\sigma_{\rm u}\approx\kappa_{\star}\sigma_{\rm u}$:
$t_{\star}\sim q\,\sigma_{\rm u}t_{\rm cross}\qquad(q\gg 1,\;\,\sigma_{\rm u}\gg 1).$ (81)
This estimate may be compared with the results of numerical simulations presented below.
Front relaxation to the steady state in the radiative regime is found by solving Equations (50) and (51) similarly to section 6.2, but now keeping the term $\dot{\gamma}_{e}\neq 0$. This term may be expressed as follows
$\dot{\gamma}_{e}={\cal R}\frac{w^{4}}{\kappa^{2}}\quad(w\gg 1),\qquad{\cal R}\equiv\left(\frac{2}{3}\right)^{5}\frac{r_{e}}{c}\,\omega^{2}.$ (82)
The value of ${\cal R}$ depends on the choice of a reference frame, which can be parameterized by the upstream motion $\beta_{\rm u}$ or $\kappa_{\rm u}$. Note that $\kappa_{\rm u}<1$ ($\beta_{\rm u}<0$) in frames moving in the $+z$ direction relative to the upstream plasma. The wave frequency $\omega$ in a chosen frame scales as $\omega\propto\kappa_{\rm u}$, so ${\cal R}\propto\kappa_{\rm u}^{2}$. Note also that the packet width $T$ depends on the choice of a frame as $\kappa_{\rm u}^{-1}$ (since $\omega T=inv$), and the radiative parameter $q\propto T{\cal R}/\kappa_{\rm u}$ is invariant.
We use the following technical tricks to avoid two complications:
(1) Radiative losses $\dot{\gamma}_{e}$ may be important only when $w\gg 1$ because of the small value of ${\cal R}$ for real FRBs. We avoid simulating huge $w$ and $\kappa$ by rescaling ${\cal R}$ to larger values, so that we can experiment with the radiative regime $q\gg 1$ at moderate $w$ and $\kappa$ accessible to simulations. In sample models, we use the wave packet described by Equation (5) with $a_{\max}=(4/3)\sqrt{7}$ (which gives $w_{\max}=4$) and set the invariant radiative parameter $T{\cal R}/\kappa_{\rm u}=1$. This choice gives $q\sim 3$.
(2) Both $\kappa$ and $\rho$ grow by a large factor during relaxation toward the steady state. The simulation is more efficient if performed in a boosted frame where $\kappa_{\rm u}\ll 1$. This helps avoid ultra-relativistic motions of the accelerated fluid along the wave propagation direction (large $\kappa\gg 1$) that would slow down the simulation. We use the frame where $\kappa_{\rm u}=0.05$. When presenting the results, we transform the fluid parameters to the upstream rest frame (where $\kappa_{\rm u}=1$), so figures will show $\kappa$ and $C=\rho/\rho_{\rm u}$ measured in the upstream frame.
Figure 6: Evolution of the compression peak $C_{\max}$ from the non-radiative steady state ($C_{\max}=C_{\max}^{\star}=13$ at $t=0$) after switching on radiative losses with $T{\cal R}/\kappa_{\rm u}=1$ (the radiative parameter ${\cal R}$ is defined in Equation (82)). The front evolves toward the new steady state with $C_{\max}^{\star}\approx 147$. The evolution is shown for three simulations with upstream magnetizations $\sigma_{\rm u}=3$, 10, 20. The initial and final states are independent of $\sigma_{\rm u}$ and shown in Figure 7.
Figure 7: The initial and final profiles of the compression front in the three simulations shown in Figure 6. The initial state is the steady non-radiative front (Figure 3, red curve); here it is shown by the black dashed curve. Radiative losses make the front evolve until it reaches a new steady state (solid curve), which reproduces the solution in Equation (58).
As an initial condition at $t=0$, we take the MHD fluid in the non-radiative steady front. In the rest frame of the upstream plasma, it is given by $\kappa_{0}(\xi)=1+w(\xi)$ and $C_{0}(\xi)=(\kappa_{0}^{2}+1)/2$. In particular, the initial compression peak is $C_{\max}(0)=13$. In the presence of radiative losses, this flow pattern is no longer a steady state, so it immediately begins to evolve at $t>0$. Both $\kappa$ and $C$ grow and eventually relax to a new steady state.
Figure 6 shows the evolution of $C_{\max}$, and Figure 7 shows the final profiles of $\kappa(\xi)$ and $C(\xi)$. The relaxation is determined by the chosen radiative coefficient ${\cal R}$ and the corresponding $q\sim 3$. In particular, $\kappa_{\max}$ grows from its initial value $51+w_{\max}=$ 5 to the final value $\kappa_{\max}=17$, consistent with the factor of $q\sim 3$, and $C_{\max}$ grows from 13 to $147$, confirming the expected increase $\sim q^{2}$.
The simulations also confirm the expected scaling of the relaxation timescale $t_{\star}\propto q\sigma_{\rm u}$ at $q\gg 1$ and $\sigma_{\rm u}\gg 1$ (Equation 81). This scaling can be seen in Figure 6, which shows three models with $\sigma_{\rm u}=3$, 10, and 20. Note that $t_{\rm cross}\propto\kappa^{2}\propto q^{2}w^{2}$, and hence $t_{\star}$ should scale as $q^{3}$ at $q\gg 1$. One can see that the relaxation timescale is indeed increased by a large factor $\sim q^{3}\sim 30$ compared with the non-radiative front evolution shown in Figure 4.
As expected, the final steady state is independent of $\sigma_{\rm u}$: it is the same in all three simulations with $\sigma_{\rm u}=3,10,20$. The steady-state profiles of $\kappa(\xi)$ and $C(\xi)$ are shown in Figure 7 and agree with Equation (58). Note that in the non-radiative regime ($q\ll 1$) the fluid comes back to rest as it exits behind the wave packet. In contrast, the fluid behind the radiative front has a large speed $\beta_{\rm D}>0$ ($\kappa\gg 1$). This is a consequence of momentum deposition by scattering during the fluid motion through the wave packet.
7 Parameters of the wind
The obtained picture of a radio wave interaction with MHD fluid can now be applied to FRBs propagating through a radial wind. The wind serves as the upstream medium for the wave packet, and we begin with a brief summary of its parameters. They are denoted with subscript “u” for consistency of notation in this paper.
The wind Lorentz factor $\gamma_{\rm u}$ and magnetization $\sigma_{\rm u}$ are estimated in Beloborodov (2020) (denoted there as $\Gamma_{\rm w}$ and $\sigma_{\rm w}$). Their typical values are $\sigma_{\rm u}\sim 10^{2}-10^{3}$ and $\gamma_{\rm u}\sim 10-30$ at radii $r\gtrsim 10^{11}\,$cm. The wind is also characterized by the particle outflow rate $\dot{\cal N}$ and power $L_{\rm w}$, which is dominated by the Poynting flux. Its isotropic equivalent (power per unit solid angle multiplied by $4\pi$) is
$L_{\rm w}\approx cr^{2}B_{\rm u}^{2}=\eta mc^{2}\dot{\cal N},\qquad\eta=\sigma_{\rm u}\gamma_{\rm u}.$ (83)
A spinning magnetar emits a spindown wind with power
$L_{\rm w}\approx\frac{\mu^{2}\Omega^{4}}{c^{3}}\approx 5.8\times 10^{37}\,\mu_{33}^{2}P^{-4}\;{\rm erg\,s}^{-1},$ (84)
where $\mu$ is the magnetic dipole of the magnetar and $P=2\pi/\Omega$ is its rotation period (in seconds). The spindown power gives a lower bound on $L_{\rm w}$, which can be enhanced during magnetar activity.
The wind forms at the light cylinder $R_{\rm LC}=c/\Omega$ and its Lorentz factor grows as $\gamma_{\rm u}(r)\approx r/R_{\rm LC}$ until it reaches $\eta^{1/3}$; at larger radii, $\gamma_{\rm u}$ grows logarithmically (Lyubarsky & Kirk, 2001). We will use a simple approximation:
$\gamma_{\rm u}\approx\left\{\begin{array}[]{cc}\vspace*{2mm}\displaystyle{\frac{r}{R_{\rm LC}}}&\quad r<\eta^{1/3}R_{\rm LC}\\ \displaystyle{\left[\frac{\eta}{\ln 2}\ln\left(1+\frac{r}{\eta^{1/3}R_{\rm LC}}\right)\right]^{1/3}}&\quad r>\eta^{1/3}R_{\rm LC}\\ \end{array}\right.$ (85)
Our description of the wind is imprecise at small radii $r\sim R_{\rm LC}$ for two reasons. (1) The wind magnetic field is assumed to be transverse to the radial direction. In fact, it has a radial component $B_{r}$ (which quickly decreases at $r>R_{\rm LC}$). (2) The wind speed is taken equal to the drift speed. This approximation would be strongly violated for ordinary radio pulsars, where plasma outflows at $R_{\rm LC}$ with a high Lorentz factor; however, it is reasonable for magnetars whose radiation field makes the outflow mildly relativistic (Beloborodov, 2020). Our wind description should hold outside a few $R_{\rm LC}$.
The parameter $\eta=\sigma_{\rm u}\gamma_{\rm u}$ equals the plasma magnetization at the light cylinder,
$\eta=\sigma_{\rm LC}\sim\frac{\mu^{2}}{4\pi R_{\rm LC}^{3}{\cal N}mc^{2}}\approx 10^{4}\,\frac{\mu_{33}^{2}}{R_{\rm LC,10}^{3}{\cal N}_{37}},$ (86)
where ${\cal N}\equiv nr^{3}$ is the density parameter of the magnetosphere; its typical value is ${\cal N}\sim 10^{37}$ (Beloborodov, 2020). The parameter $\eta$ and the Lorentz factor $\gamma_{\rm u}(r)$ determine the wind magnetization $\sigma_{\rm u}(r)=\eta/\gamma_{\rm u}$.
Note that $L_{\rm w}$ determines the wind magnetic field in the static lab frame $B_{\rm u}(r)$ according to Equation (83). The corresponding magnetic field in the rest frame of the wind (ahead of the FRB wave packet) is $\tilde{B}=B_{\rm u}/\gamma_{\rm u}$. It sets another important parameter — the gyro-frequency of the wind plasma,
$\tilde{\omega}_{B}^{\rm u}=\frac{e\tilde{B}_{\rm u}}{mc}=\frac{e}{\gamma_{\rm u}mcr}\left(\frac{L_{\rm w}}{c}\right)^{1/2},$ (87)
which should be compared with the FRB frequency. When viewed in the wind rest frame, the FRB frequency $\omega$ is redshifted to $\tilde{\omega}_{\rm u}=\omega/\kappa_{\rm u}$ where $\kappa_{\rm u}\equiv\gamma_{\rm u}(1+\beta_{\rm u})\approx 2\gamma_{\rm u}$. Then, we find
$\alpha\equiv\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}\approx\frac{2e}{mcr\omega}\left(\frac{L_{\rm w}}{c}\right)^{1/2}\approx 0.1\,\frac{L_{\rm w,37}^{1/2}}{\nu_{9}r_{12}}.$ (88)
Below we examine the compression front driven by an FRB emitted at some radius $r$ and propagating outward through the wind. We first describe the front at large $r$, where FRBs induce heating in the regular regime ($\tilde{\omega}_{B}<\tilde{\omega}$), and then consider FRBs emitted at smaller $r$, where heating is stochastic and more dramatic effects occur, including damping of the FRB.
8 Regular heating zone
Strong compression is expected at radii where the FRB has $a>1$. The wave strength parameter $a$ (Equation 1) may be written as $a=r_{1}/r$, where
$r_{1}=\left(\frac{r_{e}L}{mc\,\omega^{2}}\right)^{1/2}\approx 1.6\times 10^{13}\,L_{42}^{1/2}\nu_{9}^{-1}\,{\rm cm}.$ (89)
8.1 Saturated compression ($r>r_{\star}$)
One can quickly formulate a compression front model using the steady-state solution described by Equations (52), (57), and (75), assuming the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}$) and neglecting radiative losses. Using the definition of the proper compression factor $\tilde{C}\equiv\tilde{\rho}/\tilde{\rho}_{\rm u}$, the steady front solution can be stated as
$\tilde{C}_{\star}=\frac{\kappa_{\star}}{\kappa_{\rm u}}=1+\frac{3}{2}\left(\sqrt{a^{2}+1}-1\right).$ (90)
The corresponding compression factor for the wind density measured in the lab frame, $C\equiv\rho/\rho_{\rm u}$, is
$C_{\star}=\frac{\kappa_{\star}^{2}+1}{\kappa_{\rm u}^{2}+1}\approx\tilde{C}_{\star}^{2},$ (91)
where $\kappa\equiv\gamma_{\rm D}(1+\beta_{\rm D})$ describes the fluid drift speed $\beta_{\rm D}$, and we used $\kappa_{\rm u}\gg 1$ since the upstream medium is a relativistic wind. Outside the radius $r_{1}$ (where $a\ll 1$), one finds weak compression $C_{\star}\approx 1$ while inside $r_{1}$ (where $a\gg 1$) one finds $C_{\star}\approx(9/4)a^{2}=(9/8)a_{0}^{2}\approx a_{0}^{2}$. A simple approximate expression matching the two cases is
$C_{\star}\approx 1+a_{0}^{2}.$ (92)
Remarkably, this compression is consistent with the results of Sobacchi et al. (2024), who investigated a steady plasma flow in a (nearly monochromatic) wave packet using a different technique.
However, the simple solution stated in Equation (92) does not correctly describe compression fronts in FRBs at small radii $r\ll r_{1}$, for two reasons:
(1) At $r<r_{\star}$ defined below, the compression front does not reach the steady state, so the actual compression factor is well below its saturated value, $C\ll C_{\star}$.
(2) Compression promotes the transition to stochastic acceleration of plasma particles in the FRB wave packet. The plasma then experiences much stronger heating, and the momentum exchange between the FRB wave packet and the plasma is enhanced, changing $\kappa$ and $C$.
These two features qualitatively change the picture of the FRB compression front, as described below.
8.2 Unsaturated compression $(r<r_{\star})$
The saturated (steady-state) compression factor $C_{\star}$ may be reached only when the wave packet has propagated for a time $t>t_{\star}$. In the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}_{B}$), the timescale $t_{\star}$ is given by Equation (74) in the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$). The corresponding timescale measured in any other frame (any $\kappa_{\rm u}$) is $t_{\star}\sim\sigma_{\rm u}a^{2}\gamma_{\rm u}^{2}T$ (assuming $a^{2}>1$). So, the relaxation timescale in our static lab frame is
$t_{\star}\sim\eta\gamma_{\rm u}a^{2}T.$ (93)
[…]
$\frac{t}{t_{\star}}=\frac{r^{3}}{r_{\star}^{3}},$ (94)
$r_{\star}\approx\left(\frac{r_{e}{\cal E}\eta\gamma_{\rm u}}{m\omega^{2}}\right)^{1/3}\approx 10^{13}\,\frac{{\cal E}_{39}^{1/3}\eta_{4}^{4/9}}{\nu_{9}^{1/3}}\,{\rm cm},$ (95)
where ${\cal E}\sim LT$ is the FRB energy. This expression for $r_{\star}$ assumes $a^{2}(r_{\star})>1$, i.e. it holds as long as $r_{\star}<r_{1}$. At the transition radius $r=r_{\star}$ we find
$C(r_{\star})\approx C_{\star}(r_{\star})\approx\left[1+\frac{3}{2}\left(\sqrt{\frac{r_{1}^{2}}{r_{\star}^{2}}+1}-1\right)\right]^{2}.$ (96)
When $t_{\star}>t=r/c$, the propagating wave packet develops compression factors $C<C_{\star}$ (Figure 4). At times $\gamma_{\rm u}^{2}T\ll t\ll t_{\star}$, $C(t)$ may be approximated by
$C-1\approx(C_{\star}-1)\left(\frac{t}{t_{\star}}\right)^{\zeta}\approx\frac{9}{4}a^{2}\left(\frac{r}{\eta\gamma_{\rm u}a^{2}cT}\right)^{\zeta}.$ (97)
The numerical results in Figure 4 suggest $\zeta\approx 1/2$.
9 Stochastic heating zone
The zone of stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$) is larger than one might expect from the ratio $\tilde{\omega}_{B}^{\rm u}/\tilde{\omega}_{\rm u}$ (Equation 88). Even when $\tilde{\omega}_{B}<\tilde{\omega}$ in the upstream plasma, the condition $\tilde{\omega}_{B}>\tilde{\omega}$ can be met inside the wave packet as a result of the fluid acceleration and compression: acceleration reduces $\tilde{\omega}$ and compression increases $\tilde{\omega}_{B}$.
Fluid acceleration ($\kappa/\kappa_{\rm u}$) and compression ($\tilde{C}$ or $C$) are related by
$\tilde{C}\equiv\frac{\tilde{\rho}}{\tilde{\rho}_{\rm u}}\approx\frac{\kappa}{\kappa_{\rm u}},\qquad C\equiv\frac{\rho}{\rho_{\rm u}}\approx\frac{\kappa^{2}+1}{\kappa_{\rm u}^{2}+1}.$ (98)
These relations are exact for a steady compression front (with $C=C_{\star}$ and $\kappa=\kappa_{\star}$) and also approximately hold for unsaturated compression ($C\ll C_{\star}$ and $\kappa\ll\kappa_{\star}$), as demonstrated in Figure 5 using the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$). Equation (98) states the $\kappa$-$C$ relation viewed in any reference frame, where $\kappa_{\rm u}$ can have any value. The magnetic field is frozen in the fluid and compressed by the same factor as density:
$\frac{\tilde{B}}{\tilde{B}_{\rm u}}=\tilde{C},\qquad\frac{B}{B_{\rm u}}=C,$ (99)
where $\tilde{B}$ is measured in the local fluid rest frame, and subscript “u” indicates the upstream region ahead of the FRB. Equation (99) implies
$\tilde{\omega}_{B}\approx\frac{\kappa}{\kappa_{\rm u}}\,\tilde{\omega}_{B}^{\rm u}.$ (100)
The FRB frequency $\omega$ transforms to the local fluid frame as $\tilde{\omega}=\omega/\kappa$. Hence,
$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}\approx\frac{\kappa^{2}}{\kappa_{\rm u}^{2}}\,\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}.$ (101)
In the lab frame, the upstream medium (the wind) moves with $\kappa_{\rm u}\approx 2\gamma_{\rm u}\gg 1$, and Equation (101) may be re-written as
$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}\approx C\,\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}=C\alpha.$ (102)
Now using Equation (88) one can see that $\tilde{\omega}_{B}>\tilde{\omega}$ when
$C>C_{\rm stoch}\approx\frac{\tilde{\omega}_{\rm u}}{\tilde{\omega}_{B}^{\rm u}}\approx 10\,\frac{r_{12}\nu_{9}}{L_{\rm w,37}}.$ (103)
The stochastic regime $\tilde{\omega}_{B}>\tilde{\omega}$ appears at $r<r_{\rm stoch}$, with the transition radius $r_{\rm stoch}$ typically inside $r_{\star}$. Its value is found by equating $C$ given by Equation (97) to $C_{\rm stoch}$. In particular, for $\zeta\approx 1/2$ we find
$r_{\rm stoch}\approx\left(\frac{4r_{e}^{2}L_{\rm w}L}{m^{2}c^{3}\omega^{4}\eta\gamma_{\rm u}T}\right)^{1/3}.$ (104)
This gives $r_{\rm stoch}\sim 10^{12}$ cm for typical parameters. When $r_{\rm stoch}>\eta^{1/3}R_{\rm LC}$, one can substitute $\gamma_{\rm u}\approx\eta^{1/3}$ and then Equation (104) implies $r_{\rm stoch}\propto L^{1/3}$. If one keeps $\zeta$ as a parameter that may differ from 1/2, one still finds $r_{\rm stoch}\propto L^{1/3}$:
$r_{\rm stoch}\approx 7\times 10^{11}\,H\,L_{42}^{1/3}\,{\rm cm},$ (105)
$H\approx\left(\frac{L_{\rm w,37}^{1/2}}{T_{\rm ms}^{\zeta}\eta_{4}^{4\zeta/3}\nu_{9}^{3-2\zeta}}\right)^{1/(3-3\zeta)}.$ (106)
The expression for $H$ omits a $\zeta$-dependent prefactor, which deviates from unity when $\zeta$ deviates from 1/2.
9.1 Heating and radiative cooling in the wave packet
At radii $r<r_{\rm stoch}$, the FRB stochastically heats plasma particles to $\tilde{\gamma}\gg a$, far exceeding regular heating. In this regime, $\tilde{\gamma}$ monotonically grows in the fluid element as it crosses the wave packet (Figure 2). The growth is well described by Equation (18) with the numerical coefficient $\chi\approx 2/3$. It was derived in the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), and a similar derivation can be repeated in any frame with $\kappa_{\rm u}\neq 1$ (the only change occurs in the expression for the gyro-frequency $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$ (Equation 100)). Thus, we find
$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}\approx\frac{16\pi}{3}\frac{r_{e}U\kappa^{2/3}}{m\omega^{4/3}}\left(\frac{\tilde{\omega}_{B}^{\rm u}}{\kappa_{\rm u}}\right)^{1/3},$ (107)
where we substituted $a^{2}=4\pi r_{e}U/m\omega^{2}$. Integration for $\tilde{\gamma}(\xi)$ then gives
$\tilde{\gamma}_{\rm stoch}\approx\left[\frac{3r_{e}{\cal E}}{mc\omega r^{2}}\right]^{3/7}\!\!\left(C\alpha\right)^{1/7}\approx 10^{5}\frac{{\cal E}_{39}^{3/7}C^{1/7}L_{\rm w,37}^{1/14}}{\nu_{9}^{4/7}r_{11}},$ (108)
[…]
Equation (108) holds as long as radiative losses of the plasma are negligible. The losses $\dot{\gamma}_{e}$ are given by Equation (36) (with the coefficient 2/3 replaced by 4/3 for stochastically heated particles, see B22). Losses $\dot{\gamma}_{e}$ offset the heating rate (Equation 10) when $\tilde{\gamma}$ approaches $\tilde{\gamma}_{\rm rad}=(c/r_{e})^{3/10}\tilde{\omega}_{B}^{1/10}\tilde{\omega}^{-2/5}$. Using $\tilde{\omega}=\omega/\kappa=\tilde{\omega}_{\rm u}\kappa_{\rm u}/\kappa$ and $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$, we find
$\tilde{\gamma}_{\rm rad}\approx\left(\frac{c}{r_{e}\omega}\right)^{0.3}\!\alpha^{0.1}\,\frac{\kappa^{0.5}}{\kappa_{\rm u}^{0.2}}\approx 10^{4}\,\frac{C^{1/4}\kappa_{\rm u}^{0.3}L_{\rm w,37}^{1/20}}{\nu_{9}^{0.4}r_{11}^{0.1}}.$ (109)
In summary, stochastic heating gives
$\tilde{\gamma}=\min\{\tilde{\gamma}_{\rm stoch},\tilde{\gamma}_{\rm rad}\}.$ (110)
The radiative regime $\tilde{\gamma}=\tilde{\gamma}_{\rm rad}$ occurs at radii that satisfy
$r<\frac{6\times 10^{11}}{\gamma_{\rm u}^{1/3}}\,\frac{{\cal E}_{39}^{10/21}L_{\rm w,37}^{1/42}}{C_{2}^{5/42}\,\nu_{9}^{4/21}}\,{\rm cm}.$ (111)
9.2 Plasma compression and trapping by the FRB
Once the radio wave begins to propagate through the plasma, it immediately induces significant compression $C$ (section 6.1), which will be denoted below as $C_{0}$. When propagation occurs in the zone of stochastic heating $r<r_{\rm stoch}$, the immediate compression $C_{0}$ turns out so large that further growth of $C>C_{0}$ on timescales $t\gg t_{\rm cross}$ (front relaxation toward the steady state) is irrelevant, because $t_{\rm cross}$ becomes too long. In general, large $C$ corresponds to large $\gamma_{\rm D}\approx C^{1/2}\gamma_{\rm u}$ and implies a long timescale for the plasma to cross the wave packet $t_{\rm cross}\sim\gamma_{\rm D}^{2}T$ (Equation 69). When $t_{\rm cross}$ reaches $t=r/c$, the plasma becomes trapped in the packet, surfing the radio wave. The ratio $t/t_{\rm cross}$ may be written as
$\frac{t_{\rm cross}}{t}\sim\frac{cT\gamma_{\rm D}^{2}}{r}=\frac{\gamma_{\rm D}^{2}}{\gamma_{\rm trap}^{2}}=\frac{C}{C_{\rm trap}},$ (112)
where
$\gamma_{\rm trap}\approx\left(\frac{r}{cT}\right)^{1/2},\qquad C_{\rm trap}\approx\frac{r}{cT\gamma_{\rm u}^{2}}.$ (113)
Let us now evaluate the compression factor $C_{0}$ that immediately develops in the FRB-wind interaction, and show that it easily reaches $C_{\rm trap}$. The calculation was set up in section 6.1 using the frame where the upstream plasma is at rest. Here, we will denote this frame ${\cal K}^{\prime}$ to distinguish it from our lab frame; quantities measured in frame ${\cal K}^{\prime}$ will be denoted with a prime. We are interested in the strong compression regime $\tilde{C}_{0}\gg 1$. Then, Equation (66) gives
$\tilde{C}_{0}\approx\kappa^{\prime}\approx\frac{3\tilde{\gamma}}{2\sigma_{\rm u}}+\sqrt{\frac{\epsilon_{\rm rad}^{\prime}}{\sigma_{\rm u}}},$ (114)
where we used $w\approx(3/2)\tilde{\gamma}$.
A proper compression factor $\tilde{C}$ corresponds to $C\approx\tilde{C}^{2}$ in the lab frame (Equation 98), and radiative losses in the lab frame are ${\cal E}_{\rm rad}\approx\kappa_{\rm u}{\cal E}_{\rm rad}^{\prime}$. So, one can use two general relations
$\kappa^{\prime}\approx C^{1/2},\qquad\epsilon_{\rm rad}^{\prime}\approx\frac{\epsilon_{\rm rad}}{\kappa_{\rm u}}.$ (115)
Radiative losses of a particle interacting with the radio wave packet are given by Equation (38) with an extra factor of 2 in the stochastic heating regime (B22). Thus, we find
[…]
where we used $dt=d\xi/(1-\beta_{\rm D})$ and $\kappa^{2}(1-\beta_{\rm D})=1+\beta_{\rm D}$. Note that this expression involves the FRB energy ${\cal E}$, assuming that the particle interacts with the entire radio packet, i.e. $t_{\rm cross}\leq r/c$, which corresponds to $C\leq C_{\rm trap}$. [^8] A “deep” trapping regime $t_{\rm cross}\gg r/c$ is possible if the plasma is trapped in a small leading part of the wave packet, at $\xi_{\rm trap}\ll T$. In this case, ${\cal E}$ in Equation (116) should be replaced by the part of the FRB energy ${\cal E}_{\rm trap}$ contained in the leading part $0<\xi<\xi_{\rm trap}$. Self-consistent $\xi_{\rm trap}$ and ${\cal E}_{\rm trap}$ depend on the packet shape.
Using Equations (115) and (116) in Equation (114), we find
$C_{0}^{1/2}\sim\frac{\tilde{\gamma}}{\sigma_{\rm u}}\left(\frac{3}{2}+\psi\right),\qquad\psi\equiv\sqrt{\frac{\sigma_{\rm T}{\cal E}\sigma_{\rm u}}{4\pi r^{2}mc^{2}\kappa_{\rm u}}}.$ (117)
Here, $\tilde{\gamma}=\min\{\tilde{\gamma}_{\rm stoch},\tilde{\gamma}_{\rm rad}\}$ depends on $C=C_{0}$ (section 9.1), and one can find the consistent solution for $\tilde{\gamma}$ and $C_{0}$:
$\tilde{\gamma}=\left\{\begin{array}[]{lr}\tilde{\gamma}_{\rm stoch}&\quad\psi<0.7\\ \tilde{\gamma}_{\rm rad}&\quad\psi>0.7\end{array}\right.$ (118)
$C_{0}\sim\left[\frac{c\kappa_{\rm u}}{r_{e}\omega}\right]^{6/5}\!\!\frac{\alpha^{2/5}}{16\sigma_{\rm u}^{4}}\!\times\!\left\{\begin{array}[]{lr}(3\psi)^{12/5}(3+2\psi)^{14/5}&\psi<0.7\\ (3+2\psi)^{4}&\psi>0.7\end{array}\right.$ (119)
with the parameter $\alpha$ given by Equation (88).
The obtained solution holds if it gives $C_{0}\gg 1$, and this condition is satisfied at all $r<r_{\rm stoch}$. Furthermore, $C_{0}$ exceeds $C_{\rm trap}$. Indeed, consider the smallest $C_{0}^{1/2}\sim 3\tilde{\gamma}_{\rm stoch}/2\sigma_{\rm u}$ (Equation 117), which is found in the non-radiative regime $\psi\ll 0.7$ (small ${\cal E}$). Then, using Equation (110) for $\tilde{\gamma}_{\rm stoch}$, we find
$C_{0}\sim\frac{10^{6}\,{\cal E}_{39}^{6/5}L_{\rm w,37}^{1/5}}{\nu_{9}^{8/5}\eta_{4}^{14/5}R_{\rm LC,10}^{14/5}}\!\!\times\!\left\{\begin{array}[]{lr}1&r<R_{\rm sat}\\ (R_{\rm sat}/r)^{14/5}&r>R_{\rm sat}\end{array}\right.$ (120)
where $R_{\rm sat}=\eta^{1/3}R_{\rm LC}\approx 2\times 10^{11}\,\eta_{4}^{1/3}R_{\rm LC,10}$ cm is the radius where the wind Lorentz factor $\gamma_{\rm u}$ approaches $\eta^{1/3}$ (Equation 85). It is easy to verify that $C_{0}>C_{\rm trap}$ at all $r<r_{\rm stoch}$. So, FRBs emitted in the wind at $r<r_{\rm stoch}$ immediately trap the wind plasma.
Two aspects of the trapping regime should be noted:
(1) The large value of $C_{0}$ is found using the full FRB energy ${\cal E}\sim LT$ in the parameter $\psi$. When $C_{0}\gg C_{\rm trap}$, the plasma becomes trapped in a small leading part of the radio wave packet with ${\cal E}_{\rm trap}\ll{\cal E}$ (see footnote 8), avoiding interaction with its main part.
(2) The trapping regime holds unless the FRB becomes strongly damped by its interaction with the wind. Damping allows the plasma to move through the wave packet while consuming its energy, avoiding trapping. The parameter space where FRBs are damped is evaluated below.
10 FRB energy losses
Consider first FRB energy losses $\Delta{\cal E}^{\prime}$ in frame ${\cal K}^{\prime}$ where the upstream plasma is at rest. The corresponding momentum losses are $\Delta{\cal E}^{\prime}/c$.
Let the radio wave packet interact with $\Delta{\cal N}$ particles. The momentum lost by the radio wave is the sum of momentum gained by the MHD flow and momentum waisted by the energized plasma particles to radiation,
$\frac{\Delta{\cal E}^{\prime}}{c}=\Delta{\cal N}\,\frac{T^{t^{\prime}r^{\prime}}_{\rm MHD}}{cn^{\prime}}+\beta_{\rm D}^{\prime}\frac{{\cal E}_{\rm rad}^{\prime}}{c}.$ (121)
The radiated energy per particle is $\epsilon_{\rm rad}^{\prime}mc^{2}\equiv{\cal E}_{\rm rad}^{\prime}/\Delta{\cal N}$. This gives
$\frac{\Delta{\cal E}^{\prime}}{mc^{2}\Delta{\cal N}}\approx\beta_{\rm D}^{\prime}\left[\gamma_{\rm D}^{\prime}(1+w+\sigma_{\rm u}\kappa^{\prime})+\epsilon_{\rm rad}^{\prime}\right].$ (122)
Note that $\beta_{\rm D}^{\prime}=(\kappa^{\prime 2}-1)/(\kappa^{\prime 2}+1)$ and $\kappa^{\prime}\approx\tilde{C}$, where $\tilde{C}=\tilde{\rho}/\tilde{\rho}_{\rm u}$ is the proper compression factor. Thus, Equation (122) yields
$\frac{\Delta{\cal E}^{\prime}}{mc^{2}\Delta{\cal N}}\approx(\tilde{C}^{2}-1)\left(\frac{1+w}{2\tilde{C}}+\frac{\sigma_{\rm u}}{2}+\frac{\epsilon_{\rm rad}^{\prime}}{\tilde{C}^{2}+1}\right).$ (123)
Proper compression $\tilde{C}$ is related to compression $C$ in the lab frame by $C\approx\tilde{C}^{2}$ (Equation 98). Wave energy in the lab frame is ${\cal E}=\kappa_{\rm u}{\cal E}^{\prime}$, so $\Delta{\cal E}^{\prime}=\Delta{\cal E}/\kappa_{\rm u}$. Radiative losses in the lab frame and frame ${\cal K}^{\prime}$ are also related by ${\cal E}_{\rm rad}/{\cal E}_{\rm rad}^{\prime}\approx\kappa_{\rm u}$. Using these relations, one can rewrite Equation (123) in terms of the lab-frame quantities:
$\frac{\Delta{\cal E}}{mc^{2}\Delta{\cal N}\kappa_{\rm u}}\approx(C-1)\left[\frac{1+w}{2\sqrt{C}}+\frac{\sigma_{\rm u}}{2}+\frac{\epsilon_{\rm rad}}{(C+1)\kappa_{\rm u}}\right].$ (124)
As the FRB propagates to a radius $r$, it interacts with $\Delta{\cal N}$ wind particles given by
$\Delta{\cal N}\approx(1-\beta_{\rm u})\frac{r}{c}\dot{\cal N}\approx\frac{2rL_{\rm w}}{mc^{3}\eta\kappa_{\rm u}^{2}}.$ (125)
Equations (124) and (125) determine the FRB energy losses $\Delta{\cal E}$ anywhere in the wind, expressed as a function of the compression factor $C$. Enthalpy $w\approx(3/2)\tilde{\gamma}$ is controlled by the heating regime: $\tilde{\gamma}=(1+a^{2})^{1/2}$ at $r>r_{\rm stoch}$, and $\tilde{\gamma}$ is given by Equation (110) at $r<r_{\rm stoch}$.
It is easy to verify that losses are negligible, $\Delta{\cal E}\ll{\cal E}$, in the zone of regular heating $r>r_{\rm stoch}$. Thus, FRB damping in the wind may only occur in the zone of stochastic heating, $r<r_{\rm stoch}$.
Using $w\approx(3/2)\tilde{\gamma}\gg 1$ and $C\gg 1$, we find from Equation (124):
$\Delta{\cal E}\approx\frac{rL_{\rm w}}{2c\,\eta\gamma_{\rm u}}\left[\frac{3}{2}\tilde{\gamma}\sqrt{C}+\sigma_{\rm u}C+\frac{\sigma_{\rm T}\tilde{\gamma}^{2}{\cal E}}{4\pi r^{2}mc^{2}\gamma_{\rm u}}\right].$ (126)
For an FRB with luminosity $L$ and energy ${\cal E}\sim LT$, we find $\tilde{\gamma}(r,C)$ from Equation (110) at radii $r<r_{\rm stoch}$, and then find $\Delta{\cal E}(r,C)$ from Equation (126).
The FRB is strongly damped if $\Delta{\cal E}$ approaches ${\cal E}$. Damping begins in the leading part of the FRB $\xi\ll T$, as the plasma entering the radio wave packet consumes its local energy ${\cal E}(\xi)\sim 4\pi r^{2}U(\xi)\,c\,\xi$ and makes the deep trapping regime $C\gg C_{\rm trap}$ unsustainable. Damping succeeds in killing the FRB when it enables the plasma to keep going in $\xi$ and crossing the entire wave packet, which corresponds to $\Delta{\cal E}\sim{\cal E}$ and marginal trapping, $C\approx C_{\rm trap}$. Therefore, the condition for marginal damping is
$\frac{\Delta{\cal E}(r,C_{\rm trap})}{{\cal E}}\approx 1.$ (127)
It defines the boundary of the parameter space where FRBs are strongly damped. One can show that $\Delta{\cal E}(r,C_{\rm trap})/{\cal E}$ is a decreasing function of ${\cal E}$, so strong damping occurs for FRBs with energies ${\cal E}$ below a critical value ${\cal E}_{\rm esc}(r)$ defined by Equation (127) and the corresponding critical luminosity is $L_{\rm esc}\sim{\cal E}_{\rm esc}/T$. The damping region on the $r$-$L$ plane is shown in Figure 8.
Figure 8: Regimes of FRB propagation through the wind ($r>R_{\rm LC}$), shown on the $r$-$L$ plane, for FRBs with emission frequency $\nu=1$ GHz and duration $T=1$ ms. The two plots assume the wind energy parameter $\eta=10^{4}$ (left) and $\eta=10^{3}$ (right). The magenta region $r<R_{\rm LC}$ indicates the closed magnetosphere. FRBs emitted inside the magnetosphere experience strong radiative damping (Beloborodov 2024). FRBs emitted at radii $R_{\rm LC}<r<r_{\rm stoch}$ induce fast stochastic heating and compression of the wind. In the light purple region, FRBs trap the wind plasma, carrying it in a leading part of the wave packet and avoiding significant energy losses. In the deep purple region, FRBs are damped: their energy is consumed by heating and compressing the wind plasma, with radiative losses. The colored dots indicate the dominant damping channel: radiative losses ${\cal E}_{\rm rad}$ (red dots), energy kept by the plasma ${\cal E}_{\rm p}$ (cyan dots) and energy spent to compress the wind magnetic field ${\cal E}_{\rm mag}$ (yellow dots). In the olive zone, $r_{\rm stoch}<r<r_{1}$, the wind is still heated to a relativistic enthalpy $w\gg 1$; however, here heating occurs in the regular regime (particles oscillate in the radio wave instead of stochastically gaining energy) with the moderate $w\approx(3/2)a$ and a weaker compression. At radii $r_{\star}<r<r_{1}$, the compression factor is $C=C_{\star}\approx 1+a_{0}^{2}$. The FRB strength parameter $a=a_{0}/\sqrt{2}$ drops below unity outside $r_{1}$ (grey region).
It is useful to examine more closely the loss $\Delta{\cal E}$ stated in Equation (126). The three terms in the square brackets correspond to three contributions,
$\Delta{\cal E}={\cal E}_{\rm p}+{\cal E}_{\rm mag}+{\cal E}_{\rm rad},$ (128)
where ${\cal E}_{\rm p}\propto\tilde{\gamma}$ is the energy kept by the plasma particles, ${\cal E}_{\rm mag}\propto\sigma_{\rm u}$ is the energy used to compress the wind magnetic field, and ${\cal E}_{\rm rad}\propto\tilde{\gamma}^{2}$ is the radiative losses. Let us compare radiative losses ${\cal E}_{\rm rad}$ with ${\cal E}_{\rm p}$ and ${\cal E}_{\rm mag}$.
It is easy to see that ${\cal E}_{\rm rad}>{\cal E}_{\rm p}$ if $\tilde{\gamma}_{\rm rad}<\tilde{\gamma}_{\rm stoch}$; in this regime, the plasma radiates more energy than it keeps. This occurs at radii that satisfy the condition given in Equation (111). With $C=C_{\rm trap}$, it can be stated as
${\cal E}_{\rm rad}>{\cal E}_{\rm p}:\quad r<2\times 10^{11}\,\frac{{\cal E}_{39}^{20/47}T_{\rm ms}^{5/47}L_{\rm w,37}^{1/47}}{\gamma_{\rm u}^{4/47}\nu_{9}^{8/47}}\,{\rm cm}.$ (129)
The ratio ${\cal E}_{\rm rad}/{\cal E}_{\rm mag}$ equals the ratio of the third and second terms in the brackets in Equation (126). Using $C=C_{\rm trap}$ and $\tilde{\gamma}=\tilde{\gamma}_{\rm rad}$ (i.e. assuming ${\cal E}_{\rm rad}>{\cal E}_{\rm p}$), we find
${\cal E}_{\rm rad}>{\cal E}_{\rm mag}:\quad r<10^{11}\gamma_{\rm u}^{16/27}\frac{{\cal E}_{39}^{10/27}T_{\rm ms}^{5/27}L_{\rm w,37}^{1/27}}{\eta_{4}^{10/27}\nu_{9}^{8/27}}\,{\rm cm}.$ (130)
${\cal E}_{\rm rad}$ dominates the FRB losses at radii that satisfy both conditions stated in Equations (129) and (130).
11 Summary
The basic physics problem studied in this paper is how a strong ($a_{0}\gg 1$) electromagnetic wave packet accelerates an MHD fluid. For weak electromagnetic waves ($a_{0}\ll 1$), the acceleration effect is well known as the radiation pressure force, used in astrophysics e.g. to define the Eddington limit. For strong waves, the acceleration effect was previously calculated in unmagnetized plasmas; the $e^{\pm}$ plasma then becomes compressed in the wave packet by the factor $C\approx 1+a_{0}^{2}$. A similar problem for a magnetized plasma was poorly understood. The present paper developed a method to solve this problem, obtained sample numerical solutions, and provided a physical picture of the wave-plasma interaction. Two key facts help solve the problem: (1) the wave strength parameter $a_{0}$ controls the plasma enthalpy as shown in section 3, and (2) quantity ${\cal Q}_{-}={\cal E}-c{\cal P}$ (where ${\cal E}$ is energy and ${\cal P}$ is momentum) is conserved in the wave-plasma interaction because the wave has ${\cal Q}_{-}=0$ and cannot transfer $\Delta{\cal Q}_{-}\neq 0$ to the MHD fluid.
A first estimate would suggest that plasma with a sufficiently large magnetization $\sigma$ is not strongly compressed by the electromagnetic wave, since the deposited momentum is proportional to the small plasma density while compression requires work to be done against the large background magnetic pressure (Beloborodov 2021; see also section 6.1). The actual solution demonstrates that this expectation holds initially, when the wave packet begins to interact with the plasma. Later, the propagating wave gradually develops a large compression factor $C$ due to a curious “memory” effect described in section 6. Given a sufficient time $t_{\star}\propto C\sigma$ the compression front relaxes to a steady state with $C=C_{\star}\approx 1+a_{0}^{2}$ independent of $\sigma$. However, large $C_{\star}$ (implied by large $a_{0}$) is not reached if $t_{\star}$ exceeds the wave propagation time.
Besides the peculiar gradual evolution of the propagating wave packet, the background magnetic field brings another special feature that dramatically changes the wave-plasma interaction: compression increases the gyro-frequency and triggers strong stochastic heating of the plasma by the electromagnetic wave. The heating boosts momentum deposition, enhancing acceleration and compression of the MHD fluid in the wave packet.
We have used the obtained results to examine FRB interaction with a magnetized wind flowing from the central object (a neutron star). Our conclusions are summarized below and illustrated in Figure 8.
FRBs propagating at radii $R_{\rm LC}<r<r_{\rm stoch}$ have a huge effect on the wind medium. They heat the wind up to an average Lorentz factor $\tilde{\gamma}\sim 10^{3}a_{0}$ (measured in the fluid rest frame), and accelerate it to a bulk Lorentz factor exceeding $\gamma_{\rm trap}\sim(r/cT)^{1/2}$, which implies trapping of the wind plasma inside the radio wave packet. The energy per particle in the heated plasma $\gamma_{\rm trap}\tilde{\gamma}mc^{2}$ (measured in the static lab frame) is in the TeV range. Thus, the FRB-wind interaction converts the radio-wave energy into high energy particles, which experience radiative losses. A fraction of FRB energy is also spent to compress the wind magnetic field. Only sufficiently bright FRBs survive the energy losses (Figure 8).
The FRB-wind interaction changes at larger radii: at $r=r_{\rm stoch}$ (Equation 104) there is a sharp transition from stochastic heating to regular heating (particle oscillation in the wave), which sustains a moderate $\tilde{\gamma}\sim a_{0}$. Energy losses of the radio wave are negligible at $r>r_{\rm stoch}$. FRBs emitted in this zone still induce strong bulk acceleration of the wind plasma. Here, instead of trapping the plasma, the FRB drives an MHD compression wave in the wind. At $r>r_{\star}$ (Equation 95), the compression wave is quasisteady ($r/c\gg t_{\star}$), and the compression factor equals $C_{\star}\approx 1+a_{0}^{2}$.
[…]
The results summarized in Figure 8 imply interesting constraints on FRB emission models. In particular, they exclude scenarios that place the radio source at radii $r<r_{\rm damp}\sim 10^{11}$ cm and rely on free FRB propagation through the magnetar wind. Thus, the previously found damping zone $r<R_{\rm LC}$ (Beloborodov, 2021, 2024; Golbraikh & Lyubarsky, 2023) is extended to $r_{\rm damp}\sim 10^{11}$ cm. The exact value of $r_{\rm damp}$ varies around $10^{11}$ cm for the relevant values of wind parameter $\eta=L_{\rm w}/\dot{\cal N}mc^{2}$ around $10^{4}$. We find that $r_{\rm damp}$ slowly decreases with increasing $\eta>10^{4}$. The compression and damping effects are particularly strong in winds with $\eta<10^{4}$.
[…]
In addition to plasma acceleration and compression studied in this paper, the ambient medium may be strongly affected by the nonlinear development of self-modulation instabilities of the radio wave (Sobacchi et al., 2023), investigated so far for waves with $a_{0}<1$. Instabilities in the wind zone with $a_{0}>1$ are poorly understood and will require future work.
The author acknowledges support by NASA grants 21-ATP21-0056 and 80NSSC24K0282, NSF grant AST-2408199, and Simons Foundation grant No. 446228.

### v2

@@
 The likely central engines of FRBs are magnetars, and their ambient plasma at radii $r\gg 10^{10}$ cm is a magnetized $e^{\pm}$ wind.
-We formulate equations describing the FRB-plasma interaction, solve them numerically, and provide an approximate analytical picture of the interaction.
+We formulate basic equations of the FRB-plasma interaction, solve them numerically, and describe the physical picture of the interaction.
 We find the following: (1) FRBs emitted at $r<r_{\rm stoch}\sim 10^{12}$ cm induce fast stochastic heating and strong compression of the wind, sweeping it like a broom.
-The outcome of this interaction is determined by the energy losses of the radio wave.
-We evaluate the parameter space where the FRB survives its interaction with the wind.
-(2) At radii $r>r_{\rm stoch}$, the FRB induces regular heating to the Lorentz factor $\sim a_{0}$, where $a_{0}\propto r^{-1}$ is the wave strength parameter.
-At $r>r_{\star}\sim 10^{13}\,$cm, the FRB drives a quasisteady compression wave in the wind, with compression factor $C_{\star}\approx 1+a_{0}^{2}$.
-Both characteristic radii $r_{\rm stoch}$ and $r_{\star}$ scale with FRB luminosity $L$ as $L^{1/3}$.
+The outcome of this interaction is determined by the energy losses of the radio wave; we evaluate the parameter space where FRBs survive and escape.
+(2) At radii $r>r_{\rm stoch}$, FRB induces regular particle oscillations in the radio wave with the standard strength parameter $a$, and drives a compression wave in the wind.
+At $r>r_{\star}\sim 10^{13}\,$cm, the compression wave becomes locally quasisteady, with compression factor $1+a^{2}$.
 FRBs avoid damping if they are released into the wind medium outside $r_{\rm damp}\sim 10^{11}$ cm.
 Magnetars (992); Plasma astrophysics (1261); Radio bursts (1339); X-ray transient sources (1852); Magnetohydrodynamics (1964); Radiative processes (2055)
-1 Introduction
+I Introduction
 Observations of fast radio bursts (FRBs) demonstrate the existence of ultrastrong electromagnetic waves in the Universe.
-They have characteristic frequencies $\nu=\omega/2\pi\sim 1$ GHz and durations $T\sim 1$ ms (Petroff et al., 2019), and their dimensionless strength parameter is [^1] Throughout the paper we use notation $Q_{n}$ for a quantity $Q$ normalized to $10^{n}$ in CGS units, e.g.
-$\nu_{9}=\nu/10^{9}{\rm Hz}$.
+They have characteristic frequencies $\nu=\omega/2\pi\sim 1$ GHz and durations $T\sim 1$ ms (Petroff et al., 2019), and their dimensionless strength parameter is [^1] We use the standard notation $Q_{n}$ for a quantity $Q$ normalized to $10^{n}$ in CGS units, e.g.
+$\nu_{9}=\nu/10^{9}{\rm Hz}$ and $r_{12}=r/10^{12}{\rm cm}$.
 $a\equiv\frac{e\langle E_{\rm wave}^{2}\rangle^{1/2}}{mc\,\omega}\approx 16\,\nu_{9}^{-1}L_{42}^{1/2}r_{12}^{-1},$ (1)
-where $L=cr^{2}\langle E_{\rm wave}^{2}\rangle$ is the isotropic equivalent of wave luminosity; $\langle E_{\rm wave}^{2}\rangle^{1/2}$ is the root-mean-square of the wave electric field when the wave expands to a radius $r$.
+where $m$ is the electron mass, $c$ is the speed of light, $L=cr^{2}\langle E_{\rm wave}^{2}\rangle$ is the isotropic equivalent of wave luminosity; $\langle E_{\rm wave}^{2}\rangle^{1/2}$ is the root-mean-square of the wave electric field when the wave expands to a radius $r$.
 FRBs have to propagate through the ambient plasma surrounding the source.
 The likely central engines of FRBs are neutron stars, which are surrounded by electron-positron plasma.
-The plasma is strongly magnetized:
+The plasma is strongly magnetized: its magnetic stress $B^{2}/4\pi$ far exceeds the rest-mass energy density $\rho c^{2}$,
 $\sigma\equiv\frac{B^{2}}{4\pi\rho c^{2}}\gg 1,$ (2)
-where the magnetic field $B$ and density $\rho$ are measured in the plasma rest frame.
+where $B$ and $\rho$ are measured in the plasma rest frame.
 Near the neutron star, the plasma is locked in the closed magnetosphere, rotating with a period $P\sim 1-10$ s.
@@
 Strong radiative damping has been demonstrated for FRBs emitted near a magnetar and propagating through its equilibrium outer magnetosphere at radii $r\sim 10^{8}-10^{9}$ cm (Beloborodov, 2021, 2024).
+In addition, strong radio waves are damped through excitation of small-scale Alfvén waves in the magnetosphere (Golbraikh & Lyubarsky, 2023).
 […]
-The present paper revisits the FRB-wind interaction. The problem is formulated in section 2 and approached in two steps:
-First, we use test-particle simulations to demonstrate two types of heating, regular and stochastic, of a plasma exposed to (and compressed by) an FRB (section 3).
-When the plasma magnetic field is increased by compression, regular oscillations of particles in the radio wave are changed to stochastic resonant heating.
-This change boosts the heating and makes it irreversible.
-Then, we examine how heating by the FRB results in the bulk acceleration and compression of the plasma.
-Heating determines acceleration because energy $\Delta{\cal E}$ received from the radio wave always comes with momentum $\Delta{\cal E}/c$ (the wave loses momentum $\Delta{\cal E}/c$).
-Plasma is accelerated and compressed on timescales much longer than the Larmor period and the wave oscillation period.
-[^2] In simplest cases, the effect may be thought of as a ponderomotive force exerted by the wave.
-In the presence of strong radiative losses it is more appropriate to view it as a scattering process (B22).
-In any case, it is described from first principles by solving for the particle motion in the wave and finding $\Delta{\cal E}/c$.
-Therefore, it behaves as a fluid governed by magnetohydrodynamics (MHD).
-In section 4, we formulate MHD equations for the plasma bulk motion inside a propagating radio wave packet.
-These equations determine plasma compression by FRBs.
+The present paper investigates the FRB-wind interaction using a new method. The problem is stated in section II and then solved in two steps:
+First, we describe plasma response to an FRB on a “micro” level by tracking the motion of individual particles (section III).
+The response can be regular oscillation or stochastic heating (Beloborodov 2022; hereafter B22).
+It is best viewed in the local center-of-momentum frame of the plasma particles (“fluid frame”).
+Remarkably, the microscopic response can be described independently of the fluid motion.
+This fact is illustrated with simulations of particles driven by FRBs in static and accelerating magnetized fluids.
+The simulations demonstrate the regular and stochastic response, and the transition between the two regimes.
+Then, we examine the macroscopic fluid dynamics of plasma exposed to FRBs.
+The magnetized fluid is pushed and compressed by the FRB on timescales much longer than the particle Larmor period and the wave oscillation period, [^2] The acceleration effect may be thought of as a ponderomotive force exerted by the radio wave.
+In the presence of strong radiative losses it is more suitable to view it as a scattering process.
+In any case, it may be described from first principles by solving for particle motion in the wave and then examining the fluid dynamics on macro scales, as done in this paper.
+so it obeys magnetohydrodynamic (MHD) equations.
+The effective enthalpy of the MHD fluid is determined by particle motions in the fluid frame (i.e.
+by the microscopic response to the FRB) and controlled by the Lorentz-invariant parameter $a$.
+One can think of the micro-level response as a local heating process and the macro-level response as bulk acceleration of the fluid, leading to its compression.
+Equations for fluid dynamics inside a propagating radio wave packet are formulated in section IV.
+It is easy to see why plasma heating by FRBs is inevitably accompanied by its bulk acceleration: any energy $\Delta{\cal E}$ received from the radio wave comes with momentum $\Delta{\cal E}/c$ directed along the wave propagation direction (the wave loses momentum $\Delta{\cal E}/c$).
+Beloborodov (2021) estimated this effect for FRBs propagating through the dipole magnetosphere of a neutron star, where the FRB-plasma interaction is highly radiative and can be described as a scattering process.
+A more general and rigorous calculation should be formulated as a time-dependent fluid dynamics problem, using MHD equations and allowing weak or strong radiative losses of the heated plasma.
+This is attempted in the present paper.
+The weakly radiative regime becomes particularly relevant when the FRB propagates through the wind far outside the closed magnetosphere.
+While heating by FRBs implies push and compression, the compression in turn can change the heating process.
+Note that the MHD fluid is compressed together with its magnetic field.
+Once the magnetic field exceeds a certain value, regular particle oscillations in the radio wave change into stochastic heating, as shown below.
+This change boosts the heating rate and leads to even stronger compression, with enhanced radiative losses of the heated particles.
+The compression fronts driven by FRBs in magnetar winds involve quite dramatic effects, capable of damping the FRB itself as discussed in later sections of this paper.
 Before considering FRBs expanding through a radial wind, we examine the compression front in a plane geometry with a uniform background.
-First, we show that the problem has a simple steady-state solution (section 5).
+First, we show that the problem has a simple steady-state solution (section V), which agrees with the results of Sobacchi et al.
+(2024).
 Its compression factor $C_{\star}$ is directly related to the particle energy gained from the radio wave and becomes huge for waves with strength parameter $a\gg 1$.
 Remarkably, $C_{\star}$ is independent of the plasma magnetization $\sigma$.
-This fact casts some doubt on the relevance of the steady-state solution; in particular, it clearly becomes irrelevant in the vacuum limit ($\sigma\rightarrow\infty$) where the radio wave should have no compression effect on a background magnetic field.
-A complete picture is revealed only by the full time-dependent solution (section 6).
-It shows that the MHD flow pattern in the propagating wave packet is gradually attracted to the steady state, and this relaxation takes a long time $t_{\star}$ proportional to $\sigma$.
-For propagation times $t<t_{\star}$, the steady state with large $C_{\star}$ is not reached; we evaluate the actual compression factor $C\ll C_{\star}$ in this situation.
+This fact casts doubt on the relevance of the steady-state solution; in particular, it clearly becomes irrelevant in the vacuum limit ($\sigma\rightarrow\infty$), since the radio wave in vacuum propagates through the background magnetic field with no interaction and no compression effect.
+A complete picture is revealed by the full time-dependent solution (section VI).
+It shows that the MHD flow pattern in the propagating wave packet is gradually attracted to the steady state, and this relaxation takes a significant time $t_{\star}$ proportional to $\sigma$.
+For propagation times $t<t_{\star}$, the steady state with large $C_{\star}$ is not reached; we evaluate the compression factor $C\ll C_{\star}$ in this situation.
 These results allow one to investigate the FRB interaction with a magnetar wind in a broad range of radii.
-We describe the wind in section 7, and then investigate the compression front induced by the FRB in the wind.
+We describe the wind in section VII, and then investigate the compression front induced by the FRB in the wind.
 We identify two zones, with regular and stochastic heating, and find the transition radius $r_{\rm stoch}$ between them.
-Compression fronts from FRBs emitted in each zone are described in sections 8 and 9.
-FRB damping by its interaction with the wind is investigated in section 10.
-Our conclusions are summarized in section 11.
-2 Problem formulation
-2.1 The role of magnetic field
+Compression fronts from FRBs emitted in each zone are described in sections VIII and IX.
+FRB damping by its interaction with the wind is investigated in section X.
+Our conclusions are summarized in section XI.
+II Problem formulation
+II.1 The role of magnetic field
 For an unmagnetized plasma, particle acceleration in a strong electromagnetic wave is a textbook problem (Landau & Lifshitz, 1975) well known in the context of laser experiments (Bulanov et al., 2015).
 It was also discussed in the context of early pulsar models (Gunn & Ostriker, 1971).
-The wave accelerates initially static particles to Lorentz factors $\gamma=1+a^{2}/2$, and their average velocity is aligned with the wave propagation direction.
-This average motion gives the plasma the bulk Lorentz factor $\Gamma\approx a\gg 1$.
+The wave accelerates initially static particles to the Lorentz factor $\gamma=1+a^{2}/2$.
+The particle oscillating in the wave has a large mean velocity aligned with the wave propagation direction.
+This mean velocity gives the plasma bulk Lorentz factor $\Gamma\approx a\gg 1$.
 The problem qualitatively changes when the background plasma is magnetized (B22).
-Now, in addition to the oscillating electromagnetic field of the wave, there is a background magnetic field $\boldsymbol{B}_{\rm bg}$, which arrests the particle motion across $\boldsymbol{B}_{\rm bg}$ and changes the frame where the particles have zero average momentum (the “fluid frame”).
-In the media surrounding FRB sources, $B_{\rm bg}$ is well below the radio wave amplitude $E_{0}$ at all radii $r>R_{\rm LC}$ considered in this paper.
-At the same time, $B_{\rm bg}$ far exceeds a minimum value sufficient to “magnetize” the plasma: the particle Larmor period is much smaller than the duration of the wave packet $T\sim 1$ ms.
-Besides the wave strength parameter $a$, the problem of wave-plasma interaction has two more dimensionless parameters: the plasma magnetization $\sigma$ and the ratio of gyrofrequency $\omega_{B}=eB_{\rm bg}/mc$ to the wave frequency $\omega$.
-The ratio $\omega_{B}/\omega$ plays a particularly important role.
-B22 showed that waves with $\omega>\omega_{B}$ drive a regular particle motion in the fluid frame: fast oscillations with wave frequency $\omega$ and Lorentz factor $\gamma\approx a$, superposed with (subdominant) slow gyration in the background magnetic field.
-If $1<\omega_{B}/\omega<a$, gyration is accompanied by a resonant wave-particle interaction each Larmor period, leading to quick stochastic particle acceleration to $\gamma\gg a$.
-Oscillations in the wave then become energetically subdominant compared to the Larmor rotation with the stochastically pumped $\gamma$.
-The stochastic heating gives a well-defined mean expectation for the particle energy, which has been found analytically and confirmed numerically.
-In addition to heating the plasma, the radio wave also deposits momentum, pushing the plasma along the wave direction.
-The effective inertia of a plasma with $\sigma\gg 1$ is increased by its magnetic field, which reduces its bulk acceleration.
-The bulk acceleration leads to compression of the plasma and its magnetic field (Beloborodov, 2021).
-Dynamics of a plasma exposed to the radio wave is easiest to investigate in the plane geometry, assuming a uniform magnetized background ahead of the wave.
-Details of the radio wave are not important, however one needs to choose a concrete shape of the wave packet when simulating the plasma dynamics numerically.
-2.2 Wave packet
-As an example, consider a nearly harmonic plane electromagnetic wave with amplitude $E_{0}$, frequency $\omega$, and duration $T\gg 2\pi/\omega$. Its strength parameter (defined in Equation (1)) is
-$a=\frac{a_{0}}{\sqrt{2}},\qquad a_{0}=\frac{eE_{0}}{mc\omega}.$ (3)
-The wave speed is weakly affected by the presence of a tenuous plasma, so it propagates with speed of light $c$. It will be convenient to use the coordinate
+Now, in addition to the oscillating electromagnetic field of the wave, there is a background magnetic field $\boldsymbol{B}_{\rm bg}$, which arrests the particle motion across $\boldsymbol{B}_{\rm bg}$ and changes the frame where the particles have zero average momentum (the fluid frame).
+In the media surrounding FRB sources, $B_{\rm bg}$ is well below the radio wave amplitude $E_{0}$ at radii $r>R_{\rm LC}$ considered in this paper.
+At the same time, $B_{\rm bg}$ is sufficiently strong to “magnetize” the plasma: the particle Larmor period is far smaller than the duration of the radio wave packet $T\sim 1$ ms.
+Larmor rotation couples the plasma to $B_{\rm bg}$, and together they form an MHD fluid, which becomes heated and pushed by the radio wave.
+Besides the wave strength parameter $a$, the problem of FRB-plasma interaction has two more dimensionless parameters: the background plasma magnetization $\sigma$ and the ratio of gyrofrequency $\omega_{B}=eB_{\rm bg}/mc$ to the wave frequency $\omega$.
+The parameters $a$ and $\omega_{B}/\omega$ determine whether the plasma particles execute regular oscillations or experience stochastic heating in the radio wave.
+In addition to heating, the radio wave also deposits momentum, pushing the plasma along the wave direction.
+This effect depends on $\sigma$: $\sigma\gg 1$ implies that the effective bulk inertia of MHD fluid is significantly increased by its magnetic field, which can slow down its bulk acceleration.
+However, the acceleration effect may still be strong in waves with $a\gg 1$.
+Dynamics of a magnetized plasma exposed to the radio wave is easiest to calculate in the plane geometry, assuming a uniform magnetized background ahead of the wave packet.
+Details of the radio wave are not important; however, when simulating the plasma dynamics numerically, one needs to choose a concrete shape of the wave packet.
+II.2 Radio wave packet
+The wave speed is weakly affected by the presence of a tenuous plasma, and it will be taken equal to $c$ (this approximation is discussed in Appendix A). It is convenient to use the coordinate
 $\xi\equiv t-\frac{z}{c},$ (3)
-where the $z$-axis is chosen along the wave propagation direction.
-The radio wave packet is static in the $\xi$-coordinate and occupies the region $0<\xi<T$.
-The wave amplitude is zero at the leading edge of the packet $\xi=0$, rises on a timescale $\xi_{\rm rise}<T$, and then drops back to zero at $\xi=T$.
+where the $z$-axis is along the wave propagation direction.
+The radio wave packet is static in the $\xi$-coordinate and occupies a finite region $0<\xi<T$.
+As an example, consider a nearly harmonic plane electromagnetic wave with amplitude $E_{0}$, frequency $\omega$, and duration $T\gg 2\pi/\omega$.
+The wave electric field,
+$E_{\rm wave}=E_{0}(\xi)\sin(\omega\xi),$ (4)
+oscillates with a high $\omega$ and a gradually varying amplitude $E_{0}(\xi)$; $E_{0}=0$ at the leading edge of the packet $\xi=0$, rises on a timescale $\xi_{\rm rise}\gg\omega^{-1}$, and then drops back to zero at $\xi=T$.
+Using $cE_{\rm wave}=-dA_{\rm wave}/d\xi$ one finds the wave potential at $\xi\omega\gg 1$: $A_{\rm wave}\approx A_{0}\cos(\omega\xi)$ where $A_{0}=cE_{0}/\omega$. So,
+$a_{\rm wave}=a_{0}(\xi)\cos(\omega\xi),\qquad a_{0}=\frac{eE_{0}}{mc\omega},$ (5)
+where $a_{\rm wave}\equiv eA_{\rm wave}/mc^{2}$. The wave strength parameter is
+$a\equiv\langle a_{\rm wave}^{2}\rangle^{1/2}=\frac{a_{0}}{\sqrt{2}}.$ (6)
 In numerical examples we will use packets with the envelope
@@
 The cold unperturbed medium ahead of the packet (at $\xi<0$) will be called “upstream.” The upstream medium is described by a uniform mass density $\rho_{\rm u}$ and a uniform magnetic field $\boldsymbol{B}_{\rm u}$.
-We will focus on the problem with $\boldsymbol{B}_{\rm u}$ oriented perpendicular to the wave propagation direction (as expected in winds, which carry magnetic fields transverse to the radial direction).
+We will focus on the case with $\boldsymbol{B}_{\rm u}$ oriented perpendicular to the wave propagation direction (as expected in winds, which carry magnetic fields transverse to the radial direction).
+[^3] The problem formulation with transverse $\boldsymbol{B}_{\rm bg}$ is not special or degenerate in any way.
+One could also consider $\boldsymbol{B}_{\rm bg}$ with arbitrary angle $\alpha$ with respect to the wave propagation direction.
+The general case with any $\alpha\neq 0$ is reduced to the propagation problem with $\alpha=\pi/2$ by using a new reference frame that moves along $\boldsymbol{B}_{\rm bg}$ with Lorentz factor $\gamma=(\sin\alpha)^{-1}$ (Beloborodov, 2024).
 Two possible polarizations of the radio wave can be considered: the X-mode ($\boldsymbol{E}_{\rm wave}\perp\boldsymbol{B}_{\rm u}$) and the O-mode ($\boldsymbol{E}_{\rm wave}\parallel\boldsymbol{B}_{\rm u}$).
@@
 The X-mode has the oscillating magnetic field $\boldsymbol{B}_{\rm wave}$ parallel to the background field $\boldsymbol{B}_{\rm u}$.
-2.3 Plasma drift through the wave packet
+II.3 Plasma drift through the wave packet
 The plasma behaves as an MHD fluid on timescales longer than the Larmor period.
@@
 Here $\overline{\boldsymbol{E}}$ and $\overline{\boldsymbol{B}}$ are the MHD fields; the bar signifies averaging over a timescale longer than the Larmor time (and longer than the wave oscillation period).
-The average fields can also be denoted as $\boldsymbol{E}_{\rm bg}$ and $\boldsymbol{B}_{\rm bg}$ because they serve as the effective background fields in the problem of particle motion in the wave, as described below.
+The average fields can also be denoted as $\boldsymbol{E}_{\rm bg}$ and $\boldsymbol{B}_{\rm bg}$ because they serve as the effective background fields in the problem of test particle motion in the wave, as described below.
 A natural choice for a reference frame in the wave-plasma interaction problem is such that the upstream plasma is at rest.
-When discussing wave propagation in a wind, we will also view the problem in the lab frame where the upstream moves with a large speed $\beta_{\rm u}>0$ along the wave propagation direction.
+When discussing wave propagation in a wind, we will also view the problem in the lab frame where the upstream plasma moves with a large speed $\beta_{\rm u}>0$ along the wave propagation direction.
 In addition, frames with opposite motion, where the upstream plasma has $\beta_{\rm u}<0$, may be useful in numerical simulations.
@@
 Dynamical equations for $u_{\rm D}^{\mu}$ and final results will be formulated in a Lorentz-invariant way, so they will hold for any frame boosted along $z$ or, equivalently, any $\beta_{\rm u}$.
-Plasma with any speed $v_{\rm D}=dz/dt$ changes its $\xi$-coordinate with rate
+One can also describe fluid motion using the quantity
+$\kappa\equiv\gamma_{\rm D}(1+\beta_{\rm D})=\frac{1}{\gamma_{\rm D}(1-\beta_{\rm D})}.$ (10)
+This simplifies transformations between different frames moving along $z$.
+Relativistic addition of velocities corresponds to multiplication of $\kappa$, and changing $\beta_{\rm D}\rightarrow-\beta_{\rm D}$ corresponds to $\kappa\rightarrow\kappa^{-1}$.
+Doppler transformation of the wave frequency satisfies $\omega/\kappa={\rm inv}$.
+Transformation of other quantities follow from $a={\rm inv}$ (strength parameter of the wave) and $\omega\xi={\rm inv}$ (phase of the wave).
+In particular, transformations of the wave electric field $E_{\rm wave}$, wave energy density $U$, and coordinate $\xi$ measured in two reference frames $\cal K$} and $\tilde{\cal K}$ are
+$\frac{\tilde{E}_{\rm wave}}{E_{\rm wave}}=\frac{\xi}{\tilde{\xi}}=\frac{\tilde{\omega}}{\omega}=\frac{\tilde{\kappa}}{\kappa},\qquad\frac{\tilde{U}}{U}=\frac{\tilde{\kappa}^{2}}{\kappa^{2}}.$ (11)
+Hereafter, we choose $\tilde{\cal K}$ as the local rest frame of the fluid, [^4] Throughout this paper, we use tilde for quantities measured in the fluid frame $\tilde{\cal K}$.
+Tilde will be omitted if this does not lead to confusion: we omit tilde for thermodynamic quantities (e.g.
+pressure $P$ and enthalpy $w$, which are defined only in frame $\tilde{\cal K}$).
+which corresponds to
+$\tilde{\beta}_{\rm D}=0,\qquad\tilde{\kappa}=1.$ (12)
+Plasma with any speed $v_{\rm D}=dz/dt$ (in any frame) changes its $\xi$-coordinate with rate
 $\frac{d\xi}{dt}=1-\beta_{\rm D}>0.$ (13)
-Thus, the MHD fluid always moves toward increasing $\xi$ through the wave packet.
+Thus, MHD fluid always moves toward increasing $\xi$ through the wave packet.
 When viewed in the rest frame of the upstream medium, the initially static fluid enters the wave packet at $\xi=0$ and develops speed $\beta_{\rm D}\neq 0$ at $\xi>0$.
 We wish to find the fluid motion in the packet, $\beta_{\rm D}(\xi)$, and the fluid compression factor $C(\xi)=\rho(\xi)/\rho_{\rm u}$.
-3 Plasma heating by the wave
+III Plasma heating by the radio wave
+The response of a magnetized plasma to a propagating radio wave packet can be disentangled if we first focus on the “micro level” and consider the motion of individual particles in the local fluid frame $\tilde{\cal K}$.
+In this frame, the fluid is at rest and the wave-plasma interaction may be viewed as a local heating effect: the strong high-frequency wave energizes particles to Lorentz factors $\tilde{\gamma}\gg 1$.
+This effect was calculated in B22 by tracking the motion of test particles.
+The heating depends on the wave strength parameter $a$.
+For smooth packets, such that $a$ varies on a scale $\Delta\tilde{\xi}\gg a\tilde{\omega}_{B}$ (typically satisfied by FRBs), the results for $\tilde{\gamma}$ may be summarized as follows (see also Appendix B).
+(1) Waves with frequency $\tilde{\omega}$ exceeding the gyrofrequency $\tilde{\omega}_{B}=e\tilde{B}_{\rm bg}/mc$ drive regular particle oscillations with the average Lorentz factor
+$\langle\tilde{\gamma}\rangle=\sqrt{1+a^{2}}.$ (14)
+The local $\langle\tilde{\gamma}\rangle$ in the packet simply tracks the local strength parameter of the wave $a$.
 […]
-(1) Waves with frequency $\tilde{\omega}$ exceeding the gyrofrequency $\tilde{\omega}_{B}=e\tilde{B}_{\rm bg}/mc$ drive regular particle oscillations with
-$\tilde{\gamma}=\sqrt{1+a^{2}}\approx a\gg 1\qquad(\tilde{\omega}_{B}<\tilde{\omega}).$ (9)
-Thus, the local $\tilde{\gamma}$ in the packet simply tracks the local strength parameter of the wave $a$.
-(2) In the other heating regime, $1<\tilde{\omega}_{B}/\tilde{\omega}<a$, the resonant pumping mechanism is activated, and the wave drives stochastic particle acceleration.
-Then, the average $\tilde{\gamma}$ inside a fluid element quickly exceeds $a$ and keeps growing as the fluid moves through the wave packet (until radiative losses limit the growth or fluid exits behind the packet).
-The mean expectation for the particle Lorentz factor $\tilde{\gamma}$ grows according to the equation (B22,Beloborodov 2024)
-$\frac{d\tilde{\gamma}}{d\tilde{t}}=\chi\frac{a_{0}^{2}\tilde{\omega}_{B}^{1/3}\tilde{\omega}^{2/3}}{\tilde{\gamma}^{4/3}}\qquad(1<\tilde{\omega}_{B}/\tilde{\omega}<a_{0}),$ (10)
-where $a_{0}^{2}=2a^{2}$, $\chi\sim 1$ is a numerical coefficient, and all quantities are measured in the fluid rest frame $\tilde{\cal K}$.
+$\frac{d\langle\tilde{\gamma}\rangle_{\rm ens}}{d\tilde{t}}=\chi\frac{a^{2}\tilde{\omega}_{B}^{1/3}\tilde{\omega}^{2/3}}{\langle\tilde{\gamma}\rangle_{\rm ens}^{4/3}},$ (15)
+[…]
 Calculations in B22 were performed assuming that the background magnetic field $\boldsymbol{B}_{\rm bg}$ is static, i.e.
@@
 In general, $\overline{\boldsymbol{E}}\neq 0$ which corresponds to a moving MHD fluid $v_{\rm D}\neq 0$.
-If $v_{\rm D}$ varies (the plasma is accelerated in the wave packet), no global inertial frame exists where the fluid would be static, and one has to consider the heating effect of the wave in a moving MHD fluid.
-Then, the test particle motion should be examined with $E_{\rm bg}\neq 0$.
-Remarkably, the results found in B22 at $E_{\rm bg}=0$ also hold in the accelerating background, as long as the particle motion is measured in the local drift frame $\tilde{\cal K}$ where $\tilde{\boldsymbol{E}}_{\rm bg}=0$.
-This fact is illustrated below by direct calculations of particle motion in a wave packet with a non-static background field.
+If $v_{\rm D}$ varies (the plasma is accelerated in the wave packet), no global inertial frame exists where the fluid would be static, and one has to consider the heating effect of the radio wave in a moving MHD fluid.
+Then, the test particle motion should be examined with $E_{\rm bg}\neq 0$, as done in this section.
+One goal of our calculations below is to generalize the threshold condition for stochastic heating $b\equiv\tilde{\omega}_{B}/\tilde{\omega}\gtrsim 1$ so that it applies to accelerating fluids.
+If the upstream plasma has $b_{\rm u}\gtrsim 1$, stochastic heating is triggered immediately in the leading part of the wave packet once $a$ exceeds $\sim 1$ (B22).
+If the upstream plasma has $b_{\rm u}\ll 1$, stochastic heating can still occur, but deeper inside the wave packet where $b$ is sufficiently increased by fluid acceleration and compression.
+This can occur after the wave strength parameter has reached a high $a\gg 1$, which changes the threshold for stochastic heating.
+Remarkably, in an accelerating fluid, Equations (14) and (15) still hold, i.e.
+the evolution of $\tilde{\gamma}$ in both regimes remains unchanged.
+This fact will be illustrated below by direct calculations of particle motion in a wave packet with a non-static background field, which defines the fluid motion.
 We will prescribe smoothly varying fields $\boldsymbol{B}_{\rm bg}(\xi)$ and $\boldsymbol{E}_{\rm bg}(\xi)$ (varying on a scale much greater than the particle Larmor radius) and repeat the calculation of B22 for test particles.
-In the simulations, we will use the frame where the upstream plasma is at rest, $\beta_{\rm D}=\beta_{\rm u}=0$, and prescribe the background field of the form
-$B_{\rm bg}=\frac{\kappa^{2}+1}{2}B_{\rm bg}^{\rm u},\qquad E_{\rm bg}=\frac{\kappa^{2}-1}{2}B_{\rm bg}^{\rm u},$ (12)
-$\kappa\equiv\gamma_{\rm D}(1+\beta_{\rm D})=\frac{1}{\gamma_{\rm D}(1-\beta_{\rm D})}.$ (13)
-The upstream has $\kappa=\kappa_{\rm u}=1$, $E_{\rm bg}=0$, and $B_{\rm bg}=B_{\rm bg}^{\rm u}$.
-Fluid motion inside the wave packet is specified by a chosen function $\kappa(\xi)$.
-For the present purposes, Equation (12) may be viewed as an arbitrary prescription to test heating of fluid with a non-uniform motion $\beta_{\rm D}(\xi)$.
-As explained in next sections, any steady pattern of fluid motion in the wave packet satisfies Equation (12).
+The background fields will be parameterized as follows
+$B_{\rm bg}=\frac{\kappa^{2}+1}{\kappa_{\rm u}^{2}+1}B_{\rm bg}^{\rm u},\qquad E_{\rm bg}=\frac{\kappa^{2}-1}{\kappa_{\rm u}^{2}+1}B_{\rm bg}^{\rm u}.$ (17)
+In the simulations, we will use the frame where the upstream plasma is at rest; then, $E_{\rm bg}^{\rm u}=0$ and $\kappa_{\rm u}=1$.
+Fluid acceleration inside the wave packet is specified by a chosen function $\kappa(\xi)$.
+For the present purposes, Equation (17) may be viewed as an arbitrary prescription to test heating of fluid with non-uniform motion $\beta_{\rm D}(\xi)$, which is related to $\kappa(\xi)$ by Equation (10).
+As explained in next sections, any steady pattern of fluid motion in the wave packet actually satisfies Equation (17).
 It also satisfies $\tilde{\rho}/\kappa=const$ (uniform across the packet), where $\tilde{\rho}=\rho/\gamma_{\rm D}$ is the fluid proper density.
-Thus, the adopted prescription corresponds to a steady flow pattern with the proper compression factor $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}=\kappa$.
-The magnetic field $\tilde{B}_{\rm bg}$ is frozen in the fluid and compressed by the same factor $\kappa$ while $B_{\rm bg}=\gamma_{\rm D}\tilde{B}_{\rm bg}$ obeys Equation (12).
-Our simulation will track a test particle initially located in the static upstream region and exposed to the propagating wave packet.
+Thus, the adopted prescription corresponds to a steady flow pattern with the proper compression factor $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}$.
+The magnetic field $\tilde{B}_{\rm bg}$ is frozen in the fluid and compressed by the same factor $\kappa/\kappa_{\rm u}$ while $B_{\rm bg}=\gamma_{\rm D}\tilde{B}_{\rm bg}$ obeys Equation (17).
+Our simulations will track test particles initially located in the static, cold upstream region exposed to the propagating wave packet.
 The particle motion in the wave packet is found by solving the dynamical equation,
 $mc\,\frac{d\boldsymbol{u}}{dt}=e\left(\boldsymbol{E}+\boldsymbol{\beta}\times\boldsymbol{B}\right),$ (18)
-where vector $\boldsymbol{u}=\gamma\boldsymbol{\beta}$ is the spatial part of the particle four-velocity; radiative losses are neglected here.
+where vector $\boldsymbol{u}=\gamma\boldsymbol{\beta}$ is the three-dimensional spatial part of the particle four-velocity.
+Radiative losses are neglected in Equation (18).
 The fields $\boldsymbol{E}(\xi)$ and $\boldsymbol{B}(\xi)$ include the background and wave fields: $\boldsymbol{E}=\boldsymbol{E}_{\rm bg}+\boldsymbol{E}_{\rm wave}$ and $\boldsymbol{B}=\boldsymbol{B}_{\rm bg}+\boldsymbol{B}_{\rm wave}$.
 [^5] Notation here differs from B22 where $\boldsymbol{E}$ and $\boldsymbol{B}$ denoted the wave fields and the dynamical equation assumed $\boldsymbol{E}_{\rm bg}=0$.
-3.1 Regular heating ($\tilde{\omega}_{B}<\tilde{\omega}$)
+III.1 Regular oscillations
 Let us prescribe $B_{\rm bg}(\xi)$ and $E_{\rm bg}(\xi)$ by choosing $\kappa^{2}$ of the form
-$\kappa^{2}=1+\zeta a^{2},$ (15)
+$\frac{\kappa^{2}}{\kappa_{\rm u}^{2}}=1+\zeta a^{2},$ (19)
 where $\zeta$ is a constant coefficient.
-The prescribed $\kappa(\xi)$ implies the fastest drift of the fluid in the middle of the packet where $a(\xi)$ peaks (Equation 5); the fluid is static ahead and behind the wave packet where $a=0$.
-The case of $\zeta=0$ corresponds to a static magnetic background throughout the packet ($E_{\rm bg}=0$); in this case, the particle motion is similar to that found in B22.
-When $\zeta\neq 0$, the presence of $E_{\rm bg}\neq 0$ changes the particle motion by introducing an additional electric force.
-After solving for the oscillating particle motion in the wave packet, we calculate its Lorentz factor averaged over the wave oscillation period.
-Figure 1 shows the obtained Lorentz factor as a function of position $\xi$ in the packet, in three sample models.
-The models have the same radio wave (packet with amplitude $a_{\max}=4$) and different $\zeta=0$, 0.4, and 1.
-The results demonstrate that the average particle Lorentz factor $\tilde{\gamma}(\xi)$ measured in the local fluid frame is the same in all three models, independent of the fluid motion in the packet, and follows the solution found with a static background (Equation 9).
-In the presented numerical models, we chose $\omega_{B}/\omega=0.05$, where $\omega_{B}\equiv\tilde{\omega}_{B}^{\rm u}$ is the gyro-frequency in the upstream plasma.
-Then, the condition $\tilde{\omega}_{B}<\tilde{\omega}$ is satisfied throughout the packet, so the heating everywhere acts in the regular regime according to Equation (9).
+The simulation will be done in the reference frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$).
+The case of $\zeta=0$ corresponds to a static magnetic background throughout the packet ($\kappa=1$, $E_{\rm bg}=0$); in this case, the particle motion is similar to that found in B22.
+When $\zeta\neq 0$, the increasing $\kappa>1$ means $\beta_{\rm D}>0$ and the appearance of additional electric force $eE_{\rm bg}$, which changes the particle motion.
+The prescribed $\kappa(\xi)$ gives the fastest drift of the fluid in the middle of the packet where $a(\xi)$ peaks; the fluid is static ahead and behind the wave packet where $a=0$.
+We have calculated three sample models with $\zeta=0$, 0.4, and 1.
+In all three models, the radio wave packet has the amplitude $a_{\max}=4$, and the upstream plasma has the gyro-frequency $\tilde{\omega}_{B}^{\rm u}=0.03\omega$, where $\omega$ is the wave frequency measured in the upstream rest frame.
+These parameters represent an FRB propagating through the magnetar wind at radii $r\sim 3\times 10^{12}$ cm, as discussed in section VII below.
+Another parameter is the number of wave oscillations in the packet, $N=\omega T/2\pi$.
+Real FRBs have $N\sim 10^{6},$ and the presented simulations have $N=10^{3}$.
+We have also run models with different $N$ and verified that the results are independent of $N$ as long as $N\gg a\tilde{\omega}/\tilde{\omega}_{B}$ (this is simply the condition that the packet is much wider than the particle Larmor radius).
+For each model, we have solved for the particle motion in the radio wave packet, and then calculated its Lorentz factor averaged over the wave oscillation period: $\gamma$ in the lab frame and $\tilde{\gamma}$ in the local fluid rest frame.
+Figure 1 shows the average $\gamma$ and $\tilde{\gamma}$ as functions of position $\xi$ in the packet.
+It demonstrates that $\tilde{\gamma}=\sqrt{1+a^{2}}$ holds in all three models ($\zeta=0$, 0.4, 1).
+In the fluid frame, the particle performs same regular oscillations, independent of the fluid bulk acceleration in the packet.
+Since fluid motion $\kappa$ varies on scales much larger than the Larmor scale, Larmor rotation is not excited (B22).
+The particle orbit measured in the local fluid frame is nearly identical to the canonical “8”-shaped orbit, same as found in the absence of any background field (Appendix B).
 Figure 1: Lorentz factor $\gamma$ of a test particle moving through the wave packet (averaged over the wave oscillation period).
-The upstream plasma is at rest ($\gamma=1$) and has the gyro-frequency $\omega_{B}=0.05\omega$.
-The plasma particle moves through the wave packet in the coordinate $\xi=t-z/c$ from right to left.
+The upstream plasma is cold ($\tilde{\gamma}=1$) and static ($\kappa_{\rm u}=1$); it has gyro-frequency $\tilde{\omega}_{B}^{\rm u}=0.03\omega$ measured in the fluid frame.
+The test particle moves through the wave packet in the coordinate $\xi=t-z/c$ from right to left.
 The simulated packet consists of $10^{3}$ oscillations with an envelope described by Equation (7) with $a_{\max}=4$, and the background field is described by Equations (17) and (19).
@@
 For each model, dashed curve shows $\gamma$ measured in the fixed frame (where the upstream plasma is at rest), and solid curve shows $\tilde{\gamma}$ measured in the local fluid rest frame.
-All solid curves overlap, demonstrating that $\tilde{\gamma}(\xi)$ is independent of the local fluid velocity $\beta_{\rm D}=E_{\rm bg}/B_{\rm bg}$ and follows the solution $\tilde{\gamma}^{2}=1+a^{2}(\xi)$ (red dotted curve).
-The test-particle motion changes if $\omega_{B}/\omega$ is increased so that $\tilde{\omega}_{B}/\tilde{\omega}$ exceeds unity somewhere inside the packet, triggering stochastic acceleration.
-Compression of $\tilde{B}_{\rm bg}$ by the factor of $\kappa$ implies $\tilde{\omega}_{B}=\kappa\omega_{B}$, and the Doppler-shifted wave frequency in the fluid frame is $\tilde{\omega}=\omega/\kappa$.
-This gives
-$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}=\kappa^{2}\,\frac{\omega_{B}}{\omega}.$ (16)
+All three solid curves overlap, demonstrating that $\tilde{\gamma}(\xi)$ is independent of the local fluid velocity $\beta_{\rm D}=E_{\rm bg}/B_{\rm bg}$ and follows the solution $\tilde{\gamma}^{2}=1+a^{2}(\xi)$ (red dotted curve, overlaping the solid curves).
+III.2 Stochastic heating
+The onset of stochastic heating is controlled by two parameters: $a$ and
+$b\equiv\frac{\tilde{\omega}_{B}}{\tilde{\omega}}$. (20)
+As an example, consider a radio wave packet with amplitude $a_{\max}=30$ propagating in an ambient plasma with gyrofrequency $\tilde{\omega}_{B}^{\rm u}=0.2\tilde{\omega}_{\rm u}$, so the upstream plasma has $b_{\rm u}=0.2$.
+These parameters may represent an FRB propagating through a magnetar wind at $r\sim 5\times 10^{11}$ cm (section VII).
+In this case, the fluid is accelerated and compressed inside the wave packet by a large factor $\kappa/\kappa_{\rm u}\gg 1$, and $b\gg 1$ can be reached inside the packet.
+Indeed, Equation (17) implies compression of $\tilde{B}_{\rm bg}$ by the factor of $\kappa/\kappa_{\rm u}$, so the gyrofrequency measured in the local fluid frame $\tilde{\cal K}$ is increased to $\tilde{\omega}_{B}=(\kappa/\kappa_{\rm u})\tilde{\omega}_{B}^{\rm u}$.
+At the same time, the wave frequency measured in the fluid frame is reduced from $\tilde{\omega}_{\rm u}=\omega/\kappa_{\rm u}$ to $\tilde{\omega}=\omega/\kappa$.
+This gives $b=(\kappa/\kappa_{\rm u})^{2}b_{\rm u}$.
+The threshold condition for the onset of stochastic heating can be stated in the form $b>b_{\rm s}(a)$, as discussed in section III.3 below.
+The transition to stochastic heating occurs for any profile of $\kappa(\xi)$ as long as it reaches $b>b_{\rm s}$, i.e.
+$\kappa/\kappa_{\rm u}$ exceeds $(b_{\rm s}/b_{\rm u})^{1/2}$.
+In our example simulations, we prescribe $\kappa(\xi)$ using a modified version of Equation (19): [^6] We use the simplest modification of Equation (19) that satisfies $d\kappa/d\xi\geq 0$.
+This condition is motivated by the results of later sections that describe self-consistent calculations of $\kappa(\xi)$.
+Unlike regular oscillations (where $\tilde{\gamma}$ tracks the rise and drop of $a(\xi)$ in the wave packet), stochastic heating gives a monotonically growing expectation value $\tilde{\gamma}(\xi)$: $d\tilde{\gamma}/d\xi\geq 0$, leading to $d\kappa/d\xi\geq 0$.
+$\frac{\kappa^{2}}{\kappa_{\rm u}^{2}}=\left\{\begin{array}[]{cc}\vskip 2.84526pt1+\zeta a^{2}&\qquad\xi<T/2\\ 1+\zeta a_{\max}^{2}&\qquad\xi>T/2\end{array}\right$. (23)
+The results will be shown for two sample models: $\zeta=0$ (no fluid acceleration in the packet) and $\zeta=1$.
+Measurements of stochastic heating require calculations of many realizations of particle motion in the wave packet.
+We draw 4000 upstream particles from an initial Maxwellian distribution with a small temperature $kT=10^{-2}mc^{2}$, seeding initial perturbations for the stochastic behavior.
+The particle motion is followed for each of the 4000 realizations, and the result is averaged to find the mean value $\tilde{\gamma}(\xi)$.
+The results are shown in Figure 2.
+In the model with $\zeta=0$, the ratio $b\equiv\tilde{\omega}_{B}/\tilde{\omega}=0.2$ stays uniform and small across the wave packet.
+Then, the stochastic heating does not occur: the plasma everywhere performs regular oscillations with $\tilde{\gamma}=\sqrt{1+a^{2}}$.
+By contrast, in the model with $\zeta=1$, the background field is subject to strong compression, $\kappa/\kappa_{\rm u}=\sqrt{1+a^{2}}$, which changes the particle response to the radio wave.
+In the leading part of the packet, where $\kappa/\kappa_{\rm u}$ is still small, the particles perform regular oscillations with $\tilde{\gamma}=\sqrt{1+a^{2}}$, same as in the model with $\zeta=0$.
+The transition to stochastic heating is observed at $\xi_{\rm stoch}\approx 140\times 2\pi/\omega$, where $\tilde{\gamma}(\xi)$ shows a steep growth.
 […]
-3.2 Stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$)
-First, let us formulate the analytical prediction for stochastic heating by solving Equation (10) for $\tilde{\gamma}(\xi)$.
-The fluid proper time $\tilde{t}$ is related to time $t$ in the lab frame by $d\tilde{t}=dt/\gamma_{\rm D}$.
-During time $dt$, the fluid element moves in $\xi$ by $d\xi=(1-\beta_{\rm D})dt$, so
 $d\tilde{t}=\frac{d\xi}{\gamma_{\rm D}(1-\beta_{\rm D})}=\kappa\,d\xi.$ (24)
-Using also $\tilde{\omega}=\omega/\kappa$ and $\tilde{\omega}_{B}=\kappa\omega_{B}$ (where $\omega_{B}\equiv\tilde{\omega}_{B}^{\rm u}$ is the upstream gyrofrequency), we rewrite Equation (10) as
-$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}=\chi a_{0}^{2}\kappa^{2/3}\left(\omega_{B}\omega^{2}\right)^{1/3}.$ (18)
-For any chosen $\kappa(\xi)$, Equation (18) can be integrated for $\tilde{\gamma}(\xi)$, and this analytical prediction can be compared with simulations of test particle motion.
-Unlike regular heating (where $\tilde{\gamma}$ tracks the rise and drop of $a$), stochastic heating gives a monotonically growing mean expectation $\tilde{\gamma}(\xi)$: $d\tilde{\gamma}/d\xi\geq 0$.
-Our simulation setup is described above; in particular, $B_{\rm bg}$ and $E_{\rm bg}$ are given by Equation (12).
-We arbitrarily choose $\kappa(\xi)$ of the form, [^5] Plasma in packets with monotonically growing $\tilde{\gamma}(\xi)$ is expected to have a monotonically growing $\kappa(\xi)$.
-Therefore, in sample models of stochastic heating we choose $\kappa(\xi)$ that satisfy $d\kappa/d\xi\geq 0$.
-$\kappa=\left\{\begin{array}[]{cc}1+(\kappa_{0}-1)\sin^{2}(\pi\xi/T)&\qquad\xi<T/2\\ \kappa_{0}&\qquad\xi>T/2\end{array}\right.$ (21)
-so $\kappa(\xi)$ grows until it reaches $\kappa_{0}$ in the middle of the wave packet and stays constant in the second half of the packet.
-We have calculated particle motion in packets with $\kappa_{0}=1$ (no fluid motion, $\beta_{\rm D}=0$) and $\kappa_{0}=3$.
-Figure 2: Average Lorentz factor $\tilde{\gamma}$ (measured in the local fluid rest frame) of particles crossing the wave packet in the regime of stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$).
+Using also $\tilde{\omega}=\omega/\kappa$ and $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$, we rewrite Equation (15) as
+$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}=\chi\,a^{2}\kappa^{2/3}\left(\frac{\tilde{\omega}_{B}^{\rm u}\omega^{2}}{\kappa_{\rm u}}\right)^{1/3}.$ (25)
+For any chosen $\kappa(\xi)$, Equation (25) can be integrated for $\tilde{\gamma}(\xi)$.
+As one can see in Figure 2, this analytical calculation reproduces the mean value $\tilde{\gamma}(\xi)$ found in the full dynamical simulation with the ensemble of 4000 particles.
+We have also calculated models with different $a_{\max}$, $\tilde{\omega}_{B}^{\rm u}/\tilde{\omega}_{\rm u}$, and $\zeta$, and found that Equation (25) with $\chi\approx 0.8$ well describes stochastic heating in all of them.
+Figure 2: Average Lorentz factor $\tilde{\gamma}$ (measured in the local fluid rest frame) of particles interacting with the wave packet with amplitude $a_{\max}=30$.
+The upstream plasma is at rest ($\kappa_{\rm u}=1$) and has a small temperature $kT=10^{-2}mc^{2}$; its gyro-frequency is $\tilde{\omega}_{B}^{\rm u}=0.2\tilde{\omega}_{\rm u}$.
 The average $\tilde{\gamma}$ is calculated by tracking $4\times 10^{3}$ test particles.
-The upstream plasma is at rest ($\gamma_{\rm D}=1$) and has a small temperature $kT=10^{-2}mc^{2}$; its gyro-frequency is $\omega_{B}=2\omega$.
-The packet $0<\xi<T$ consists of $N=4\times 10^{3}$ oscillations; its envelope $a_{0}=\sqrt{2}a$ is described by Equation (5) with $a_{0}^{\max}=30$ (blue curve).
-Background field is described by Equations (12) and (21).
-As one can see in the figure, the plasma experiences heating as it moves through the wave ($\tilde{\gamma}$ grows with $\xi$).
-Black solid curves show the simulation results for $\kappa_{0}=1$ (static background field, no bulk motion of the plasma) and $\kappa_{0}=3$ ($\beta_{\rm D}^{\max}=0.8$).
-The results approximately agree with the analytical prediction (Equation (18), red dotted curves) with $\chi=0.66$ for $\kappa_{0}=1$ and $\chi=0.67$ for $\kappa_{0}=3$.
-The stochastic nature of heating requires calculations of many realizations of particle motion in the packet.
-We draw 4000 particles from an initial Maxwellian distribution with a small temperature $kT=10^{-2}mc^{2}$, seeding small initial perturbations for the stochastic behavior.
-The particle motion is followed for each of the 4000 realizations, and the result is averaged to find the mean expectation for $\tilde{\gamma}(\xi)$.
-The results are shown in Figure 2 for $\omega_{B}/\omega=2$ and $a_{0}=30$.
-In both shown cases, $\kappa_{0}=1$ and $\kappa_{0}=3$, the found $\tilde{\gamma}(\xi)$ is in good agreement with the analytical prediction given by Equation (18) with $\chi\approx 2/3$.
-In summary, the results of B22 for regular and stochastic heating hold in the local frame $\tilde{\cal K}$ regardless of fluid motion, as long as the fluid speed varies on timescales much longer than the Larmor period of particles.
-The Lorentz factor $\tilde{\gamma}\gg 1$ implies a strong heating effect, giving the fluid a relativistic effective temperature.
-The energized particles gyrate in the plane perpendicular to $\boldsymbol{B}_{\rm bg}$, so the plasma behaves as a two-dimensional relativistic gas.
-Its internal energy density $U_{\rm p}$ and pressure $P_{\rm p}$ (measured in frame $\tilde{\cal K}$) are
-$U_{\rm p}=2P_{\rm p}=(\tilde{\gamma}-1)\tilde{\rho}c^{2},\qquad\tilde{\rho}=\frac{\rho}{\gamma_{\rm D}}.$ (22)
-The heated particles gyrate with the Larmor frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}\ll\tilde{\omega}_{B}$.
-It is below the wave frequency $\tilde{\omega}$ in both heating regimes, $\tilde{\omega}_{B}<\tilde{\omega}$ and $\tilde{\omega}_{B}>\tilde{\omega}$ (B22).
-4 MHD equations for fluid motion
-In the test-particle numerical experiments described above, we prescribed an arbitrary pattern of the fluid drift speed $\beta_{\rm D}(\xi)$.
+Background fields $B_{\rm bg}$ and $E_{\rm bg}$ are described by Equations (17) and (23).
+The model with $\zeta=0$ corresponds to no fluid acceleration in the wave packet and shows regular oscillations with $\tilde{\gamma}=\sqrt{1+a^{2}}$.
+The model with $\zeta=1$ demonstrates the transition to stochastic heating caused by the strong acceleration and compression of the fluid.
+Red dotted curves show the analytical predictions for regular oscillations ($\tilde{\gamma}=\sqrt{1+a^{2}}$) and stochastic heating (Equation (25) with $\chi=0.8$).
+Figure 3: Transition curve separating the regimes of regular oscillations and stochastic heating on the $a$-$b$ plane.
+Black circles show the results of 12 simulations: each circle shows $(a,b)$ at which the transition from regular oscillations to stochastic heating occurred as the particles moved through the wave packet.
+Dotted curve shows $b=(1/3)\sqrt{1+a^{2}}$.
+The region above the solid curve satisfies $b>2\sqrt{2}\,a$, which corresponds to wave amplitude $\tilde{E}_{0}<\tilde{B}_{\rm bg}/2$.
+In this region, the electromagnetic field satisfies $E^{2}<B^{2}$ throughout the radio wave oscillation, and stochastic heating is not expected.
+The radio wave cannot compress the background far above the solid line, as this would require energy exceeding the available energy of the wave.
+As illustrated in Figures 1 and 2, the particle response to the radio wave admits a simple analytical description in frame $\tilde{\cal K}$ regardless of how this frame moves in the radio wave packet.
+This universality is foreseeable.
+The speed of frame $\tilde{\cal K}$ varies on timescales much longer than the particle Larmor period; so, in the problem of particle response to the radio wave, $\tilde{\cal K}$ is not too different from an inertial frame.
+As a result, $\tilde{\gamma}$ evolves with the same rate as in a non-accelerating fluid: it follows Equation (14) in the regular oscillation regime and Equation (25) in the stochastic heating regime.
+The large Lorentz factor $\tilde{\gamma}\gg 1$ may be thought of as a heating effect, giving the fluid a relativistic effective enthalpy and increasing its effective inertial mass.
+Note that the Larmor frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}$ remains below the wave frequency $\tilde{\omega}$ for both regular oscillation and stochastic heating regimes (B22).
+The ratio $\tilde{\omega}_{\rm L}/\tilde{\omega}$ reaches a maximum value of $\approx 1/3$ at the transition between the two regimes, which is described below.
+III.3 Transition curve on the $a$-$b$ plane
+Each test-particle simulation that shows the transition to stochastic heating inside the wave packet at some point $\xi_{\rm stoch}$ gives a measurement of $a(\xi_{\rm stoch})$ and $b(\xi_{\rm stoch})$.
+These measurements are shown in Figure 3 for a set of simulations with various $b_{\rm u}$, $a_{\max}$, and $\zeta$.
+We kept $b_{\rm u}$ below 0.4 in all cases; then, stochastic heating is not activated without accelerating/compressing the background.
+Compression enables the transition by increasing $b$ from $b_{\rm u}$ by the factor $b/b_{\rm u}=\kappa^{2}/\kappa_{\rm u}^{2}=1+\zeta a^{2}$.
+As one can see in Figure 3, the transition occurs on a well defined curve in the $a$-$b$ plane.
+It is independent of the shape of the wave packet $a(\xi)$ or the profile of background compression in the packet (here parameterized by $\zeta$); it is also independent of $b_{\rm u}$.
+Basically, the transition from regular oscillations to stochastic heating is controlled by local $a$ and $b$.
+This is expected, since the local plasma response to the wave viewed in the fluid frame is independent of the drift of this frame or the global shape of the wave packet.
+The transition curve extends from $a\sim 0.3$ to $a\gg 1$, and approximately follows the relation
+$b=b_{\rm s}\approx\frac{1}{3}\sqrt{1+a^{2}}\qquad{\rm(transition)}.$ (26)
+At $b<b_{\rm s}$, the plasma performs regular oscillations in the radio wave with frequency $\tilde{\omega}$ while Larmor rotation (with frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}=\tilde{\omega}_{B}/\sqrt{1+a^{2}}<\tilde{\omega}/3$) is not excited.
+Excitation of Larmor rotation and transition to stochastic heating occur when $\tilde{\omega}_{\rm L}$ exceeds $\tilde{\omega}/3$ (Larmor radius $\tilde{r}_{\rm L}=c/\tilde{\omega}_{\rm L}$ exceeds half wavelength $\pi c/\tilde{\omega}$), which corresponds to $b>b_{\rm s}$.
+At this point, the quick growth of $\tilde{\gamma}$ is triggered, with $\tilde{\gamma}$ reaching very high values and $\tilde{\omega}_{\rm L}$ decreasing far below $\tilde{\omega}$.
+Within a numerical factor, Equation (26) can be obtained by combining $\tilde{\gamma}\sim a$ (achieved in regular oscillations) with the condition for triggering stochastic heating $\tilde{\gamma}\gtrsim\tilde{\gamma}_{\star}\approx(a^{3}/b)^{1/2}$ (B22).
+The transition to stochastic heating at $\tilde{\gamma}\sim\tilde{\gamma}_{\star}$ may be thought of as the onset of synchrotron absorption of the wave.
+IV MHD equations for fluid motion
+In the test-particle numerical experiments described above, we prescribed an arbitrary pattern of the fluid drift speed $\beta_{\rm D}(\xi)$ or, equivalently, $\kappa(\xi)$.
 Our goal now is to find the actual pattern of fluid motion in the radio wave packet.
 It can be quantified by the profiles of two quantities: $\kappa$ (related to $\beta_{\rm D}$) and $\tilde{C}=\tilde{\rho}/\tilde{\rho}_{\rm u}$ (the proper compression factor).
-They are equal, $\tilde{C}=\kappa$, when a steady pattern of fluid motion is established in the propagating wave packet, i.e.
-when $\kappa$ and $\tilde{\rho}$ are functions of $\xi$ only (see below).
+They turn out equal, $\tilde{C}=\kappa$, when a steady pattern of fluid motion is established in the propagating wave packet, i.e.
+when $\kappa$ and $\tilde{\rho}$ are functions of $\xi$ only.
 The full problem, however, involves evolving profiles.
@@
 We first state the equations in the usual variables $t,z$ and later switch to the more convenient coordinates $t,\xi$.
-4.1 Conservation laws
+IV.1 Conservation laws
 Fluid motion in the wave packet obeys conservation of mass, energy, and momentum.
@@
 $T^{\mu\nu}=T^{\mu\nu}_{\rm p}+T^{\mu\nu}_{\rm f}+T^{\mu\nu}_{\rm w}.$ (28)
-The relevant components of $T^{\mu\nu}$ are $T^{tt}$, $T^{tz}$, and $T^{zz}$. In particular, the wave stress-energy tensor is
+The relevant components of $T^{\mu\nu}$ are $T^{tt}$, $T^{tz}$, and $T^{zz}$. In particular, the stress-energy tensor of the radio wave is
 $T^{zz}_{\rm w}=T^{tz}_{\rm w}=T^{tt}_{\rm w}=U=\frac{\langle E_{\rm wave}^{2}\rangle}{4\pi}=\frac{m\,\omega^{2}a^{2}}{4\pi r_{e}},$ (29)
 where $r_{e}=e^{2}/mc^{2}$ is the classical electron radius.
-The plasma stress-energy tensor is
-$T^{\mu\nu}_{\rm p}=(\tilde{\rho}c^{2}+U_{\rm p}+P_{\rm p})u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P_{\rm p},$ (26)
-where $\eta^{\mu\nu}$ is the Minkowski metric.
-The plasma pressure $P_{\rm p}$ and internal energy $U_{\rm p}$ (Equation 22) are controlled by the wave strength $a(\xi)$.
-In particular, $\tilde{\gamma}=\sqrt{1+a^{2}}$ in the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}$).
-Relevant components of the stress-energy tensor of the MHD electromagnetic field are
-$T^{tt}_{\rm f}=T^{zz}_{\rm f}=\frac{B^{2}+E^{2}}{8\pi},\qquad T^{tz}_{\rm f}=\frac{EB}{4\pi},$ (27)
+Consider now the relevant components of the stress-energy tensor of the MHD electromagnetic field:
+$T^{tt}_{\rm f}=T^{zz}_{\rm f}=\frac{B^{2}+E^{2}}{8\pi},\qquad T^{tz}_{\rm f}=\frac{EB}{4\pi}.$ (30)
 Hereafter, we drop the bar in $\overline{E}$ and $\overline{B}$ — in all remaining sections, $\boldsymbol{E}$ and $\boldsymbol{B}$ will have the meaning of MHD fields averaged over the Larmor timescale.
@@
 Magnetic flux freezing in each fluid element implies
-$\frac{\tilde{B}}{B_{\rm u}}=\frac{\tilde{\rho}}{\rho_{\rm u}}.$ (30)
-The total MHD stress-energy tensor may be summarized as follows
-$T^{\mu\nu}_{\rm MHD}=T^{\mu\nu}_{\rm p}+T^{\mu\nu}_{\rm f}=h\tilde{\rho}c^{2}u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P,$ (31)
-where
-$h=1+w+\sigma,\qquad P=\tilde{\rho}c^{2}\left(\frac{w}{3}+\frac{\sigma}{2}\right),$ (32)
-$w\equiv\frac{U_{\rm p}+P_{\rm p}}{\tilde{\rho}c^{2}}=\frac{3}{2}(\tilde{\gamma}-1),\quad\sigma\equiv\frac{\tilde{B}^{2}}{4\pi\tilde{\rho}c^{2}}=\frac{\tilde{\rho}}{\tilde{\rho}_{\rm u}}\,\sigma_{\rm u}.$ (33)
-Note that the stress-energy tensor of the MHD fluid depends on the strength parameter $a$ of the wave it interacts with.
-In particular, in the regular heating regime, Equation (9) implies that the dimensionless enthalpy $w$ tracks the local strength of the wave $a(\xi)$:
-$w\approx\frac{3}{2}a\qquad\left({\rm when~{}}\,\tilde{\omega}_{B}<\tilde{\omega}~{}{\rm and}~{}a\gg 1\right).$ (34)
+$\frac{\tilde{B}}{B_{\rm u}}=\frac{\tilde{\rho}}{\rho_{\rm u}}=\frac{\sigma}{\sigma_{\rm u}},\qquad\sigma\equiv\frac{\tilde{B}^{2}}{4\pi\tilde{\rho}c^{2}}.$ (33)
+The remaining component of the system is the plasma, which is energized by the radio wave.
+The plasma stress-energy tensor $T_{\rm p}^{\mu\nu}$ is derived in Appendix C for both possible regimes: stochastic heating and regular oscillation in the wave (Equations (C6) and (C8), respectively).
+The total MHD stress-energy tensor is
+$T^{\mu\nu}_{\rm MHD}=T^{\mu\nu}_{\rm p}+T^{\mu\nu}_{\rm f}.$ (34)
 Interaction of the wave with the MHD fluid obeys energy and momentum conservation,
-$\partial_{\nu}\left(T^{\mu\nu}_{\rm MHD}+T^{\mu\nu}_{\rm w}\right)=-\tilde{n}\dot{{\cal E}}_{e}u_{\rm D}^{\mu}.$ (35)
-The term $\tilde{n}\dot{{\cal E}}_{e}u^{\mu}$ is the radiative loss rate (per unit volume) of the plasma particles, which oscillate and gyrate in the wave; $\tilde{n}=\tilde{\rho}/m$ is the particle number density and ˙$\dot{{\cal E}}_{e}$ is the emitted power per particle, which is a Lorentz-invariant quantity.
-The power ˙$\dot{{\cal E}}_{e}$ can be calculated in the fluid frame $\tilde{\cal K}$. When $\tilde{\omega}_{B}<\tilde{\omega}$, it is given by
-$\dot{\gamma}_{e}\equiv\frac{\dot{{\cal E}}_{e}}{mc^{2}}\approx\frac{2}{3}\frac{r_{e}}{c}\tilde{\omega}^{2}a^{2}\tilde{\gamma}^{2}.$ (36)
-A similar expression (with the numerical coefficient $4/3$ instead of $2/3$) holds in the regime of stochastic heating, $\tilde{\omega}_{B}>\tilde{\omega}$ (B22).
-The strength parameter of the radio wave $a$ is invariant, and its frequency $\omega$ and electric field $E_{\rm wave}$ transform to the fluid frame as
-$\frac{\tilde{\omega}}{\omega}=\frac{\tilde{E}_{\rm wave}}{E_{\rm wave}}=\kappa^{-1},$ (37)
-where $\kappa$ is defined in Equation (13). Equation (36) may also be written in terms of the wave energy density $U$,
-$\dot{\gamma}_{e}\approx\frac{\sigma_{\rm T}\tilde{U}}{mc}\,\tilde{\gamma}^{2}=\frac{\sigma_{\rm T}U}{mc\,\kappa^{2}}\,\tilde{\gamma}^{2},$ (38)
-where $\sigma_{\rm T}=(8\pi/3)r_{e}^{2}$ is the Thomson cross section.
-4.2 MHD equations in coordinates $t,\xi$
+$\partial_{\nu}\left(T^{\mu\nu}_{\rm MHD}+T^{\mu\nu}_{\rm w}\right)=-n\dot{{\cal P}}_{e}^{\mu},$ (35)
+where $n=\rho/m$ is the particle number density, and $\dot{{\cal P}}_{e}^{\mu}=(\dot{{\cal E}}_{e}/c,\dot{\boldsymbol{{\cal P}}}_{e})$ is the average emission rate of four-momentum per particle, which is derived in Appendix D.
+It is given by Equations (D9), (D10) in the regular oscillation regime and by Equations (D19), (D20) in the stochastic heating regime.
+IV.2 MHD equations in coordinates $t,\xi$
 The conservation laws can also be stated using fluxes through the surfaces of $\xi=const$, which move with speed $c$. This will considerably simplify the analysis of wave-fluid interaction.
@@
 $\left.\frac{\partial\rho}{\partial t}\right|_{\xi}+\left.\frac{\partial F_{\rm m}}{\partial\xi}\right|_{t}=0.$ (37)
-Similarly, the flux of four-momentum ${\cal P}^{\mu}$ through a surface $\xi=const$ is $T^{\mu t}-T^{\mu z}$.
+Similarly, the flux of four-momentum ${\cal P}^{\mu}$ through a surface $\xi=const$ is $F^{\mu}=T^{\mu t}-T^{\mu z}$.
 Conservation of four-momentum (Equation 35) in the $t,\xi$-coordinates becomes [^7] This equation can also be derived formally by transforming tensor $T^{\mu\nu}$ to new coordinates: $(t,z)\rightarrow(t,\xi)$.
 The resulting $T^{t\xi}$ and $T^{z\xi}$ are the energy and momentum fluxes crossing $\xi=const$.
-$\left.\frac{\partial T^{\mu t}}{\partial t}\right|_{\xi}+\frac{\partial}{\partial\xi}\left(T^{\mu t}-T^{\mu z}\right)_{t}=-\tilde{n}\dot{{\cal E}}_{e}u_{\rm D}^{\mu}.$ (41)
-Note that the wave is static in $\xi$, so it creates no flux of energy or momentum across a surface of $\xi=const$:
-$T^{\mu t}_{\rm w}-T^{\mu z}_{\rm w}=U-U=0\qquad(\mu=t,z).$ (42)
-Only the MHD fluid contributes to the fluxes. Thus, the fluxes of energy and momentum through a surface of $\xi=const$ are
-$\frac{F_{\rm en}}{c}$ $=T^{tt}_{\rm MHD}-T^{tz}_{\rm MHD}=\frac{c^{2}\rho h}{\kappa}-P,$ (43)
-$F_{\rm mom}$ $=T^{zt}_{\rm MHD}-T^{zz}_{\rm MHD}=\frac{c^{2}\rho h}{\kappa}\beta_{\rm D}-P.$ (44)
-4.3 Conservation of ${\cal Q}_{\pm}$
-Instead of energy ${\cal E}$ and momentum ${\cal P}$ it will be helpful to use quantities
-${\cal Q}_{\pm}={\cal E}\pm c{\cal P}.$ (45)
-The fluxes of ${\cal Q}_{\pm}$ through a surface of $\xi=const$ are
-$F_{-}$ $=F_{\rm en}-cF_{\rm mom}=\frac{c^{3}\tilde{\rho}h}{\kappa^{2}},$ (46)
-$F_{+}$ $=F_{\rm en}+cF_{\rm mom}=c^{3}\tilde{\rho}h-2cP.$ (47)
-Conservation of ${\cal Q}_{\pm}$ is given by the sum/difference of equations describing energy and momentum conservation (Equation (41) with $\mu=t$ and $z$):
-$\left.\frac{\partial}{\partial t}\left(T^{tt}\pm T^{tz}\right)\right|_{\xi}+\frac{\partial}{\partial\xi}\left(\frac{F_{\pm}}{c}\right)_{t}=-\tilde{n}\dot{{\cal E}}_{e}(u_{\rm D}^{t}\pm u_{\rm D}^{z}).$ (48)
-Note that the stress-energy tensor of the radio wave disappears from the conservation law for ${\cal Q}_{-}$.
-This has a simple physical reason: the energy and momentum of the wave satisfy ${\cal E}_{\rm w}=c{\cal P}_{\rm w}$, so the wave carries ${\cal Q}_{-}=0$.
-There is no exchange of ${\cal Q}_{-}$ in the wave-fluid interaction: whenever the radio wave loses energy $\Delta{\cal E}$ to the MHD fluid, it also loses momentum $\Delta{\cal P}=\Delta{\cal E}/c$, so
-$\Delta{\cal Q}_{-}=\Delta{\cal E}-c\,\Delta{\cal P}=0.$ (49)
+$\left.\frac{\partial T^{\mu t}}{\partial t}\right|_{\xi}+\frac{\partial}{\partial\xi}\left(T^{\mu t}-T^{\mu z}\right)_{t}=-cn\dot{{\cal P}}_{e}^{\mu}.$ (38)
+Note that the wave is static in $\xi$, so it creates no flux of energy or momentum across a surface of $\xi=const$: [^8] This statement holds for waves propagating with speed $c$.
+The exact propagation speed $v_{\rm gr}<c$ (Appendix A) implies that the radio wave carries a non-zero energy flux in $\xi$: $F_{\rm w}=U(c-v_{\rm gr})$.
+In particular, in the fluid frame $\tilde{\cal K}$, $\tilde{F}_{\rm w}\sim c\tilde{U}\tilde{\omega}_{\rm p}^{2}/2\tilde{\omega}^{2}\sim\tilde{\rho}c^{3}a^{2}/2\tilde{\gamma}$.
+One may neglect $\tilde{F}_{\rm w}$ compared to flux $\tilde{F}$ created by the MHD fluid (effectively setting $v_{\rm gr}=c$) if $a^{2}/\tilde{\gamma}\ll h$, where $h$ is given in Equation (44) below.
+This condition turns out satisfied if the plasma has $\sigma_{\rm u}\gg 1$.
+$F^{\mu}_{\rm w}=T^{\mu t}_{\rm w}-T^{\mu z}_{\rm w}=U-U=0\qquad(\mu=t,z).$ (39)
+Then, only the MHD fluid contributes to the fluxes.
+IV.3 Conservation of $\cal Q$$^{\pm}$
+[…]
+${\cal Q}^{\pm}={\cal E}\pm c{\cal P}.$ (40)
+Conservation of $\cal Q$$^{\pm}$ is given by the sum/difference of equations stating energy and momentum conservation (Equation (38) with $\mu=t$ and $z$):
+$\frac{\partial}{\partial t}(T^{tt}+T^{tz})+\frac{\partial}{\partial\xi}(T^{tt}-T^{zz})$ $=-\dot{\cal Q}^{+},$ (41)
+$\frac{\partial}{\partial t}(T^{tt}-T^{tz})+\frac{\partial}{\partial\xi}(T^{tt}-2T^{tz}+T^{zz})$ $=-\dot{\cal Q}^{-},$ (42)
+where $\dot{\cal Q}^{\pm}=cn(\dot{{\cal P}}_{e}^{t}\pm\dot{{\cal P}}_{e}^{z})$.
+[…]
+$\Delta{\cal Q}^{-}=\Delta{\cal E}-c\,\Delta{\cal P}=0.$ (43)
 This observation significantly simplifies the problem.
-As long as the shape of the propagating radio wave $a(\xi)$ is approximately steady (i.e.
-the wave is not damped), the fluid dynamics in the packet is fully described by two equations that express conservation of mass and ${\cal Q}_{-}$:
+As long as the propagating radio wave packet $a(\xi)$ is approximately steady (i.e.
+the wave is not damped), the fluid dynamics in the packet is fully described by two equations that express conservation of mass and $\cal Q$$^{-}$.
+The remaining independent equation for $Q^{+}$ is needed if one aims to explicitly follow the gradual evolution of the wave packet $a(t,\xi)$.
+It is not needed for FRBs that are weakly affected by their interaction with the wind.
+Here, we will assume this condition to be satisfied, and later (in section X) will estimate the parameter space where strong FRB damping is expected.
+Another observation concerns the form of the plasma stress-energy tensor $T_{\rm p}^{\mu\nu}$ derived in Appendix C.
+In the stochastic heating regime, it takes the usual ideal-fluid form (Equation C8).
+By contrast, in the regular oscillation regime, $T_{\rm p}^{\mu\nu}$ looks different because of the first term in Equation (C6).
+However, this term is the same in all three components $T_{\rm p}^{tt}$, $T_{\rm p}^{tz}$, $T_{\rm p}^{zz}$ and drops out from Equation (42).
+Thus, effectively, only the second term in $T_{\rm p}^{\mu\nu}$ contributes to the $Q^{-}$ equation, i.e.
+the plasma stress-energy tensor can be replaced by $T_{\rm p}^{\mu\nu}=(1+a^{2})^{1/2}\tilde{\rho}c^{2}u_{\rm D}^{\mu}u_{\rm D}^{\nu}$.
+It has the ideal-fluid form with zero pressure and non-zero dimensionless enthalpy $w$ defined by $1+w=(1+a^{2})^{1/2}$.
+Recall that the MHD electromagnetic field also has $T_{\rm f}^{\mu\nu}$ of an ideal-fluid form (Equation 31), with effective pressure $\tilde{\rho}c^{2}\sigma/2$ and enthalpy $\tilde{\rho}c^{2}\sigma$.
+As a result, one can replace $T^{\mu\nu}$ in Equation (42) with the effective $T_{\rm MHD}^{\mu\nu}=\tilde{\rho}c^{2}hu_{\rm D}^{\mu}u_{\rm D}^{\nu}+P\eta^{\mu\nu}$ with the following definitions of $h$ and $P$:
+$h=1+w+\sigma,$ (44)
+$w=\sqrt{1+a^{2}}-1,\qquad\frac{P}{\tilde{\rho}c^{2}}=\frac{\sigma}{2}\qquad({\rm regular})$ (45a)
+$w=\frac{3}{2}(\tilde{\gamma}-1),\qquad\frac{P}{\tilde{\rho}c^{2}}=\frac{\sigma}{2}+\frac{w}{3}\quad\;({\rm stochastic})$ (45b)
+In the regular oscillation regime, the plasma makes no contribution to pressure $P$; the oscillations only increase the effective inertial mass of each particle by the factor $\tilde{\gamma}=\sqrt{1+a^{2}}$.
+Unlike stochastic heating, coherent oscillations are not true heating.
+However, it is convenient to define effective enthalpy $w$ in both regimes, so that they have the common form of $T_{\rm MHD}^{\mu\nu}$.
+The equations of mass and $Q^{-}$ conservation can now be stated as follows
 $\partial_{t}\rho$ $=-\partial_{\xi}\left(\frac{\tilde{\rho}}{\kappa}\right),$ (46)
-$\partial_{t}\left(\frac{\rho h}{\kappa}-\frac{P}{c^{2}}\right)$ $=-\partial_{\xi}\left(\frac{\tilde{\rho}h}{\kappa^{2}}\right)-\frac{\tilde{\rho}}{\kappa}\dot{\gamma}_{e}.$ (51)
-The radio wave does not contribute to these conservation laws, because it carries no mass or ${\cal Q}_{-}$.
-The wave strength parameter $a(\xi)$ still enters the ${\cal Q}_{-}$ equation through $h=1+w+\sigma$ (since $a$ controls the fluid enthalpy $w$) and through $\dot{\gamma}_{e}\propto a^{2}$.
-As long as $a(\xi)$ is steady, it can be treated as a given function in the wave-fluid interaction problem, so Equations (50) and (51) form a closed set for two unknowns $\rho(t,\xi)$ and $\kappa(t,\xi)$.
-For a given radio wave packet, the equations can be solved in the domain $0<\xi<T$ with given initial conditions $\rho(0,\xi)$ and $\kappa(0,\xi)$, and boundary conditions $\rho(t,0)=\rho_{\rm u}$ and $\kappa(t,0)=\kappa_{\rm u}$.
-Equations (50) and (51) allow one to focus on the fluid dynamics in the wave packet without considering the evolution of the radio wave itself, assuming that the evolution of $a(\xi)$ is negligible on timescales of interest.
-The consistency of this assumption can be checked later using the remaining independent equation that states conservation of ${\cal Q}_{+}$ and governs the evolution of $a(t,\xi)$.
-A simpler consistency check can be done by comparing the energy exchanged in the wave-plasma interaction with the wave energy (Section 10).
-Note also that Equations (50) and (51) are Lorentz-invariant: they have the same form in any frame boosted along $z$ [^7] Lorentz transformation of initial conditions is not, however, straightforward, since hypersurfaces of simultaneity ($t=const$) are different in different frames.
+$\partial_{t}\left(\frac{\rho h}{\kappa}-\frac{P}{c^{2}}\right)$ $=-\partial_{\xi}\left(\frac{\tilde{\rho}h}{\kappa^{2}}\right)-\frac{n\dot{{\cal Q}}^{-}_{e}}{c^{2}}.$ (47)
+The term $\dot{{\cal Q}}^{-}_{e}=c(\dot{{\cal P}}_{e}^{t}-\dot{{\cal P}}_{e}^{z})$ describes radiative losses per particle.
+It is determined by Equations (D9), (D10) and Equations (D19), (D20) in the regimes of regular oscillation and stochastic heating:
+$\!\!\!\!\!\!\dot{\cal Q}^{-}_{e}\!=\!\frac{c\sigma_{\rm T}U}{\gamma_{\rm D}\kappa^{3}}\!\!\times\!\!\left\{\begin{array}[]{lr}\vskip 2.84526pt1+a^{2}&\!\!\!\!\!{\rm regular}\\ \displaystyle{\frac{5}{2}\tilde{\gamma}^{2}}&\!\!\!\!{\rm stochastic}\end{array}\!\right\}\!\approx\!\frac{c\sigma_{\rm T}U}{\gamma_{\rm D}\kappa^{3}}(1+w)^{2}.$ (50)
+where $\sigma_{\rm T}=(8\pi/3)r_{e}^{2}$ is the Thomson cross section.
+The radio wave does not contribute to Equations (46) and (47), because it carries no mass and negligible $\cal Q$$^{-}$.
+The wave strength parameter $a(\xi)$ still enters the $\cal Q$$^{-}$ equation through $h=1+w+\sigma$ (since $a$ controls the fluid enthalpy $w$) and through $\dot{{\cal Q}}_{e}^{-}\propto U\propto a^{2}$.
+As long as $a(\xi)$ is steady (not damped), it can be treated as a given function in the wave-fluid interaction problem, and Equations (46) and (47) form a closed set for two unknowns $\rho(t,\xi)$ and $\kappa(t,\xi)$.
+The equations can be solved in the domain $0<\xi<T$ (i.e.
+inside the radio wave packet) with given initial conditions $\rho(0,\xi)$ and $\kappa(0,\xi)$, and boundary conditions $\rho(t,0)=\rho_{\rm u}$ and $\kappa(t,0)=\kappa_{\rm u}$.
+Equations (46) and (47) are Lorentz-invariant: they have the same form in any frame boosted along zz.
+[^9] Lorentz transformation of initial conditions is not, however, trivial, since hypersurfaces of simultaneity ($t=const$) are different in different frames.
 They can be solved e.g.
 in the frame where the upstream plasma is at rest: $\kappa_{\rm u}=1$ and $\rho_{\rm u}=\tilde{\rho}_{\rm u}$.
-The problem can also be solved in other frames where $\kappa_{\rm u}\neq 1$ and the upstream plasma has speed $\beta_{\rm u}=(\kappa_{\rm u}^{2}-1)/(\kappa_{\rm u}^{2}+1)$, Lorentz factor $\gamma_{\rm u}=(\kappa_{\rm u}^{2}+1)/2\kappa_{\rm u}$, and proper density $\tilde{\rho}_{\rm u}=\rho_{\rm u}/\gamma_{\rm u}$.
-5 Steady compression front
+The problem can also be solved in other frames where $\kappa_{\rm u}\neq 1$ and the upstream plasma has speed $\beta_{\rm u}=(\kappa_{\rm u}^{2}-1)/(\kappa_{\rm u}^{2}+1)$, Lorentz factor $\gamma_{\rm u}=(\kappa_{\rm u}^{2}+1)/2\kappa_{\rm u}$, and density $\rho_{\rm u}=\gamma_{\rm u}\tilde{\rho}_{\rm u}$.
+V Steady compression front
 The pattern of fluid motion inside the propagating wave packet will be called below the “front” for brevity.
@@
 Here, we first show that there is a unique steady flow pattern, where all quantities depend on $\xi$ only, with no evolution with $t$.
-As verified in section 6, the steady state is an attractor for the front evolution.
-It is reached under two conditions: (1) the front propagation time is long enough for the flow pattern to relax to the steady state, and (2) the propagation time is short enough to neglect damping of the wave, so $a(\xi)$ can be treated as constant in time.
+As verified in section VI, the steady state is an attractor for the front evolution.
+It is reached under two conditions: (1) the front propagation time is long enough for the flow pattern to relax to the steady state, and (2) the propagation time is short enough to neglect damping of the radio wave, so $a(\xi)$ can be treated as constant in time.
 The steady front solution is found by setting $\left.\partial_{t}\right|_{\xi}=0$ and $\partial_{\xi}=d/d\xi$ in the MHD equations. Then, Equations (46) and (47) give
 $\frac{d}{d\xi}\left(\frac{\tilde{\rho}}{\kappa}\right)$ $=0\quad\Rightarrow\quad\tilde{\rho}=\frac{\kappa}{\kappa_{\rm u}}\tilde{\rho}_{\rm u},$ (51)
-$\frac{d}{d\xi}\left(\frac{h}{\kappa}\right)$ $=-\dot{\gamma}_{e}.$ (53)
+$\frac{d}{d\xi}\left(\frac{h}{\kappa}\right)$ $=-\frac{\kappa\gamma_{\rm D}\dot{{\cal Q}}^{-}_{e}}{mc^{2}}.$ (52)
 The corresponding lab-frame density $\rho=\gamma_{\rm D}\tilde{\rho}$ can be expressed using
@@
 $\frac{\rho}{\rho_{\rm u}}=\frac{1-\beta_{\rm u}}{1-\beta}=\frac{\kappa^{2}+1}{\kappa_{\rm u}^{2}+1}.$ (54)
-The solution $\tilde{\rho}/\rho_{\rm u}=\kappa/\kappa_{\rm u}$ also implies $\sigma=\sigma_{\rm u}\kappa/\kappa_{\rm u}$ and $d/d\xi(h/\kappa)=d/d\xi[(1+w)/\kappa]$. Thus, Equation (53) becomes
-$\frac{d}{d\xi}\left(\frac{1+w}{\kappa}\right)=-\dot{\gamma}_{e}=-\frac{\sigma_{\rm T}U}{mc}\,\frac{\tilde{\gamma}^{2}}{\kappa^{2}}.$ (56)
+The solution $\tilde{\rho}/\tilde{\rho}_{\rm u}=\kappa/\kappa_{\rm u}$ also implies $\sigma=\sigma_{\rm u}\kappa/\kappa_{\rm u}$ and $d/d\xi(h/\kappa)=d/d\xi[(1+w)/\kappa]$.
+As a result, Equation (52) becomes
+$\frac{d}{d\xi}\left(\frac{1+w}{\kappa}\right)=-\frac{\sigma_{\rm T}U(1+w)^{2}}{mc\,\kappa^{2}}.$ (55)
 Note that the fluid magnetization $\sigma$ has dropped out.
-Remarkably, the fluid motion and compression in a steady front are independent of $\sigma_{\rm u}$.
-In particular, in the case of negligible radiative losses, we find
-$\kappa=(1+w)\kappa_{\rm u}\qquad({\rm when}~{}\,\dot{\gamma}_{e}=0),$ (57)
-where we used the cold upstream condition $w_{\rm u}=0$.
-One can also solve for $\kappa(\xi)$ without neglecting radiative losses.
-In particular, consider the simpler regime of regular heating ($\tilde{\omega}_{B}<\tilde{\omega}_{B}$).
-We are interested in waves with $a\gg 1$, which heat the plasma to $w\approx(3/2)\tilde{\gamma}\gg 1$.
-Equation (56) can be integrated for $(1+w)/\kappa$, and we find
-$\kappa=(1+w)\kappa_{\rm u}(1+q),$ (58)
+Remarkably, fluid motion and compression in a steady front are unaffected by the magnetic field, regardless of how strong it is.
+Integrating Equation (55) for $(1+w)/\kappa$ and using the cold upstream condition $w_{\rm u}=0$, we find
+$\kappa=(1+q)(1+w)\kappa_{\rm u},$ (56)
 where
-[…]
-The importance of radiative losses is controlled by the (Lorentz-invariant) dimensionless parameter $q$, which is determined by the column energy density of the wave packet, $\Sigma$.
-Equation (57) holds when $q\ll 1$.
-6 Relaxation to the steady state
-A simple argument shows that the timescale to approach the steady state is long if the plasma magnetization $\sigma_{\rm u}$ is large.
-Consider the vacuum limit of $\rho_{\rm u}\rightarrow 0$ (which corresponds to $\sigma_{\rm u}\rightarrow\infty$).
+$q(\xi)=\frac{\sigma_{\rm T}\Sigma(\xi)}{mc^{2}\kappa_{\rm u}},\qquad\Sigma(\xi)\equiv\int_{0}^{\xi}U(\xi^{\prime})c\,d\xi^{\prime}.$ (57)
+The effect of radiative losses on the fluid motion is controlled by the (Lorentz-invariant) dimensionless parameter $q$, which is determined by the column energy density of the radio wave packet, $\Sigma$.
+Losses weakly affect the front when $q\ll 1$.
+In the regular oscillation regime with $q\ll 1$, the obtained steady-state solution for the compression front simplifies to
+$\kappa=\kappa_{\rm u}\sqrt{1+a^{2}}\qquad\mathrm{(regular\;\&\;non\mathrm{-}radiative)}.$ (58)
+In this limit, the solution agrees with the result of Sobacchi et al.
+(2024).
+This agreement is remarkable, since Sobacchi et al.
+(2024) used a different approach to the steady-state compression front: they studied plasma in the frame moving with the group speed of a nearly monochromatic wave packet, and employed an expansion in Fourier harmonics, assuming that the upstream plasma ahead of the packet has $\tilde{\omega}_{B}<\tilde{\omega}/a$.
+VI Relaxation to the steady state
+A simple argument shows that the approach to the steady state takes the front a long propagation time if the plasma magnetization $\sigma_{\rm u}$ is large.
+Consider the vacuum limit of $\rho_{\rm u}\rightarrow 0$, which corresponds to $\sigma_{\rm u}\rightarrow\infty$.
+[^10] Recall that we consider radio waves with frequencies exceeding the plasma Larmor frequency.
+In the opposite case, they would be MHD waves (where magnetized particles move with $\beta_{\rm D}=E/B$), and $\rho_{\rm u}\rightarrow 0$ would correspond to force-free electrodynamics.
 Then, the background field must remain unchanged from its upstream state: in a vacuum, the electromagnetic wave goes through the background field without interaction, so no $E_{\rm bg}$ is generated.
 Thus, at $\sigma_{\rm u}\rightarrow\infty$ the magnetized plasma must remain static and will not be compressed by the radio wave packet.
-The steady-state front with large $C$ still provides a consistent solution, but relaxing to it takes an infinite time, so it becomes irrelevant at $\sigma_{\rm u}\rightarrow\infty$.
-A similar conclusion is reached by considering energy.
+The steady-state front with a large compression factor $C=\rho/\rho_{\rm u}$ still provides a consistent solution, but relaxing to it takes an infinite time, so it becomes irrelevant at $\sigma_{\rm u}\rightarrow\infty$.
+The same conclusion is reached by considering energy.
 Compression of the magnetically dominated plasma requires deposition of energy $\propto\sigma_{\rm u}$, much greater than the energy of plasma particles that interact with the wave.
@@
 It increases the effective mass per particle from $m$ to $\sigma_{\rm u}m$, which makes it heavy and impossible to accelerate in the limit of $\sigma_{\rm u}\rightarrow\infty$.
-An analogy can be drawn with collision of a photon of energy ${\cal E}_{\rm ph}$ with a particle of mass $\sigma m$.
-Using the frame where the particle is initially at rest, one finds that the particle gains speed $\beta\approx{\cal E}_{\rm ph}/\sigma mc^{2}\ll 1$ if $\sigma\gg{\cal E}_{\rm ph}/mc^{2}$.
-It is worth looking in more detail at a similar collision problem formulated for the wave-plasma system.
-6.1 Wave-plasma collision
-Suppose a radio wave packet of energy ${\cal E}$ passes through an initially static, cold slab of plasma with magnetization $\sigma_{\rm u}$.
-Some energy $\Delta{\cal E}$ and momentum $\Delta{\cal E}/c$ are lost by the radio wave and received by the slab.
-As a result, the slab is accelerated to speed $\beta_{\rm D}$ and compressed by the factor $1-\beta_{\rm D}$ (cf.
-Equation (55)).
-Compression increases the slab magnetic field from initial $B_{\rm u}$ to $B=B_{\rm u}/(1-\beta_{\rm D})$.
-In the simplified collision picture, the slab is treated as a single body characterized by its volume $V$, particle number $N$, inertial mass of the plasma $M=Nm(1+w)$ (increased from $M_{\rm u}=Nm$ by deposited heat), magnetic field $B$, and electric field $E=\beta_{\rm D}B$.
-Energy conservation gives
-$M_{\rm u}c^{2}+V_{\rm u}\frac{B_{\rm u}^{2}}{8\pi}+\Delta{\cal E}=\gamma_{\rm D}Mc^{2}+V\frac{B^{2}+E^{2}}{8\pi}+{\cal E}_{\rm rad},$ (60)
-where we included possible radiative losses of the plasma ${\cal E}_{\rm rad}$.
-The radiated momentum ${\cal P}^{\prime}_{\rm rad}=0$ in the fluid rest frame implies ${\cal P}_{\rm rad}=\beta_{\rm D}{\cal E}_{\rm rad}/c$ in the lab frame.
-Momentum conservation equation (multiplied by $c$) is
-$\Delta{\cal E}=\gamma_{\rm D}\beta_{\rm D}Mc^{2}+V\frac{EB}{4\pi}+\beta_{\rm D}{\cal E}_{\rm rad}.$ (61)
-Subtracting the momentum equation from the energy equation one can exclude $\Delta{\cal E}$:
-$M_{\rm u}c^{2}+V_{\rm u}\frac{B_{\rm u}^{2}}{8\pi}=\frac{Mc^{2}}{\kappa}+V\frac{(B-E)^{2}}{8\pi}+(1-\beta_{\rm D}){\cal E}_{\rm rad}.$ (62)
-We now substitute $E=\beta_{\rm D}B$, $B=B_{\rm u}/(1-\beta_{\rm D})$, $VB=V_{\rm u}B_{\rm u}$, and $M=(1+w)M_{\rm u}$, and obtain
-$1=\frac{1+w}{\kappa}-\frac{\beta_{\rm D}\sigma_{\rm u}}{2}+(1-\beta_{\rm D})\epsilon_{\rm rad},$ (63)
-where
-$\sigma_{\rm u}\equiv\frac{V_{\rm u}B_{\rm u}^{2}}{4\pi M_{\rm u}c^{2}},\qquad\epsilon_{\rm rad}\equiv\frac{{\cal E}_{\rm rad}}{M_{\rm u}c^{2}}.$ (64)
-Radiative losses can be described by the effective scattering cross section $\sigma_{\rm sc}$ per electron, with $\sigma_{\rm sc}\sim w^{2}\sigma_{\rm T}$ within a factor of $\sim 2$ (B22). It implies
-$\epsilon_{\rm rad}\sim\frac{\sigma_{\rm T}w^{2}\Sigma}{mc^{2}},$ (65)
-where $\Sigma$ is the column energy density of the radio wave packet.
-Using the identity $1-\beta_{\rm D}=2/(\kappa^{2}+1)$ one can rewrite Equation (63) as a cubic equation for $\kappa$.
-Its solution simplifies in two limits: $\kappa\gg 1$ ($\beta_{\rm D}\approx 1$) and $\beta_{\rm D}\ll 1$ ($\kappa\approx 1$).
-Assuming $\sigma_{\rm u}\gg 1$, we find
-$\kappa-1\approx\left\{\begin{array}[]{lr}\vspace*{1mm}\displaystyle{\frac{2}{\sigma_{\rm u}}(w+\epsilon_{\rm rad})}&{\rm when}\;\displaystyle{w+\epsilon_{\rm rad}\ll\sigma_{\rm u}}\\ \displaystyle{\frac{w}{\sigma_{\rm u}}+\sqrt{\frac{\epsilon_{\rm rad}}{\sigma_{\rm u}}}}&{\rm when}\;\displaystyle{w+\epsilon_{\rm rad}\gg\sigma_{\rm u}}\end{array}\right.$ (66)
-It also gives the proper compression factor $\tilde{C}\approx\kappa$.
-Once $\kappa$ is known, we can also find the energy $\Delta{\cal E}$ taken from the radio wave packet. Equation (61) gives
-$\frac{\Delta{\cal E}}{M_{\rm u}c^{2}}=\beta_{\rm D}\left[\gamma_{\rm D}(1+w+\sigma_{\rm u}\kappa)+\epsilon_{\rm rad}\right],$ (67)
-and using Equation (66) we find
-$\frac{\Delta{\cal E}}{M_{\rm u}c^{2}}\approx\left\{\begin{array}[]{lr}\vspace*{1mm}\displaystyle{2(w+\epsilon_{\rm rad})},&\displaystyle{w+\epsilon_{\rm rad}\ll\sigma_{\rm u}}\\ \displaystyle{\frac{w^{2}}{\sigma_{\rm u}}+\frac{3}{2}w\sqrt{\frac{\epsilon_{\rm rad}}{\sigma_{\rm u}}}+\frac{3}{2}\epsilon_{\rm rad}},&\displaystyle{w+\epsilon_{\rm rad}\gg\sigma_{\rm u}}\end{array}\right.$ (68)
-The upstream plasma is composed of many slabs.
-One after another, they fill the propagating packet and exit behind it, and one might conclude that Equation (66) approximately describes $\kappa$ in a steady propagating front.
-In fact, Equation (66) is very far from the correct MHD steady-state solution given in section 5.
-When viewed in the upstream rest frame, a non-radiative steady front should have $\kappa=1+w$ regardless of $\sigma_{\rm u}$ (Equation 57).
-In contrast, the collision model described above yields $\kappa\approx 1+(2w/\sigma_{\rm u})\approx 1$ when $\sigma_{\rm u}\gg w$.
-The simplified collision model (Equation 66) works well at the beginning of packet propagation through the plasma; however, it fails to describe the gradual relaxation of the compression front to the correct steady state after many interactions with many plasma slabs.
-When $\sigma_{\rm u}\gg 1$, this relaxation takes a long time and occurs due to a slow growth of magnetic pressure in the front, as demonstrated below.
-The grown magnetic pressure drops out from the final steady pattern of the MHD flow because it occurs with a zero gradient of $\sigma/\kappa$ (section 5).
-This behavior is not captured by the simplified collision model, which treats each plasma slab passing through the packet as a single body and does not consider any gradients.
-By contrast, in full MHD, the gradients across the front play an essential role: they endow the propagating front with a “memory” of its earlier interactions with the fluid that has been disposed behind the front.
+An analogy can be drawn with absorption of a photon with energy ${\cal E}_{\rm ph}$ by a particle of mass $\sigma m$.
+Using the frame where the particle is initially at rest, one finds that the particle gains speed $\beta\approx{\cal E}_{\rm ph}/\sigma mc^{2}\ll 1$ if $\sigma\gg{\cal E}_{\rm ph}/mc^{2}$ (and ${\cal E}_{\rm ph}$ transforms mainly into the particle’s internal energy).
+Similarly, one might expect that the interaction of a radio wave packet with the MHD fluid gives it a small bulk velocity,
+$\beta_{\rm D}\sim\frac{w}{\sigma_{\rm u}}\ll 1,$ (59)
+when the interaction generates fluid enthalpy $w\ll\sigma_{\rm u}$.
+This would mean that energy deposition by the radio wave heats the fluid without significant bulk acceleration or compression.
+[^11] A toy problem of wave-plasma collision may be formulated treating the magnetized fluid as a single body carrying energy and momentum.
+A consistent formulation, however, requires some care: the MHD fluid would need to be confined by walls, which carry stress and energy to be included in the conservation laws.
+This estimate clearly differs from the steady-front solution (Equation 56), which gives $\kappa\gg 1$ (i.e.
+$\beta_{\rm D}\approx 1$) at $w\gg 1$ regardless of $\sigma_{\rm u}$.
+The estimate $\beta_{\rm D}\sim w/\sigma_{\rm u}$ may be reasonable at the beginning of the packet propagation through the plasma with $\sigma_{\rm u}\gg w$; however, it fails to describe the gradual relaxation of the compression front to the steady state.
+When $\sigma_{\rm u}\gg 1$, this relaxation takes a long time and involves a slow growth of the compression factor and the magnetic pressure in the front, as demonstrated below.
+Pressure gradients across the front play an essential role: they endow the propagating front with a “memory” of its earlier interactions with the fluid that has been disposed behind the front.
 The accumulating memory enables the front relaxation to the steady state.
-Below we calculate this relaxation by solving the full time-dependent MHD problem described by Equations (50) and (51).
-We solve it first in the non-radiative regime (neglecting $\dot{\gamma}_{e}$ in Equation (51)) and then in the radiative regime.
-In the remainder of this section, we assume regular heating ($\tilde{\omega}_{B}>\tilde{\omega}$).
-A similar relaxation trend holds with stochastic heating ($\tilde{\omega}_{B}<\tilde{\omega}$); however, this regime gives a huge enthalpy $w$ and relaxation takes a very long time, so steady fronts with stochastic heating are not expected in FRBs (section 9).
-6.2 Evolution of a non-radiative compression front
+Below we demonstrate this relaxation by solving the full time-dependent MHD problem described by Equations (46) and (47).
+We solve it first in the non-radiative regime (neglecting $\dot{{\cal Q}}_{e}^{-}$ in Equation (47)) and then in the radiative regime.
+In the remainder of this section, we assume regular oscillations of particles in the radio wave.
+A similar relaxation trend holds with stochastic heating; however, this regime gives a huge enthalpy $w$ and relaxation would take a very long time, so steady fronts with stochastic heating do not form in FRBs (section IX).
+VI.1 Evolution of a non-radiative compression front
 Before solving numerically the time-depending MHD equations, we make a heuristic estimate of the timescale $t_{\star}$ for the front relaxation toward the steady state.
 First, note the time it takes a fluid element to cross the entire wave packet $0<\xi<T$:
-[…]
-In particular, in a steady state, $\kappa=\kappa_{\star}\approx 1+(3/2)a$. Then, for a packet with $a_{\max}\gg 1$, one finds
-[…]
-Hereafter, star indicates the steady state.
+$t_{\rm cross}=\int_{0}^{T}\frac{d\xi}{1-\beta_{\rm D}}=\int_{0}^{T}\frac{\kappa^{2}+1}{2}\,d\xi.$ (60)
+In particular, in a steady state and in the regular oscillation regime, $\kappa^{2}=(1+a^{2})\kappa_{\rm u}^{2}$. Then, for a packet with $a_{\max}\gg 1$, one finds
+$t_{\rm cross}^{\star}\approx\int_{0}^{T}[a^{2}(\xi)\,\kappa_{\rm u}^{2}+1]\frac{d\xi}{2}=\left(\frac{3}{8}a_{\max}^{2}\kappa_{\rm u}^{2}+1\right)\frac{T}{2},$ (61)
+where we assumed a wave packet $a(\xi)$ of the form given in Equation (7). Hereafter, star indicates a steady state.
 Now consider the momentum per particle carried by the MHD fluid inside the wave packet,
 ${\cal P}_{\rm MHD}=\frac{T^{tz}_{\rm MHD}}{nc}=h\gamma_{\rm D}\beta_{\rm D}mc.$ (62)
-It grows when the fluid receives momentum from the wave.
-Recall that the wave interacts with the plasma particles only, and the plasma shares the received momentum with the background field (through gyration).
+It grows when the fluid receives momentum from the radio wave.
+Recall that the wave directly interacts only with the plasma particles, and the plasma shares the received momentum with the background electromagnetic field (through gyration about $\boldsymbol{B}$).
 Therefore, the budget for the wave-fluid momentum exchange (per particle crossing the packet) is set by $T^{tz}_{\rm p}$ rather than $T^{tz}_{\rm MHD}$:
@@
 where $h=1+w+\sigma$.
-Let us now specialize to the frame where the upstream fluid is at rest ($\kappa_{\rm u}=1$), and the steady state has $\kappa=\kappa_{\star}=1+w$ and $\sigma=\kappa_{\star}\sigma_{\rm u}=(1+w)\sigma_{\rm u}$.
-In this frame, the expected timescale to approach the steady state is
-$t_{\star}\sim\frac{1+w+\kappa_{\star}\sigma_{\rm u}}{1+w}\,t_{\rm cross}^{\star}=(1+\sigma_{\rm u})t_{\rm cross}^{\star}\quad(\kappa_{\rm u}=1).$ (74)
+Let us specialize to the frame where the upstream fluid is at rest ($\kappa_{\rm u}=1$).
+The steady front has $\kappa=\kappa_{\star}=1+w$ and $\sigma=\kappa_{\star}\sigma_{\rm u}=(1+w)\sigma_{\rm u}$.
+In this frame,
+$t_{\rm cross}^{\star}\approx\frac{3}{16}a_{\max}^{2}T\qquad(a_{\max}\gg 1),$ (65)
+and the expected timescale to approach the steady state is
+$t_{\star}\sim\frac{1+w+\kappa_{\star}\sigma_{\rm u}}{1+w}\,t_{\rm cross}^{\star}=(1+\sigma_{\rm u})t_{\rm cross}^{\star}.$ (66)
 Figure 4: Top: relaxation of the non-radiative compression front $C(\xi)$ toward the steady-state $C_{\star}(\xi)$ (red curve, Equation (67)) from an initial state with a uniform static plasma, $C=1$ and $\kappa=1$.
-Black curves show snapshots of the front, with time $t$ indicated next to each curve (in units of packet duration $T$).
-The wave packet (Equation 5) has $a_{\max}=(4/3)\sqrt{7}$, equivalent to $w_{\max}=4$.
+Black curves show snapshots of the front at times $t$ indicated next to each curve in units of packet duration $T$.
+The wave packet (Equation 7) has $a_{\max}=\sqrt{24}$, which gives $w_{\max}=4$.
 The upstream magnetization is $\sigma_{\rm u}=10$.
-Bottom: evolution of $\kappa$ from the initial $\kappa=1$ toward $\kappa_{\star}$ in the same simulation.
-This simple estimate agrees with our numerical solutions for the front evolution.
+Bottom: evolution of $\kappa$ (black) from the initial $\kappa=1$ toward $\kappa_{\star}$ (red) in the same simulation.
+This simple estimate agrees with our numerical solution for the front evolution.
 We find the evolution by solving Equations (46) and (47) for $\kappa(t,\xi)$ and $\rho(t,\xi)$ in a given wave packet $a(\xi)$ propagating in a background plasma with a given upstream magnetization $\sigma_{\rm u}$.
-The fluid enthalpy inside the packet, $w\equiv(U_{\rm p}+P_{\rm p})/\tilde{\rho}c^{2}$, is determined by Equations (9) and (22):
-$w(\xi)=\frac{3}{2}\left[\sqrt{a^{2}(\xi)+1}-1\right].$ (75)
-In sample models, we use the packet described by Equation (5) with $a_{\max}=(4/3)\sqrt{7}$ which corresponds to $w_{\max}=4$.
-As an initial state at $t=0$, we take a static plasma, $\kappa(\xi)=1$ and $\rho(\xi)=\rho_{\rm u}$, and follow the evolution of the flow pattern inside the propagating wave packet until it relaxes to the steady-state solution.
-In the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), the steady-state solution is
-$\kappa_{\star}=1+w,\quad\rho_{\star}=\frac{\kappa_{\star}^{2}+1}{2}\rho_{\rm u}\qquad(\kappa_{\rm u}=1).$ (76)
-The model with $\sigma_{\rm u}=10$ is shown in Figure 3.
-At the beginning of the simulation we observe that $\kappa(\xi)$ starts to grow from unity and develops a peaked profile — the MHD fluid is accelerated in the first half of the packet and decelerated in the second half, before exiting behind the packet.
-Over time, $\kappa(\xi)$ approaches $\kappa_{\star}(\xi)$, which is an attractor for the flow pattern.
+The fluid enthalpy inside the packet is given by Equation (45a).
+In sample models, we use the packet described by Equation (7) with $a_{\max}=\sqrt{24}$, which corresponds to $w_{\max}=4$.
+As an initial state at $t=0$, we take a static plasma, $\kappa(\xi)=1$ and $\rho(\xi)=\rho_{\rm u}$, and follow the evolution of the flow pattern in the propagating wave packet until it relaxes to a steady-state.
+In the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), the expected steady-state solution is
+$\kappa_{\star}=\sqrt{1+a^{2}},\qquad C_{\star}\equiv\frac{\rho_{\star}}{\rho_{\rm u}}=\frac{\kappa_{\star}^{2}+1}{2}=1+\frac{a^{2}}{2}.$ (67)
+The sample model with $\sigma_{\rm u}=10$ is shown in Figure 4.
+At the beginning of the simulation we observe that $\kappa(\xi)$ starts to grow from unity and develops a peaked profile — the MHD fluid is accelerated in the leading part of the packet (at $\xi<\xi_{\star}$) and decelerated in the later part (at $\xi>\xi_{\star}$), before exiting behind the packet.
+The compression factor in the acceleration part is well described by the steady-state solution:
+$\kappa\approx\kappa_{\star}\;\;{\rm and}\;\;C\equiv\frac{\rho}{\rho_{\rm u}}\approx C_{\star}\quad{\rm at~}\xi<\xi_{\star}.$ (68)
+The maximum compression factor is reached at $\xi_{\star}$:
+$C_{\max}=C(\xi_{\star})\approx C_{\star}(\xi_{\star}).$ (69)
+The peak $C_{\max}$ gradually grows with time as its location $\xi_{\star}$ shifts toward the middle of the wave packet.
+Figure 5: Relation between $\kappa=\gamma_{\rm D}(1+\beta_{\rm D})$ and $C=\rho/\rho_{\rm u}$ (here measured at the peak inside the compression front) during relaxation toward the final steady state.
+Black curve shows the simulation result (the model with $\sigma_{\rm u}=10$), and red dashed curve shows the relation $C=(\kappa^{2}+1)/2$ (equivalent to $\tilde{\rho}/\kappa=\rho_{\rm u}$ and uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)$).
+Fluid density in the front closely tracks the steady-state solution $\tilde{\rho}=\kappa\rho_{\rm u}$ defined for an instantaneous profile of $\kappa(\xi)$, which slowly grows during relaxation.
+The correlated growth of $\kappa$ and $C$ stops when they reach $\kappa_{\max}^{\star}=5$ and $C_{\max}^{\star}=13$.
+Then, the fluxes of mass and $Q^{-}$ both become exactly uniform across the front, so a true steady state is achieved.
+When $\xi_{\star}$ reaches the middle of the wave packet, the entire compression front $\kappa(\xi)$ approaches $\kappa_{\star}(\xi)=\sqrt{1+a^{2}}$, which is an attractor for the flow pattern.
 In particular, the maximum of the $\kappa$-profile, $\kappa_{\max}$, eventually approaches the expected $\kappa_{\max}^{\star}=1+w_{\max}=5$.
-The maximum compression factor $C_{\max}=\rho_{\max}/\rho$ approaches
-$C_{\max}^{\star}=\frac{1}{2}\left[(1+w_{\max})^{2}+1\right].$ (77)
-The evolution of compression peak $C_{\max}(t)$ provides a simple way to quantify the front relaxation toward the steady-state.
+The maximum compression factor $C_{\max}=\rho_{\max}/\rho_{\rm u}$ approaches $C_{\max}^{\star}=1+a_{\max}^{2}/2=13$.
+We also observe in the simulation that the gradual evolution of the front proceeds through a sequence of quasi-steady states: each snapshot of the front shows a nearly uniform mass flux in $\xi$: $F_{\rm m}=\rho(c-v_{\rm D})\approx const$, so $\rho(1-\beta_{\rm D})\approx\rho_{\rm u}$ and $C\approx(1-\beta_{\rm D})^{-1}$.
+Thus, the relation $C\approx(\kappa^{2}+1)/2$ is sustained during the evolution, i.e.
+$\kappa$ and $C$ grow in a correlated way (Figure 5).
+Evolution of the compression peak $C_{\max}(t)$ provides a simple way to quantify the front relaxation toward the final, exact steady-state.
 The evolution of $C_{\max}$ is shown in Figure 6 for $\sigma_{\rm u}=3$, 10, 20.
 One can see that the timescale $t_{\star}$ for approaching the steady-state $C_{\max}^{\star}=13$ increases with $\sigma_{\rm u}$ in agreement with Equation (66).
-The slow evolution occurs through a sequence of quasi-steady states with a nearly uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)\approx const$.
-It implies a correlated growth of $\kappa$ and $C$: the relation $C\approx(\kappa^{2}+1)/2$ is sustained during the evolution (Figure 5).
 Figure 6: Evolution of the compression peak $C_{\max}$ from the initial $C_{\max}=1$ to the steady-state $C_{\max}^{\star}=13$ (indicated by the horizontal dashed line) for the non-radiative fronts with $\sigma_{\rm u}=3$, 10, and 20.
-Figure 5: Relation between $\kappa=\gamma_{\rm D}(1+\beta_{\rm D})$ and $C=\rho/\rho_{\rm u}$ (here measured at the peak inside the compression front) during relaxation toward the final steady state.
-Black curve shows the simulation result (the model with $\sigma_{\rm u}=20$), and red dashed curve shows the relation $C=(\kappa^{2}+1)/2$, equivalent to $\tilde{\rho}/\kappa=\rho_{\rm u}$ (a uniform mass flux in $\xi$, $F_{\rm m}=\rho(c-v)$).
-Fluid density in the front closely tracks the steady-state solution $\tilde{\rho}=\kappa\rho_{\rm u}$ defined for an instantaneous profile of $\kappa(\xi)$, which slowly grows during relaxation.
-The correlated growth of $\kappa$ and $C$ stops when they reach $\kappa_{\max}=5$ and $C_{\max}=13$.
-Then, the fluxes $F_{\rm m}$ and $F_{-}$ both become exactly uniform across the front, and a true steady state is achieved.
-We conclude that at any large but finite $\sigma_{\rm u}$, the wave packet approaches the steady state on a long timescale $t_{\star}\sim\sigma_{\rm u}t_{\rm cross}$ (assuming negligible damping of the radio wave during this time).
-The conclusion might seem counter-intuitive: one might expect that the front has no memory on timescales longer than $t_{\rm cross}$ since the plasma interacting with the wave packet is continually refreshed on the timescale $t_{\rm cross}$.
-Nevertheless, the conservation laws require the propagating wave to retain some memory of its previous interactions with the plasma.
-Given enough time, the packet gradually builds up the strong compression $\tilde{\rho}/\rho_{\rm u}=\kappa\approx w$ — the pattern of the MHD flow in the wave is “groomed” toward the steady state.
-Remarkably, this gradual evolution leads to a strong magnetic field in the packet, so that the plasma magnetization approaches $\sigma\approx w\sigma_{\rm u}$.
-Thus, tremendous magnetic pressure develops in the propagating front while it finds a self-consistent steady pattern.
-At the same time, the final pattern of fluid motion established in the wave packet is the same as it would be in the absence of magnetic pressure, as the background magnetic field drops out from the steady-state solution $\kappa_{\star}(\xi)$.
-6.3 Evolution of a radiative compression front
-In the regime of $q\gg 1$, the fluid element moving through the wave packet experiences strong radiative losses.
-The radio wave sustains the fluid enthalpy $w\approx(3/2)a$ by continually depositing energy to offset the losses (B22).
-Thus, the received power per particle (measured in the local fluid frame) equals the emitted power ˙$\dot{{\cal E}}_{e}$ (Lorentz invariant).
-The wave-fluid interaction is basically a scattering process with the cross section $\sigma_{\rm sc}=\dot{{\cal E}}_{e}/F$, where $F=cU$ is the energy flux of the radio wave.
-The scattering is accompanied by a radiative pressure force $\dot{\cal P}$, which is Lorentz-invariant and easiest to find in the local fluid frame where $\dot{\cal P}=\dot{{\cal E}}_{e}/c$, so
-$\dot{\cal P}=\frac{\dot{{\cal E}}_{e}}{c}=\sigma_{\rm T}\tilde{U}\,\tilde{\gamma}^{2}=\frac{\sigma_{\rm T}U}{\kappa^{2}}\,\tilde{\gamma}^{2}.$ (78)
+We conclude that at any large but finite $\sigma_{\rm u}$, the wave packet approaches the steady state on a long timescale $t_{\star}\sim\sigma_{\rm u}t_{\rm cross}^{\star}$ (assuming negligible damping of the radio wave during this time).
+This conclusion may seem counter-intuitive: one might expect that the front has no memory on timescales longer than $t_{\rm cross}$ since the plasma interacting with the wave packet is continually refreshed on the timescale $t_{\rm cross}$.
+Nevertheless, the conservation laws require the propagating front to retain some memory of its earlier structure.
+Given enough time, the front gradually builds up the strong compression factor $\tilde{\rho}/\rho_{\rm u}=\kappa\approx 1+w$ across the entire front — the pattern of the MHD flow in the wave is “groomed” toward the steady state.
+Remarkably, this gradual evolution leads to a strong magnetic field in the packet, as the plasma magnetization approaches $\sigma\approx(1+w)\sigma_{\rm u}$.
+Thus, huge magnetic pressure develops in the propagating front while it finds a self-consistent steady pattern.
+At the same time, the final pattern of fluid motion in the wave packet is the same as found for radio waves propagating in an unmagnetized plasma.
+The magnetic pressure drops out from the steady-state solution $\kappa=\kappa_{\star}(\xi)$ because $\sigma/\kappa$ is exactly uniform across the compression front (section V).
+The relaxation of the compression front can be viewed as the evolution of a driven MHD wave in the plasma.
+The FRB drives a smooth compression wave of width $cT$, and it takes some time $t\sim t_{\star}$ for the wave pattern to find its equilibrium shape that propagates together with its driver without further change.
+Recall that in MHD compression tends to be redistributed across the plasma with Lorentz factor $\tilde{\gamma}_{\rm s}\approx\sqrt{1+\sigma_{\rm eff}}$ (Appendix A).
+The compression pattern evolves due to the driving effect of the FRB and the communication of MHD information across the pattern, which occurs along characteristics $\cal C$$^{-}$ and $\cal C$$^{+}$.
+The $\cal C$$^{\pm}$ characteristics propagate in $z$ with speeds $\tilde{\beta}_{\pm}=\pm\tilde{\beta}_{\rm s}$ relative to the fluid, which itself moves with speed $\beta_{\rm D}$ relative to a fixed lab frame.
+The $\cal C$$^{\pm}$ speeds in the lab frame $\beta_{\pm}$ are determined by the relativistic velocity transformation of $\tilde{\beta}_{\pm}$ from the fluid frame $\tilde{\cal K}$ to the lab frame.
+Instead of speeds $\beta_{\pm}$, it is convenient to use $\kappa_{\pm}\equiv\gamma_{\pm}(1+\beta_{\pm})$, similar to $\kappa$ defined for the fluid speed $\beta_{\rm D}$ in Equation (10).
+Then, the transformation between the frames takes the form
+[…]
+Note that $\tilde{\kappa}_{+}\tilde{\kappa}_{-}=1$, since $\tilde{\beta}_{+}=\tilde{\beta}_{\rm s}=-\tilde{\beta}_{-}$.
+Magnetically dominated plasma with $\sigma_{\rm eff}\gg 1$ has $\tilde{\gamma}_{\rm s}\approx\sqrt{\sigma_{\rm eff}}$, $\tilde{\kappa}_{+}\approx 2\tilde{\gamma}_{\rm s}$ and $\tilde{\kappa}_{-}\approx(2\tilde{\gamma}_{\rm s})^{-1}$.
+The MHD characteristics $\cal C$$^{\pm}$ drift in $\xi$ with rate
+$\frac{d\xi_{\pm}}{dt}=1-\beta_{\pm}=\frac{2}{\kappa_{\pm}^{2}+1}.$ (71)
+Note that MHD information flows across the compression front in the direction of increasing $\xi$: $d\xi_{\pm}/dt>0$.
+As one can see in Figure 4, the front evolution is basically the gradual spreading of the relaxed zone $\kappa=\kappa_{\star}$ from the leading edge $\xi=0$ toward larger $\xi$.
+The steady state is first established near the leading edge of the front where the relaxation time $t_{\star}(\xi)$ is short because $t_{\rm cross}(\xi)$ is short for a small leading part $\xi\ll T$.
+The middle part of the front takes longer to relax; however, its difference from the steady-state solution, $\kappa\neq\kappa_{\star}$, cannot perturb the already relaxed leading part, since MHD information does not spread to smaller $\xi$.
+In a plasma with $\sigma_{\rm eff}\gg 1$, MHD information carried by $\cal C$$^{+}$ crosses the front much slower than $\cal C$$^{-}$:
+$\frac{d\xi_{+}}{dt}\approx\frac{2}{\kappa_{+}^{2}}\approx\frac{1}{2\sigma_{\rm eff}\kappa^{2}}\ll\frac{d\xi_{-}}{dt}.$ (72)
+Thus, the timescale for complete MHD relaxation is related to $\cal C$$^{+}$ propagation across the front rather than $\cal C$$^{-}$.
+The drift $d\xi_{+}/dt\approx 2/\kappa_{+}^{2}$ is slower than the plasma drift $d\xi/dt\approx 2/\kappa^{2}$ by the factor of $\tilde{\kappa}_{+}^{2}\approx 4\sigma_{\rm eff}$ (when using a lab frame where the plasma has $\kappa\gg 1$).
+Therefore, the timescale $t_{+}$ for MHD information to cross the compression front is $\approx 4\sigma_{\rm eff}$ times longer than the timescale for the plasma to cross the front:
+$t_{+}\approx 4\sigma_{\rm eff}t_{\rm cross}.$ (73)
+When the entire front approaches the steady state, $\kappa\approx\kappa_{\star}=\tilde{\gamma}\kappa_{\rm u}$ implies $\sigma_{\rm eff}\approx\sigma_{\rm u}$ (uniform across the front) and $t_{+}^{\star}\approx 4\sigma_{\rm u}t_{\rm cross}^{\star}$, where star indicates a steady state.
+This expression is similar to the heuristic estimate for relaxation time $t_{\star}$ given in Equation (66), within the numerical factor of $4$.
+VI.2 Evolution of a radiative compression front
+If $q\gg 1$, the fluid element moving through the wave packet experiences strong radiative losses.
+In this case, the radio wave continually deposits energy to offset the losses and keep the average $\tilde{\gamma}=\sqrt{1+a^{2}}$ (B22).
+The deposited power per particle in the fluid frame approximately equals the average emitted power $\delta\tilde{{\cal E}}_{e}/\delta\tilde{t}$ (given by Equation (D9) with $\beta_{\rm D}=0$), so the wave-fluid interaction at $q\gg 1$ is essentially a scattering process.
+The scattering exerts a force on the fluid.
+The wave passes momentum with rate $\delta\tilde{{\cal E}}_{e}/c\delta\tilde{t}$, and the fluid radiates momentum with rate $\delta\tilde{{\cal P}}_{e}^{z}/\delta\tilde{t}$, so the net rate of momentum deposition into the MHD fluid is
+$\frac{\delta\tilde{{\cal P}}}{\delta\tilde{t}}=\frac{\,\delta\tilde{{\cal E}}_{e}}{c\delta\tilde{t}}-\frac{\delta\tilde{{\cal P}}_{e}^{z}}{\delta\tilde{t}}$. (74)
+Transformation from the local fluid frame to a lab frame gives $\delta{\cal P}/\delta t=\delta\tilde{{\cal P}}/\delta\tilde{t}$ since $\delta t=\gamma\delta\tilde{t}$ and $\delta{\cal P}=\gamma\delta\tilde{\cal P}$ (from Lorentz transformation, using $\delta\tilde{\cal E}/\delta\tilde{\cal P}=\tilde{v}_{\rm D}=0$).
+Thus, Equation (74) gives the Lorentz-invariant force $\dot{\cal P}_{\rm sc}=\delta{\cal P}/\delta t$ applied to the MHD fluid by the scattering of the radio wave.
+Substitution of Equations (D9) and (D10) (with $\beta_{\rm D}=0$ in frame $\tilde{\cal K}$) yields
+$\dot{\cal P}_{\rm sc}=\frac{\sigma_{\rm T}U}{\kappa^{2}}(1+a^{2}).$ (75)
 A fluid element crossing the front gains from scattering the following net momentum per particle:
-[…]
+$\Delta{\cal P}_{\rm sc}$ $=\int_{0}^{t_{\rm cross}}\dot{\cal P}_{\rm sc}\,dt=\int_{0}^{T}\frac{\dot{\cal P}_{\rm sc}\,d\xi}{1-\beta_{\rm D}}$
 $\sim\left(1+\frac{1}{\kappa^{2}}\right)\sigma_{\rm T}Ua^{2}T,$ (76)
 with $a\sim a_{\max}$.
-This sets the budget for the fluid momentum exchange with the wave per crossing time $t_{\rm cross}$.
+This sets the budget for the fluid momentum gain per crossing time $t_{\rm cross}$.
 The gained momentum is shared with the magnetic field frozen in the fluid, increasing ${\cal P}_{\rm MHD}$.
@@
 $t_{\cal P}\sim\frac{{\cal P}_{\rm MHD}}{\Delta{\cal P}_{\rm sc}}\,t_{\rm cross}\sim\frac{h\kappa}{qw^{2}}\,t_{\rm cross},$ (77)
-where $q$ is evaluated using the entire column energy of the wave packet $\Sigma(T)$ (Equation 59).
-Equation (80) gives the timescale for approaching the steady state, $t_{\star}=t_{\cal P}(\kappa_{\star})$ if we substitute $\kappa\sim\kappa_{\star}\approx qw$ (Equation 58) and $h_{\star}=1+w+\kappa_{\star}\sigma_{\rm u}\approx\kappa_{\star}\sigma_{\rm u}$:
-$t_{\star}\sim q\,\sigma_{\rm u}t_{\rm cross}\qquad(q\gg 1,\;\,\sigma_{\rm u}\gg 1).$ (81)
+where $q$ may be evaluated using the column energy $\Sigma$ in the middle of the wave packet (Equation 57).
+Equation (77) gives the timescale for approaching the steady state, $t_{\star}=t_{\cal P}(\kappa_{\star})$ if we substitute $\kappa\sim\kappa_{\star}=(1+q)(1+w)$ (Equation 56) and $h_{\star}=1+w+\kappa_{\star}\sigma_{\rm u}\approx\kappa_{\star}\sigma_{\rm u}$.
+Interpolating between the limits of $q\ll 1$ and $q\gg 1$, we find
+$t_{\star}\sim(1+q)\,\sigma_{\rm u}t_{\rm cross}^{\star}\qquad(\sigma_{\rm u}\gg 1).$ (78)
 This estimate may be compared with the results of numerical simulations presented below.
-Front relaxation to the steady state in the radiative regime is found by solving Equations (50) and (51) similarly to section 6.2, but now keeping the term $\dot{\gamma}_{e}\neq 0$.
-This term may be expressed as follows
-$\dot{\gamma}_{e}={\cal R}\frac{w^{4}}{\kappa^{2}}\quad(w\gg 1),\qquad{\cal R}\equiv\left(\frac{2}{3}\right)^{5}\frac{r_{e}}{c}\,\omega^{2}.$ (82)
-The value of ${\cal R}$ depends on the choice of a reference frame, which can be parameterized by the upstream motion $\beta_{\rm u}$ or $\kappa_{\rm u}$.
-Note that $\kappa_{\rm u}<1$ ($\beta_{\rm u}<0$) in frames moving in the $+z$ direction relative to the upstream plasma.
-The wave frequency $\omega$ in a chosen frame scales as $\omega\propto\kappa_{\rm u}$, so ${\cal R}\propto\kappa_{\rm u}^{2}$.
-Note also that the packet width $T$ depends on the choice of a frame as $\kappa_{\rm u}^{-1}$ (since $\omega T=inv$), and the radiative parameter $q\propto T{\cal R}/\kappa_{\rm u}$ is invariant.
+Front relaxation to the steady state in the radiative regime is found by solving Equations (46) and (47) similarly to section VI.1, but now keeping the term $n\dot{\cal Q}^{-}_{e}/c^{2}$.
+Using Equation (50) and $U=m\,\omega^{2}a^{2}/4\pi r_{e}$ (Equation 28), we rewrite $\dot{\cal Q}^{-}_{e}$ as follows
+$\frac{\dot{\cal Q}^{-}_{e}}{mc^{2}}={\cal R}\,\frac{a^{2}(1+a^{2})}{\gamma_{\rm D}\kappa^{3}},\qquad{\cal R}\equiv\frac{2r_{e}}{3c}\,\omega^{2}.$ (79)
+[…]
+$q(T)=\frac{{\cal R}}{\kappa_{\rm u}}\int_{0}^{T}a^{2}d\xi=\frac{3}{8}a_{\max}^{2}\frac{T{\cal R}}{\kappa_{\rm u}}.$ (80)
+A characteristic value of $q$ in the middle of the front is $q(T/2)=q(T)/2$.
+[…]
 We use the following technical tricks to avoid two complications:
-(1) Radiative losses $\dot{\gamma}_{e}$ may be important only when $w\gg 1$ because of the small value of ${\cal R}$ for real FRBs.
-We avoid simulating huge $w$ and $\kappa$ by rescaling ${\cal R}$ to larger values, so that we can experiment with the radiative regime $q\gg 1$ at moderate $w$ and $\kappa$ accessible to simulations.
-In sample models, we use the wave packet described by Equation (5) with $a_{\max}=(4/3)\sqrt{7}$ (which gives $w_{\max}=4$) and set the invariant radiative parameter $T{\cal R}/\kappa_{\rm u}=1$.
-This choice gives $q\sim 3$.
+[…]
 (2) Both $\kappa$ and $\rho$ grow by a large factor during relaxation toward the steady state.
-The simulation is more efficient if performed in a boosted frame where $\kappa_{\rm u}\ll 1$.
+The simulation more efficiently performed in a boosted frame where $\kappa_{\rm u}\ll 1$.
 This helps avoid ultra-relativistic motions of the accelerated fluid along the wave propagation direction (large $\kappa\gg 1$) that would slow down the simulation.
@@
 When presenting the results, we transform the fluid parameters to the upstream rest frame (where $\kappa_{\rm u}=1$), so figures will show $\kappa$ and $C=\rho/\rho_{\rm u}$ measured in the upstream frame.
-Figure 6: Evolution of the compression peak $C_{\max}$ from the non-radiative steady state ($C_{\max}=C_{\max}^{\star}=13$ at $t=0$) after switching on radiative losses with $T{\cal R}/\kappa_{\rm u}=1$ (the radiative parameter ${\cal R}$ is defined in Equation (82)).
-The front evolves toward the new steady state with $C_{\max}^{\star}\approx 147$.
-The evolution is shown for three simulations with upstream magnetizations $\sigma_{\rm u}=3$, 10, 20.
-The initial and final states are independent of $\sigma_{\rm u}$ and shown in Figure 7.
-Figure 7: The initial and final profiles of the compression front in the three simulations shown in Figure 6.
-The initial state is the steady non-radiative front (Figure 3, red curve); here it is shown by the black dashed curve.
-Radiative losses make the front evolve until it reaches a new steady state (solid curve), which reproduces the solution in Equation (58).
-As an initial condition at $t=0$, we take the MHD fluid in the non-radiative steady front.
-In the rest frame of the upstream plasma, it is given by $\kappa_{0}(\xi)=1+w(\xi)$ and $C_{0}(\xi)=(\kappa_{0}^{2}+1)/2$.
-In particular, the initial compression peak is $C_{\max}(0)=13$.
-In the presence of radiative losses, this flow pattern is no longer a steady state, so it immediately begins to evolve at $t>0$.
+[…]
+Figure 8: Initial (dashed) and final (solid) profiles of the compression front in the three simulations shown in Figure 7.
+The initial state is the steady non-radiative front (Figure 4, red curve).
+Radiative losses make the front evolve until it reaches the new steady state, which reproduces the analytical solution (Equation 56).
+As an initial condition at $t=0$, we take the steady flow pattern in the non-radiative front.
+In the rest frame of the upstream plasma, it is given by $\kappa_{0}(\xi)=\sqrt{1+a^{2}(\xi)}$ and $C_{0}(\xi)=1+a^{2}(\xi)/2$, so the initial compression peak is $C_{\max}(0)=13$.
+In the presence of radiative losses, this flow pattern is no longer a steady state, and it immediately begins to evolve at $t>0$.
 Both $\kappa$ and $C$ grow and eventually relax to a new steady state.
-Figure 6 shows the evolution of $C_{\max}$, and Figure 7 shows the final profiles of $\kappa(\xi)$ and $C(\xi)$.
-The relaxation is determined by the chosen radiative coefficient ${\cal R}$ and the corresponding $q\sim 3$.
-In particular, $\kappa_{\max}$ grows from its initial value $51+w_{\max}=$ 5 to the final value $\kappa_{\max}=17$, consistent with the factor of $q\sim 3$, and $C_{\max}$ grows from 13 to $147$, confirming the expected increase $\sim q^{2}$.
-The simulations also confirm the expected scaling of the relaxation timescale $t_{\star}\propto q\sigma_{\rm u}$ at $q\gg 1$ and $\sigma_{\rm u}\gg 1$ (Equation 81).
-This scaling can be seen in Figure 6, which shows three models with $\sigma_{\rm u}=3$, 10, and 20.
-Note that $t_{\rm cross}\propto\kappa^{2}\propto q^{2}w^{2}$, and hence $t_{\star}$ should scale as $q^{3}$ at $q\gg 1$.
-One can see that the relaxation timescale is indeed increased by a large factor $\sim q^{3}\sim 30$ compared with the non-radiative front evolution shown in Figure 4.
-As expected, the final steady state is independent of $\sigma_{\rm u}$: it is the same in all three simulations with $\sigma_{\rm u}=3,10,20$.
-The steady-state profiles of $\kappa(\xi)$ and $C(\xi)$ are shown in Figure 7 and agree with Equation (58).
-Note that in the non-radiative regime ($q\ll 1$) the fluid comes back to rest as it exits behind the wave packet.
-In contrast, the fluid behind the radiative front has a large speed $\beta_{\rm D}>0$ ($\kappa\gg 1$).
-This is a consequence of momentum deposition by scattering during the fluid motion through the wave packet.
-7 Parameters of the wind
-The obtained picture of a radio wave interaction with MHD fluid can now be applied to FRBs propagating through a radial wind.
-The wind serves as the upstream medium for the wave packet, and we begin with a brief summary of its parameters.
+[…]
+The simulations also confirm the expected scaling of the relaxation timescale $t_{\star}\sim(1+q)\sigma_{\rm u}t_{\rm cross}^{\star}$ at $\sigma_{\rm u}\gg 1$ (Equation 78).
+The scaling $t_{\star}\propto\sigma_{\rm u}$ is seen in Figure 7, which shows three models with $\sigma_{\rm u}=3$, 10, and 20.
+Note also that $t_{\rm cross}^{\star}\propto\kappa_{\star}^{2}\propto(1+q)^{2}$ and hence $t_{\star}$ is expected to scale as $(1+q)^{3}$.
+We observe that the relaxation timescale is indeed increased by a large factor of $\approx(1+q)^{3}/2$ compared to the non-radiative front relaxation shown in Figure 6.
+As expected, the final steady state (Figure 8) is independent of $\sigma_{\rm u}$: it is the same in all three simulations with $\sigma_{\rm u}=3,10,20$.
+The observed final state satisfies Equation (56) with $\kappa_{\rm u}=1$ (since we view the results in the upstream rest frame).
+Note that in the non-radiative regime ($q\ll 1$) fluid comes back to rest as it exits behind the wave packet.
+By contrast, fluid behind the radiative front ($q\gg 1$) retains a large speed $\beta_{\rm D}\sim 1$ ($\kappa\gg 1$).
+This is the result of irreversible momentum deposition by the scattering of the radio wave.
+VII Parameters of the wind
+The obtained picture of the radio wave interaction with MHD fluid can now be applied to FRBs propagating through a radial wind carrying a transverse magnetic field, with some magnetization parameter $\sigma_{\rm u}$.
+The wind serves as the upstream medium for the radio wave packet, and we begin with a brief summary of its parameters.
 They are denoted with subscript “u” for consistency of notation in this paper.
@@
 The spindown power gives a lower bound on $L_{\rm w}$, which can be enhanced during magnetar activity.
-The wind forms at the light cylinder $R_{\rm LC}=c/\Omega$ and its Lorentz factor grows as $\gamma_{\rm u}(r)\approx r/R_{\rm LC}$ until it reaches $\eta^{1/3}$; at larger radii, $\gamma_{\rm u}$ grows logarithmically (Lyubarsky & Kirk, 2001).
+The wind forms at the light cylinder $R_{\rm LC}=c/\Omega$ and its Lorentz factor grows as $\gamma_{\rm u}(r)\approx r/R_{\rm LC}$ until it reaches $\eta^{1/3}$; at larger radii, $\gamma_{\rm u}$ grows logarithmically (Beskin et al., 1998; Lyubarsky & Kirk, 2001).
 We will use a simple approximation:
-$\gamma_{\rm u}\approx\left\{\begin{array}[]{cc}\vspace*{2mm}\displaystyle{\frac{r}{R_{\rm LC}}}&\quad r<\eta^{1/3}R_{\rm LC}\\ \displaystyle{\left[\frac{\eta}{\ln 2}\ln\left(1+\frac{r}{\eta^{1/3}R_{\rm LC}}\right)\right]^{1/3}}&\quad r>\eta^{1/3}R_{\rm LC}\\ \end{array}\right.$ (85)
+[…]
 Our description of the wind is imprecise at small radii $r\sim R_{\rm LC}$ for two reasons.
@@
 Note that $L_{\rm w}$ determines the wind magnetic field in the static lab frame $B_{\rm u}(r)$ according to Equation (81).
-The corresponding magnetic field in the rest frame of the wind (ahead of the FRB wave packet) is $\tilde{B}=B_{\rm u}/\gamma_{\rm u}$.
+The corresponding magnetic field in the rest frame of the wind (ahead of the FRB wave packet) is $\tilde{B}_{\rm u}=B_{\rm u}/\gamma_{\rm u}$.
 It sets another important parameter — the gyro-frequency of the wind plasma,
@@
 Then, we find
-$\alpha\equiv\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}\approx\frac{2e}{mcr\omega}\left(\frac{L_{\rm w}}{c}\right)^{1/2}\approx 0.1\,\frac{L_{\rm w,37}^{1/2}}{\nu_{9}r_{12}}.$ (88)
+$b_{\rm u}\equiv\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}=\frac{r_{b}}{r},\quad\;\,r_{b}=\frac{2eL_{\rm w}^{1/2}}{mc^{3/2}\omega}\approx 10^{11}\frac{L_{\rm w,37}^{1/2}}{\nu_{9}}\,{\rm cm}.$ (86)
 Below we examine the compression front driven by an FRB emitted at some radius $r$ and propagating outward through the wind.
-We first describe the front at large $r$, where FRBs induce heating in the regular regime ($\tilde{\omega}_{B}<\tilde{\omega}$), and then consider FRBs emitted at smaller $r$, where heating is stochastic and more dramatic effects occur, including damping of the FRB.
-8 Regular heating zone
-Strong compression is expected at radii where the FRB has $a>1$. The wave strength parameter $a$ (Equation 1) may be written as $a=r_{1}/r$, where
-$r_{1}=\left(\frac{r_{e}L}{mc\,\omega^{2}}\right)^{1/2}\approx 1.6\times 10^{13}\,L_{42}^{1/2}\nu_{9}^{-1}\,{\rm cm}.$ (89)
-8.1 Saturated compression ($r>r_{\star}$)
-One can quickly formulate a compression front model using the steady-state solution described by Equations (52), (57), and (75), assuming the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}$) and neglecting radiative losses.
-Using the definition of the proper compression factor $\tilde{C}\equiv\tilde{\rho}/\tilde{\rho}_{\rm u}$, the steady front solution can be stated as
-$\tilde{C}_{\star}=\frac{\kappa_{\star}}{\kappa_{\rm u}}=1+\frac{3}{2}\left(\sqrt{a^{2}+1}-1\right).$ (90)
-The corresponding compression factor for the wind density measured in the lab frame, $C\equiv\rho/\rho_{\rm u}$, is
-$C_{\star}=\frac{\kappa_{\star}^{2}+1}{\kappa_{\rm u}^{2}+1}\approx\tilde{C}_{\star}^{2},$ (91)
+We first describe the front at large $r$, where plasma responds to the radio wave in the regular oscillation regime.
+Then, we consider FRBs emitted at smaller $r$, where stochastic heating occurs and leads to more dramatic effects, including damping of the FRB.
+VIII Regular oscillation zone
+Strong compression is expected at radii where the FRB wave packet has $a_{\max}>1$. The strength parameter $a_{\max}$ (Equation 1) may be written as
+$a_{\max}=\frac{r_{1}}{r},$ (87)
+where
+$r_{1}=\left(\frac{r_{e}L}{mc\,\omega^{2}}\right)^{1/2}\approx 1.6\times 10^{13}\,L_{42}^{1/2}\nu_{9}^{-1}\,{\rm cm},$ (88)
+and $L$ is the peak luminosity of the FRB.
+VIII.1 Saturated compression ($r>r_{\star}$)
+One can quickly formulate a compression front model using the steady-state solution described by Equations (51) and (58), which assumes regular oscillations of the plasma in the radio wave.
+The shape of the wave packet $a(\xi)$ then determines the profile of the proper compression factor in the front, $\tilde{C}\equiv\tilde{\rho}/\tilde{\rho}_{\rm u}$,
+$\tilde{C}_{\star}=\frac{\kappa_{\star}}{\kappa_{\rm u}}=\sqrt{1+a^{2}},$ (89)
+where star indicates that the solution is stated for the steady front. The corresponding compression of the wind density measured in the lab frame, $C\equiv\rho/\rho_{\rm u}$, is
+$C_{\star}(\xi)=\frac{\kappa_{\star}^{2}+1}{\kappa_{\rm u}^{2}+1}\approx 1+a^{2},$ (90)
 where $\kappa\equiv\gamma_{\rm D}(1+\beta_{\rm D})$ describes the fluid drift speed $\beta_{\rm D}$, and we used $\kappa_{\rm u}\gg 1$ since the upstream medium is a relativistic wind.
-Outside the radius $r_{1}$ (where $a\ll 1$), one finds weak compression $C_{\star}\approx 1$ while inside $r_{1}$ (where $a\gg 1$) one finds $C_{\star}\approx(9/4)a^{2}=(9/8)a_{0}^{2}\approx a_{0}^{2}$.
-A simple approximate expression matching the two cases is
-$C_{\star}\approx 1+a_{0}^{2}.$ (92)
-Remarkably, this compression is consistent with the results of Sobacchi et al. (2024), who investigated a steady plasma flow in a (nearly monochromatic) wave packet using a different technique.
-However, the simple solution stated in Equation (92) does not correctly describe compression fronts in FRBs at small radii $r\ll r_{1}$, for two reasons:
-(1) At $r<r_{\star}$ defined below, the compression front does not reach the steady state, so the actual compression factor is well below its saturated value, $C\ll C_{\star}$.
-(2) Compression promotes the transition to stochastic acceleration of plasma particles in the FRB wave packet.
-The plasma then experiences much stronger heating, and the momentum exchange between the FRB wave packet and the plasma is enhanced, changing $\kappa$ and $C$.
-These two features qualitatively change the picture of the FRB compression front, as described below.
-8.2 Unsaturated compression $(r<r_{\star})$
-The saturated (steady-state) compression factor $C_{\star}$ may be reached only when the wave packet has propagated for a time $t>t_{\star}$.
-In the regular heating regime ($\tilde{\omega}_{B}<\tilde{\omega}_{B}$), the timescale $t_{\star}$ is given by Equation (74) in the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$).
-The corresponding timescale measured in any other frame (any $\kappa_{\rm u}$) is $t_{\star}\sim\sigma_{\rm u}a^{2}\gamma_{\rm u}^{2}T$ (assuming $a^{2}>1$).
-So, the relaxation timescale in our static lab frame is
-$t_{\star}\sim\eta\gamma_{\rm u}a^{2}T.$ (93)
+The maximum compression in the steady-state front is at the peak of the wave packet: $C_{\star}^{\max}\approx 1+a_{\max}^{2}$.
+The simple solution stated in Equation (90) is valid only at sufficiently large radii $r>r_{\star}$ evaluated below. It does not describe compression fronts for FRBs at smaller radii for two reasons:
+(1) At $r<r_{\star}$, the compression front does not reach the steady state (section VIII.2). Then, the actual compression factor is well below its saturated value, $C\ll C_{\star}$.
+(2) Compression promotes the transition to stochastic heating of plasma particles by the radio wave (section IX).
+The plasma then receives much more energy (and hence much more momentum), which changes $\kappa$ and $C$.
+VIII.2 Unsaturated compression $(r<r_{\star})$
+The saturated (steady-state) compression factor $C_{\max}^{\star}$ may be reached only when the wave packet has propagated for a time $t>t_{\star}\approx(1+\sigma_{\rm u})t_{\rm cross}^{\star}$ (section VI.1).
+The magnetar wind has $\sigma_{\rm u}\gg 1$ and moves with Lorentz factor $\gamma_{\rm u}\approx\kappa_{\rm u}/2\gg 1$.
+Then, the relaxation timescale $t_{\star}$ may be expressed as
+$t_{\star}\approx\frac{3}{4}\eta\gamma_{\rm u}a_{\max}^{2}T,$ (91)
+where we used Equation (61) for $t_{\rm cross}^{\star}$ and $\sigma_{\rm u}\gamma_{\rm u}=\eta$.
 […]
 $\frac{t}{t_{\star}}=\frac{r^{3}}{r_{\star}^{3}},$ (92)
-$r_{\star}\approx\left(\frac{r_{e}{\cal E}\eta\gamma_{\rm u}}{m\omega^{2}}\right)^{1/3}\approx 10^{13}\,\frac{{\cal E}_{39}^{1/3}\eta_{4}^{4/9}}{\nu_{9}^{1/3}}\,{\rm cm},$ (95)
-where ${\cal E}\sim LT$ is the FRB energy. This expression for $r_{\star}$ assumes $a^{2}(r_{\star})>1$, i.e. it holds as long as $r_{\star}<r_{1}$. At the transition radius $r=r_{\star}$ we find
-$C(r_{\star})\approx C_{\star}(r_{\star})\approx\left[1+\frac{3}{2}\left(\sqrt{\frac{r_{1}^{2}}{r_{\star}^{2}}+1}-1\right)\right]^{2}.$ (96)
-When $t_{\star}>t=r/c$, the propagating wave packet develops compression factors $C<C_{\star}$ (Figure 4). At times $\gamma_{\rm u}^{2}T\ll t\ll t_{\star}$, $C(t)$ may be approximated by
-$C-1\approx(C_{\star}-1)\left(\frac{t}{t_{\star}}\right)^{\zeta}\approx\frac{9}{4}a^{2}\left(\frac{r}{\eta\gamma_{\rm u}a^{2}cT}\right)^{\zeta}.$ (97)
-The numerical results in Figure 4 suggest $\zeta\approx 1/2$.
-9 Stochastic heating zone
-The zone of stochastic heating ($\tilde{\omega}_{B}>\tilde{\omega}$) is larger than one might expect from the ratio $\tilde{\omega}_{B}^{\rm u}/\tilde{\omega}_{\rm u}$ (Equation 88).
-Even when $\tilde{\omega}_{B}<\tilde{\omega}$ in the upstream plasma, the condition $\tilde{\omega}_{B}>\tilde{\omega}$ can be met inside the wave packet as a result of the fluid acceleration and compression: acceleration reduces $\tilde{\omega}$ and compression increases $\tilde{\omega}_{B}$.
+$r_{\star}\approx\left(\frac{2r_{e}{\cal E}\eta\gamma_{\rm u}}{m\omega^{2}}\right)^{1/3}\approx 1.5\times 10^{13}\,\frac{{\cal E}_{39}^{1/3}\eta_{4}^{4/9}}{\nu_{9}^{2/3}}\,{\rm cm}.$ (93)
+[…]
+${\cal E}=\frac{L}{a_{\max}^{2}}\int_{0}^{T}a^{2}d\xi=\frac{3}{8}LT,$ (94)
+where the numerical factor $3/8$ is found for the concrete shape of the radio wave packet given by Equation (7).
+The expression for $r_{\star}$ assumes $a_{\max}^{2}(r_{\star})>1$, i.e.
+it holds as long as $r_{\star}<r_{1}$.
+At the transition radius $r_{\star}$, we find
+$C_{\max}^{\star}(r_{\star})\approx 1+\frac{r_{1}^{2}}{r_{\star}^{2}}.$ (95)
+At radii $r<r_{\star}$, the propagation time $t<t_{\star}$ implies that the wave packet develops compression factors $C_{\max}(t)<C_{\max}^{\star}$ (Figure 6).
+In this regime, $C_{\max}$ may be approximated by
+$C_{\max}-1\approx(C_{\max}^{\star}-1)\left(\frac{t}{t_{\star}}\right)^{\alpha}=a_{\max}^{2}\left(\frac{r}{ct_{\star}}\right)^{\alpha}.$ (96)
+Numerical results in Figure 6 suggest $\alpha\approx 1/2$. The value of $\alpha$ depends on the assumed shape of the wave packet.
+IX Stochastic heating zone
+IX.1 Boundary of the stochastic heating zone
+The transition from regular oscillations to stochastic heating is well defined on the $a$-$b$ plane (Figure 3).
+The transition curve $b=b_{\rm s}(a)$ is approximately described by Equation (26).
+Even when $b_{\rm u}\ll 1$ in the upstream plasma ahead of the FRB, the condition $b>b_{\rm s}$ can be met inside the wave packet where the fluid is accelerated and compressed.
+Compression increases $\tilde{\omega}_{B}$ and acceleration reduces $\tilde{\omega}$, so their ratio $b=\tilde{\omega}_{B}/\tilde{\omega}$ is increased.
 Fluid acceleration ($\kappa/\kappa_{\rm u}$) and compression ($\tilde{C}$ or $C$) are related by
 $\tilde{C}\equiv\frac{\tilde{\rho}}{\tilde{\rho}_{\rm u}}\approx\frac{\kappa}{\kappa_{\rm u}},\qquad C\equiv\frac{\rho}{\rho_{\rm u}}\approx\frac{\kappa^{2}+1}{\kappa_{\rm u}^{2}+1}.$ (97)
-These relations are exact for a steady compression front (with $C=C_{\star}$ and $\kappa=\kappa_{\star}$) and also approximately hold for unsaturated compression ($C\ll C_{\star}$ and $\kappa\ll\kappa_{\star}$), as demonstrated in Figure 5 using the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$).
-Equation (98) states the $\kappa$-$C$ relation viewed in any reference frame, where $\kappa_{\rm u}$ can have any value.
+These relations are exact for a steady compression front (with $C=C_{\star}$ and $\kappa=\kappa_{\star}$) and also approximately hold for unsaturated compression, as demonstrated in Figure 5 using the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$).
+Equation (97) states the $\kappa$-$C$ relation viewed in other frames, where $\kappa_{\rm u}$ can have any value.
 The magnetic field is frozen in the fluid and compressed by the same factor as density:
-$\frac{\tilde{B}}{\tilde{B}_{\rm u}}=\tilde{C},\qquad\frac{B}{B_{\rm u}}=C,$ (99)
-where $\tilde{B}$ is measured in the local fluid rest frame, and subscript “u” indicates the upstream region ahead of the FRB. Equation (99) implies
-$\tilde{\omega}_{B}\approx\frac{\kappa}{\kappa_{\rm u}}\,\tilde{\omega}_{B}^{\rm u}.$ (100)
-The FRB frequency $\omega$ transforms to the local fluid frame as $\tilde{\omega}=\omega/\kappa$. Hence,
-$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}\approx\frac{\kappa^{2}}{\kappa_{\rm u}^{2}}\,\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}.$ (101)
-In the lab frame, the upstream medium (the wind) moves with $\kappa_{\rm u}\approx 2\gamma_{\rm u}\gg 1$, and Equation (101) may be re-written as
-$\frac{\tilde{\omega}_{B}}{\tilde{\omega}}\approx C\,\frac{\tilde{\omega}_{B}^{\rm u}}{\tilde{\omega}_{\rm u}}=C\alpha.$ (102)
-Now using Equation (88) one can see that $\tilde{\omega}_{B}>\tilde{\omega}$ when
-$C>C_{\rm stoch}\approx\frac{\tilde{\omega}_{\rm u}}{\tilde{\omega}_{B}^{\rm u}}\approx 10\,\frac{r_{12}\nu_{9}}{L_{\rm w,37}}.$ (103)
-The stochastic regime $\tilde{\omega}_{B}>\tilde{\omega}$ appears at $r<r_{\rm stoch}$, with the transition radius $r_{\rm stoch}$ typically inside $r_{\star}$.
-Its value is found by equating $C$ given by Equation (97) to $C_{\rm stoch}$.
-In particular, for $\zeta\approx 1/2$ we find
-$r_{\rm stoch}\approx\left(\frac{4r_{e}^{2}L_{\rm w}L}{m^{2}c^{3}\omega^{4}\eta\gamma_{\rm u}T}\right)^{1/3}.$ (104)
-This gives $r_{\rm stoch}\sim 10^{12}$ cm for typical parameters.
-When $r_{\rm stoch}>\eta^{1/3}R_{\rm LC}$, one can substitute $\gamma_{\rm u}\approx\eta^{1/3}$ and then Equation (104) implies $r_{\rm stoch}\propto L^{1/3}$.
-If one keeps $\zeta$ as a parameter that may differ from 1/2, one still finds $r_{\rm stoch}\propto L^{1/3}$:
-$r_{\rm stoch}\approx 7\times 10^{11}\,H\,L_{42}^{1/3}\,{\rm cm},$ (105)
-$H\approx\left(\frac{L_{\rm w,37}^{1/2}}{T_{\rm ms}^{\zeta}\eta_{4}^{4\zeta/3}\nu_{9}^{3-2\zeta}}\right)^{1/(3-3\zeta)}.$ (106)
-The expression for $H$ omits a $\zeta$-dependent prefactor, which deviates from unity when $\zeta$ deviates from 1/2.
-9.1 Heating and radiative cooling in the wave packet
-At radii $r<r_{\rm stoch}$, the FRB stochastically heats plasma particles to $\tilde{\gamma}\gg a$, far exceeding regular heating.
+$\frac{\tilde{B}}{\tilde{B}_{\rm u}}=\tilde{C},\qquad\frac{B}{B_{\rm u}}=C.$ (98)
+This implies $\tilde{\omega}_{B}\approx(\kappa/\kappa_{\rm u})\,\tilde{\omega}_{B}^{\rm u}$ while the FRB frequency $\omega$ transforms to the local fluid frame as $\tilde{\omega}=\omega/\kappa$, so one finds $b\approx(\kappa/\kappa_{\rm u})^{2}b_{\rm u}$.
+In the lab frame, the upstream medium (the wind) moves with $\kappa_{\rm u}\approx 2\gamma_{\rm u}\gg 1$.
+Then, $(\kappa/\kappa_{\rm u})^{2}\approx C$ and
+$b\approx Cb_{\rm u}.$ (99)
+Large compression $C$ inside the radio wave packet can lead to $b>b_{\rm s}$, enabling stochastic heating even when $b_{\rm u}\ll 1$.
+Therefore, the zone of stochastic heating $r<r_{\rm stoch}$ extends far outside $r_{b}$ defined in Equation (86).
+The boundary $r_{\rm stoch}$ is inside $r_{\star}$, and plasma compression by FRB at $r<r_{\star}$ occurs in the unsaturated regime: the peak $C_{\max}$ of $C(\xi)$ is reached at $\xi=\xi_{\star}<T/2$ (section III.2).
+The transition from regular oscillations to stochastic heating can only occur at $\xi<\xi_{\star}$ where $C(\xi)$ is rising.
+In this leading part of the wave packet, $C(\xi)$ tracks the steady-state $C_{\star}(\xi)$ (Figure 4), so
+$C\approx 1+a^{2},\qquad b\approx(1+a^{2})b_{\rm u}.$ (100)
+Stochastic heating is triggered where $b$ reaches $b_{\rm s}(a)$:
+$(1+a^{2})b_{\rm u}\approx b_{\rm s}(a).$ (101)
+This equation defines $a_{\rm stoch}$ and $C_{\rm stoch}\approx 1+a_{\rm stoch}^{2}$ at the transition from regular oscillations to stochastic heating.
+Figure 9 shows how fluid exposed to an FRB moves in the $a$-$b$ plane.
+At radii $r<r_{\rm stoch}$, the trajectory $b(a)$ reaches $b_{\rm s}(a)$, bringing the fluid to stochastic heating.
+Figure 9: Blue curves with arrows show the evolution of parameter $b\equiv\tilde{\omega}_{B}/\tilde{\omega}$ in a fluid element exposed to an FRB.
+Fluid is initially in the upstream region where $a=0$ and $b=b_{\rm u}=r_{b}/r$ (Equation 86).
+Parameter $b$ increases inside the radio wave packet in response to the rising $a$ (Equation 100) and can eventually reach $b_{\rm s}$, triggering stochastic heating.
+The four blue curves correspond to radii $r/r_{b}=3$, 5, 10, 20.
+Reaching the stochastic heating regime at large $r$ requires high FRB luminosities $L$ that give $r_{\rm stoch}>r$ (Equation 105).
+The transition point $b=b_{\rm s}$ is reached if the wave packet has $C_{\max}\geq C_{\rm stoch}$.
+Using $C_{\max}(r)$ given in Equation (96), one can state the condition $C_{\max}\geq C_{\rm stoch}$ as
+$a_{\max}^{2}\left(\frac{4r}{3cT\eta\gamma_{\rm u}a_{\max}^{2}}\right)^{\alpha}\geq a_{\rm stoch}^{2},$ (102)
+with $\alpha\approx 1/2$ suggested by our numerical models in section VI.1.
+Equation (102) defines the zone $r\leq r_{\rm stoch}$ where stochastic heating is triggered.
+Its boundary $r_{\rm stoch}$ is found using the known $b_{\rm s}(a)$ (Figure 3), the wind parameters $\eta$, $\gamma_{\rm u}(r)$, and $b_{\rm u}(r)$, and the known $a_{\max}(r)$ for a given FRB.
+An approximate expression for $r_{\rm stoch}$ may be obtained using Equation (26) for $b_{\rm s}(a)$. Equation (101) then gives
+$a_{\rm stoch}^{2}\approx\frac{1}{9b_{\rm u}^{2}}-1=\left(\frac{r}{3r_{b}}\right)^{2}-1.$ (103)
+The boundary $r=r_{\rm stoch}$ of the stochastic heating zone (Equation 102) satisfies
+$\frac{r_{1}}{r}\left(\frac{4r}{3cT\eta^{4/3}}\right)^{1/2}\approx\frac{r^{2}}{9r_{b}^{2}}-1,$ (104)
+where we substituted $a_{\max}=r_{1}/r$ and $\gamma_{\rm u}\approx\eta^{1/3}$. For bright FRBs, $r_{\rm stoch}\gg 3r_{b}$ and Equation (104) gives
+$r_{\rm stoch}\approx 10^{12}\frac{L_{\rm w,37}^{2/5}L_{42}^{1/5}}{\eta_{4}^{4/15}\nu_{9}^{6/5}T_{\rm ms}^{1/5}}\,{\rm cm}.$ (105)
+This expression approximately holds for FRBs with $L>10^{40}\,L_{\rm w,37}^{-1/2}\eta_{4}^{4/3}\nu_{9}T_{\rm ms}$ erg s-1.
+IX.2 Heating and radiative cooling in the wave packet
+At radii $r<r_{\rm stoch}$, the FRB stochastically heats plasma particles to $\tilde{\gamma}\gg a$, far exceeding the regular oscillation energy.
 In this regime, $\tilde{\gamma}$ monotonically grows in the fluid element as it crosses the wave packet (Figure 2).
-The growth is well described by Equation (18) with the numerical coefficient $\chi\approx 2/3$.
-It was derived in the frame where the upstream plasma is at rest ($\kappa_{\rm u}=1$), and a similar derivation can be repeated in any frame with $\kappa_{\rm u}\neq 1$ (the only change occurs in the expression for the gyro-frequency $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$ (Equation 100)).
-Thus, we find
-$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}\approx\frac{16\pi}{3}\frac{r_{e}U\kappa^{2/3}}{m\omega^{4/3}}\left(\frac{\tilde{\omega}_{B}^{\rm u}}{\kappa_{\rm u}}\right)^{1/3},$ (107)
+The growth is well described by Equation (25) with the numerical coefficient $\chi\approx 0.8$:
+$\tilde{\gamma}^{4/3}\,\frac{d\tilde{\gamma}}{d\xi}\approx 10\,\frac{r_{e}U\kappa^{2/3}}{m\omega^{4/3}}\left(\frac{\tilde{\omega}_{B}^{\rm u}}{\kappa_{\rm u}}\right)^{1/3},$ (106)
 where we substituted $a^{2}=4\pi r_{e}U/m\omega^{2}$. Integration for $\tilde{\gamma}(\xi)$ then gives
-$\tilde{\gamma}_{\rm stoch}\approx\left[\frac{3r_{e}{\cal E}}{mc\omega r^{2}}\right]^{3/7}\!\!\left(C\alpha\right)^{1/7}\approx 10^{5}\frac{{\cal E}_{39}^{3/7}C^{1/7}L_{\rm w,37}^{1/14}}{\nu_{9}^{4/7}r_{11}},$ (108)
-[…]
-Equation (108) holds as long as radiative losses of the plasma are negligible.
-The losses $\dot{\gamma}_{e}$ are given by Equation (36) (with the coefficient 2/3 replaced by 4/3 for stochastically heated particles, see B22).
-Losses $\dot{\gamma}_{e}$ offset the heating rate (Equation 10) when $\tilde{\gamma}$ approaches $\tilde{\gamma}_{\rm rad}=(c/r_{e})^{3/10}\tilde{\omega}_{B}^{1/10}\tilde{\omega}^{-2/5}$.
-Using $\tilde{\omega}=\omega/\kappa=\tilde{\omega}_{\rm u}\kappa_{\rm u}/\kappa$ and $\tilde{\omega}_{B}=\tilde{\omega}_{B}^{\rm u}\kappa/\kappa_{\rm u}$, we find
-$\tilde{\gamma}_{\rm rad}\approx\left(\frac{c}{r_{e}\omega}\right)^{0.3}\!\alpha^{0.1}\,\frac{\kappa^{0.5}}{\kappa_{\rm u}^{0.2}}\approx 10^{4}\,\frac{C^{1/4}\kappa_{\rm u}^{0.3}L_{\rm w,37}^{1/20}}{\nu_{9}^{0.4}r_{11}^{0.1}}.$ (109)
+$\tilde{\gamma}_{\rm stoch}\approx\left[\frac{2r_{e}{\cal E}}{mc\omega r^{2}}\right]^{3/7}\!\!b^{1/7}\approx 10^{5}\,\frac{{\cal E}_{39}^{3/7}L_{\rm w,37}^{1/14}\,C^{1/7}}{\nu_{9}^{4/7}r_{11}},$ (107)
+where ${\cal E}=4\pi r^{2}\Sigma=\int 4\pi r^{2}Uc\,d\xi$ is the energy of the radio wave (isotropic equivalent), and we substituted $b=Cb_{\rm u}$ (Equation 99) and used Equation (86) for $b_{\rm u}$.
+One can see that $\tilde{\gamma}$ gained from stochastic heating exceeds the regular oscillation $\tilde{\gamma}\sim a$ by a factor of $\sim 10^{3}$.
+Equation (107) holds if radiative losses of the heated plasma are negligible.
+Energy losses in the fluid frame $mc^{2}\tilde{\dot{\gamma}}_{e}=\Delta\tilde{\cal E}_{e}/\Delta\tilde{t}$ are given by Equation (D16), which we rewrite as
+$\tilde{\dot{\gamma}}_{e}=\frac{3\sigma_{\rm T}\tilde{U}}{2mc}\,\tilde{\gamma}^{2}.$ (108)
+The losses offset the heating rate (Equation 15) when $\tilde{\gamma}$ approaches
+$\tilde{\gamma}_{\rm rad}\approx\left(\frac{c}{r_{e}\tilde{\omega}}\right)^{\!0.3}\!b^{0.1}\approx 10^{4}\,\frac{C^{1/4}\kappa_{\rm u}^{0.3}L_{\rm w,37}^{1/20}}{\nu_{9}^{0.4}r_{11}^{0.1}},$ (109)
+where we used $b\approx Cb_{\rm u}$ and $\tilde{\omega}=\omega/\kappa=(\omega/\kappa_{\rm u}\sqrt{C})$.
 In summary, stochastic heating gives
@@
 $r<\frac{6\times 10^{11}}{\gamma_{\rm u}^{1/3}}\,\frac{{\cal E}_{39}^{10/21}L_{\rm w,37}^{1/42}}{C_{2}^{5/42}\,\nu_{9}^{4/21}}\,{\rm cm}.$ (111)
-9.2 Plasma compression and trapping by the FRB
-Once the radio wave begins to propagate through the plasma, it immediately induces significant compression $C$ (section 6.1), which will be denoted below as $C_{0}$.
-When propagation occurs in the zone of stochastic heating $r<r_{\rm stoch}$, the immediate compression $C_{0}$ turns out so large that further growth of $C>C_{0}$ on timescales $t\gg t_{\rm cross}$ (front relaxation toward the steady state) is irrelevant, because $t_{\rm cross}$ becomes too long.
-In general, large $C$ corresponds to large $\gamma_{\rm D}\approx C^{1/2}\gamma_{\rm u}$ and implies a long timescale for the plasma to cross the wave packet $t_{\rm cross}\sim\gamma_{\rm D}^{2}T$ (Equation 69).
-When $t_{\rm cross}$ reaches $t=r/c$, the plasma becomes trapped in the packet, surfing the radio wave.
-The ratio $t/t_{\rm cross}$ may be written as
-$\frac{t_{\rm cross}}{t}\sim\frac{cT\gamma_{\rm D}^{2}}{r}=\frac{\gamma_{\rm D}^{2}}{\gamma_{\rm trap}^{2}}=\frac{C}{C_{\rm trap}},$ (112)
+IX.3 Plasma trapping by the FRB
+When FRB propagation through the wind occurs in the zone of stochastic heating $r<r_{\rm stoch}$, the plasma acceleration and compression turn out so large that it becomes trapped inside the radio wave packet.
+In general, large compression $C$ corresponds to large $\gamma_{\rm D}\approx C^{1/2}\gamma_{\rm u}$ (Equation 97) and implies a long timescale for the plasma to cross the wave packet $t_{\rm cross}\sim\gamma_{\rm D}^{2}T$ (Equation 60).
+If $t_{\rm cross}$ exceeds $t=r/c$, the plasma becomes trapped, surfing the radio wave.
+The ratio $t_{\rm cross}/t$ may be written as
+$\frac{t_{\rm cross}}{t}\sim\frac{cT\gamma_{\rm D}^{2}}{r}\approx\frac{\gamma_{\rm D}^{2}}{\gamma_{\rm trap}^{2}}\approx\frac{C}{C_{\rm trap}},$ (112)
 where
-$\gamma_{\rm trap}\approx\left(\frac{r}{cT}\right)^{1/2},\qquad C_{\rm trap}\approx\frac{r}{cT\gamma_{\rm u}^{2}}.$ (113)
-Let us now evaluate the compression factor $C_{0}$ that immediately develops in the FRB-wind interaction, and show that it easily reaches $C_{\rm trap}$.
-The calculation was set up in section 6.1 using the frame where the upstream plasma is at rest.
-Here, we will denote this frame ${\cal K}^{\prime}$ to distinguish it from our lab frame; quantities measured in frame ${\cal K}^{\prime}$ will be denoted with a prime.
-We are interested in the strong compression regime $\tilde{C}_{0}\gg 1$.
-Then, Equation (66) gives
-$\tilde{C}_{0}\approx\kappa^{\prime}\approx\frac{3\tilde{\gamma}}{2\sigma_{\rm u}}+\sqrt{\frac{\epsilon_{\rm rad}^{\prime}}{\sigma_{\rm u}}},$ (114)
-where we used $w\approx(3/2)\tilde{\gamma}$.
-A proper compression factor $\tilde{C}$ corresponds to $C\approx\tilde{C}^{2}$ in the lab frame (Equation 98), and radiative losses in the lab frame are ${\cal E}_{\rm rad}\approx\kappa_{\rm u}{\cal E}_{\rm rad}^{\prime}$.
-So, one can use two general relations
-$\kappa^{\prime}\approx C^{1/2},\qquad\epsilon_{\rm rad}^{\prime}\approx\frac{\epsilon_{\rm rad}}{\kappa_{\rm u}}.$ (115)
-Radiative losses of a particle interacting with the radio wave packet are given by Equation (38) with an extra factor of 2 in the stochastic heating regime (B22). Thus, we find
+$\gamma_{\rm trap}\approx\left(\frac{r}{cT}\right)^{1/2},\quad C_{\rm trap}\approx\frac{r}{cT\gamma_{\rm u}^{2}}\approx\frac{30\,r_{11}}{T_{\rm ms}(\gamma_{\rm u}/10)^{2}}.$ (113)
+In the regular oscillation zone $r>r_{\rm stoch}$, trapping does not occur: $C_{\max}(r)$ is everywhere below $C_{\rm trap}(r)$.
+Note that $C_{\max}$ is a decreasing function of radius while $C_{\rm trap}$ increases with radius, so the ratio $C_{\max}/C_{\rm trap}$ is largest at the inner boundary of the regular oscillation zone $r_{\rm stoch}$.
+It is easy to verify that $C_{\max}(r_{\rm stoch})=C_{\rm stoch}<C_{\rm trap}(r_{\rm stoch})$, where $C_{\rm stoch}$ is given by Equation (103).
+By contrast, in the stochastic heating zone $r<r_{\rm stoch}$, the FRB deposits enormous energy and momentum into the plasma, resulting in huge compression $C$.
+It would greatly exceed $C_{\rm trap}$ if one assumes that the plasma crosses the wave packet.
+[^12] A lower limit on the proper compression factor $\tilde{C}\approx C^{1/2}\approx\tilde{\kappa}/\tilde{\kappa}_{\rm u}$ may be estimated as $\tilde{\gamma}/\sigma_{\rm u}$, which gives $C\gg C_{\rm trap}$.
+So, the plasma cannot cross the packet at $r<r_{\rm stoch}$; it gets stuck inside the packet at some $\xi_{\rm trap}<T$.
+The trapping regime holds unless the FRB itself becomes strongly damped by its interaction with the wind.
+Damping allows the plasma to move through the wave packet while consuming its energy until $\xi_{\rm trap}$ approaches $T$ and the entire FRB loses its energy.
+The parameter space where this happens is evaluated below.
+X FRB energy losses
+Consider an FRB wave packet overtaking the wind and accelerating it from Lorentz factor $\gamma_{\rm u}$ to some $\gamma_{\rm D}$.
+Acceleration by the factor $\gamma_{\rm u}/\gamma_{\rm D}\approx\kappa_{\rm u}/\kappa\gg 1$ is accompanied by compression given in Equation (97).
 […]
-where we used $dt=d\xi/(1-\beta_{\rm D})$ and $\kappa^{2}(1-\beta_{\rm D})=1+\beta_{\rm D}$.
-Note that this expression involves the FRB energy ${\cal E}$, assuming that the particle interacts with the entire radio packet, i.e.
-$t_{\rm cross}\leq r/c$, which corresponds to $C\leq C_{\rm trap}$.
-[^8] A “deep” trapping regime $t_{\rm cross}\gg r/c$ is possible if the plasma is trapped in a small leading part of the wave packet, at $\xi_{\rm trap}\ll T$.
-In this case, ${\cal E}$ in Equation (116) should be replaced by the part of the FRB energy ${\cal E}_{\rm trap}$ contained in the leading part $0<\xi<\xi_{\rm trap}$.
-Self-consistent $\xi_{\rm trap}$ and ${\cal E}_{\rm trap}$ depend on the packet shape.
-Using Equations (115) and (116) in Equation (114), we find
-$C_{0}^{1/2}\sim\frac{\tilde{\gamma}}{\sigma_{\rm u}}\left(\frac{3}{2}+\psi\right),\qquad\psi\equiv\sqrt{\frac{\sigma_{\rm T}{\cal E}\sigma_{\rm u}}{4\pi r^{2}mc^{2}\kappa_{\rm u}}}.$ (117)
-Here, $\tilde{\gamma}=\min\{\tilde{\gamma}_{\rm stoch},\tilde{\gamma}_{\rm rad}\}$ depends on $C=C_{0}$ (section 9.1), and one can find the consistent solution for $\tilde{\gamma}$ and $C_{0}$:
-$\tilde{\gamma}=\left\{\begin{array}[]{lr}\tilde{\gamma}_{\rm stoch}&\quad\psi<0.7\\ \tilde{\gamma}_{\rm rad}&\quad\psi>0.7\end{array}\right.$ (118)
-$C_{0}\sim\left[\frac{c\kappa_{\rm u}}{r_{e}\omega}\right]^{6/5}\!\!\frac{\alpha^{2/5}}{16\sigma_{\rm u}^{4}}\!\times\!\left\{\begin{array}[]{lr}(3\psi)^{12/5}(3+2\psi)^{14/5}&\psi<0.7\\ (3+2\psi)^{4}&\psi>0.7\end{array}\right.$ (119)
-with the parameter $\alpha$ given by Equation (88).
-The obtained solution holds if it gives $C_{0}\gg 1$, and this condition is satisfied at all $r<r_{\rm stoch}$.
-Furthermore, $C_{0}$ exceeds $C_{\rm trap}$.
-Indeed, consider the smallest $C_{0}^{1/2}\sim 3\tilde{\gamma}_{\rm stoch}/2\sigma_{\rm u}$ (Equation 117), which is found in the non-radiative regime $\psi\ll 0.7$ (small ${\cal E}$).
-Then, using Equation (110) for $\tilde{\gamma}_{\rm stoch}$, we find
-$C_{0}\sim\frac{10^{6}\,{\cal E}_{39}^{6/5}L_{\rm w,37}^{1/5}}{\nu_{9}^{8/5}\eta_{4}^{14/5}R_{\rm LC,10}^{14/5}}\!\!\times\!\left\{\begin{array}[]{lr}1&r<R_{\rm sat}\\ (R_{\rm sat}/r)^{14/5}&r>R_{\rm sat}\end{array}\right.$ (120)
-where $R_{\rm sat}=\eta^{1/3}R_{\rm LC}\approx 2\times 10^{11}\,\eta_{4}^{1/3}R_{\rm LC,10}$ cm is the radius where the wind Lorentz factor $\gamma_{\rm u}$ approaches $\eta^{1/3}$ (Equation 85).
-It is easy to verify that $C_{0}>C_{\rm trap}$ at all $r<r_{\rm stoch}$.
-So, FRBs emitted in the wind at $r<r_{\rm stoch}$ immediately trap the wind plasma.
-Two aspects of the trapping regime should be noted:
-(1) The large value of $C_{0}$ is found using the full FRB energy ${\cal E}\sim LT$ in the parameter $\psi$.
-When $C_{0}\gg C_{\rm trap}$, the plasma becomes trapped in a small leading part of the radio wave packet with ${\cal E}_{\rm trap}\ll{\cal E}$ (see footnote 8), avoiding interaction with its main part.
-(2) The trapping regime holds unless the FRB becomes strongly damped by its interaction with the wind.
-Damping allows the plasma to move through the wave packet while consuming its energy, avoiding trapping.
-The parameter space where FRBs are damped is evaluated below.
-10 FRB energy losses
-Consider first FRB energy losses $\Delta{\cal E}^{\prime}$ in frame ${\cal K}^{\prime}$ where the upstream plasma is at rest. The corresponding momentum losses are $\Delta{\cal E}^{\prime}/c$.
-Let the radio wave packet interact with $\Delta{\cal N}$ particles.
-The momentum lost by the radio wave is the sum of momentum gained by the MHD flow and momentum waisted by the energized plasma particles to radiation,
-$\frac{\Delta{\cal E}^{\prime}}{c}=\Delta{\cal N}\,\frac{T^{t^{\prime}r^{\prime}}_{\rm MHD}}{cn^{\prime}}+\beta_{\rm D}^{\prime}\frac{{\cal E}_{\rm rad}^{\prime}}{c}.$ (121)
-The radiated energy per particle is $\epsilon_{\rm rad}^{\prime}mc^{2}\equiv{\cal E}_{\rm rad}^{\prime}/\Delta{\cal N}$. This gives
-$\frac{\Delta{\cal E}^{\prime}}{mc^{2}\Delta{\cal N}}\approx\beta_{\rm D}^{\prime}\left[\gamma_{\rm D}^{\prime}(1+w+\sigma_{\rm u}\kappa^{\prime})+\epsilon_{\rm rad}^{\prime}\right].$ (122)
-Note that $\beta_{\rm D}^{\prime}=(\kappa^{\prime 2}-1)/(\kappa^{\prime 2}+1)$ and $\kappa^{\prime}\approx\tilde{C}$, where $\tilde{C}=\tilde{\rho}/\tilde{\rho}_{\rm u}$ is the proper compression factor.
-Thus, Equation (122) yields
-$\frac{\Delta{\cal E}^{\prime}}{mc^{2}\Delta{\cal N}}\approx(\tilde{C}^{2}-1)\left(\frac{1+w}{2\tilde{C}}+\frac{\sigma_{\rm u}}{2}+\frac{\epsilon_{\rm rad}^{\prime}}{\tilde{C}^{2}+1}\right).$ (123)
-Proper compression $\tilde{C}$ is related to compression $C$ in the lab frame by $C\approx\tilde{C}^{2}$ (Equation 98).
-Wave energy in the lab frame is ${\cal E}=\kappa_{\rm u}{\cal E}^{\prime}$, so $\Delta{\cal E}^{\prime}=\Delta{\cal E}/\kappa_{\rm u}$.
-Radiative losses in the lab frame and frame ${\cal K}^{\prime}$ are also related by ${\cal E}_{\rm rad}/{\cal E}_{\rm rad}^{\prime}\approx\kappa_{\rm u}$.
-Using these relations, one can rewrite Equation (123) in terms of the lab-frame quantities:
-$\frac{\Delta{\cal E}}{mc^{2}\Delta{\cal N}\kappa_{\rm u}}\approx(C-1)\left[\frac{1+w}{2\sqrt{C}}+\frac{\sigma_{\rm u}}{2}+\frac{\epsilon_{\rm rad}}{(C+1)\kappa_{\rm u}}\right].$ (124)
+$\Delta{\cal E}\approx\Delta{\cal N}\,\frac{T^{tt}_{\rm MHD}}{n}+\Delta{\cal N}{\cal E}_{e},$ (114)
+where ${\cal E}_{e}$ is the radiated energy per particle.
+In this relation, we have neglected the initial energy of the wind compared to its gain from the FRB.
+Using $T_{\rm MHD}^{tt}=\tilde{\rho}c^{2}h\gamma_{\rm D}^{2}-P$ and dropping the term $P$ (it is smaller than the main term by the factor $\sim\gamma_{\rm D}^{-2}\ll 1$) one finds
+$\frac{T^{tt}_{\rm MHD}}{n}\approx mc^{2}\gamma_{\rm D}(1+w+\sigma).$ (115)
+Radiative losses of a particle interacting with the radio wave packet are derived in Appendix D in the regimes of regular oscillation and stochastic heating.
+For the MHD flow with $\gamma_{\rm D}\gg 1$ and $\tilde{\gamma}\gg 1$ they simplify to
+$\frac{\Delta{\cal E}_{e}}{\Delta t}\approx\frac{c\sigma_{\rm T}U\tilde{\gamma}^{2}}{2\kappa^{2}}.$ (116)
+Then, we find
+${\cal E}_{e}=\!\int\!\frac{\Delta{\cal E}_{e}}{\Delta t}\,dt\approx\!\int\!\frac{c\sigma_{\rm T}\tilde{\gamma}^{2}Ud\xi}{4}\approx\frac{\sigma_{\rm T}\tilde{\gamma}^{2}{\cal E}}{16\pi r^{2}},$ (117)
+[…]
+Using $\gamma_{\rm D}/\gamma_{\rm u}\approx\sqrt{C}\gg 1$ and $\sigma/\sigma_{\rm u}=\tilde{\rho}/\tilde{\rho}_{\rm u}\approx\sqrt{C}$ (Equation 97), we obtain
+$\frac{\Delta{\cal E}}{\Delta{\cal N}mc^{2}\gamma_{\rm u}}\approx\sqrt{C}\,w+C\sigma_{\rm u}+\frac{\sigma_{\rm T}{\cal E}\,\tilde{\gamma}^{2}}{16\pi r^{2}mc^{2}\gamma_{\rm u}}.$ (118)
 As the FRB propagates to a radius $r$, it interacts with $\Delta{\cal N}$ wind particles given by
 $\Delta{\cal N}\approx(1-\beta_{\rm u})\frac{r}{c}\dot{\cal N}\approx\frac{2rL_{\rm w}}{mc^{3}\eta\kappa_{\rm u}^{2}}.$ (119)
-Equations (124) and (125) determine the FRB energy losses $\Delta{\cal E}$ anywhere in the wind, expressed as a function of the compression factor $C$.
-Enthalpy $w\approx(3/2)\tilde{\gamma}$ is controlled by the heating regime: $\tilde{\gamma}=(1+a^{2})^{1/2}$ at $r>r_{\rm stoch}$, and $\tilde{\gamma}$ is given by Equation (110) at $r<r_{\rm stoch}$.
+Equations (118) and (119) determine the FRB energy loss $\Delta{\cal E}$, expressed as a function of the compression factor $C$.
 It is easy to verify that losses are negligible, $\Delta{\cal E}\ll{\cal E}$, in the zone of regular heating $r>r_{\rm stoch}$.
 Thus, FRB damping in the wind may only occur in the zone of stochastic heating, $r<r_{\rm stoch}$.
-Using $w\approx(3/2)\tilde{\gamma}\gg 1$ and $C\gg 1$, we find from Equation (124):
-$\Delta{\cal E}\approx\frac{rL_{\rm w}}{2c\,\eta\gamma_{\rm u}}\left[\frac{3}{2}\tilde{\gamma}\sqrt{C}+\sigma_{\rm u}C+\frac{\sigma_{\rm T}\tilde{\gamma}^{2}{\cal E}}{4\pi r^{2}mc^{2}\gamma_{\rm u}}\right].$ (126)
-For an FRB with luminosity $L$ and energy ${\cal E}\sim LT$, we find $\tilde{\gamma}(r,C)$ from Equation (110) at radii $r<r_{\rm stoch}$, and then find $\Delta{\cal E}(r,C)$ from Equation (126).
-The FRB is strongly damped if $\Delta{\cal E}$ approaches ${\cal E}$.
-Damping begins in the leading part of the FRB $\xi\ll T$, as the plasma entering the radio wave packet consumes its local energy ${\cal E}(\xi)\sim 4\pi r^{2}U(\xi)\,c\,\xi$ and makes the deep trapping regime $C\gg C_{\rm trap}$ unsustainable.
-Damping succeeds in killing the FRB when it enables the plasma to keep going in $\xi$ and crossing the entire wave packet, which corresponds to $\Delta{\cal E}\sim{\cal E}$ and marginal trapping, $C\approx C_{\rm trap}$.
-Therefore, the condition for marginal damping is
+Then, $w\approx(3/2)\tilde{\gamma}$ and Equation (118) gives
+$\Delta{\cal E}\approx\frac{rL_{\rm w}}{2c\,\eta\gamma_{\rm u}}\left[\frac{3}{2}\tilde{\gamma}\sqrt{C}+\sigma_{\rm u}C+\frac{\sigma_{\rm T}\tilde{\gamma}^{2}{\cal E}}{16\pi r^{2}mc^{2}\gamma_{\rm u}}\right].$ (120)
+[…]
+Figure 10: Regimes of FRB propagation through the wind ($r>R_{\rm LC}$), shown on the $r$-$L$ plane, for FRBs with frequency $\nu=1$ GHz and duration $T=1$ ms.
+The two plots assume the wind energy parameter $\eta=10^{4}$ (left) and $\eta=10^{3}$ (right).
+Magenta region $r<R_{\rm LC}$ indicates the closed magnetosphere; FRBs emitted in this region are known to experience strong radiative damping (Beloborodov, 2024).
+FRBs are also damped in the wind zone $r>R_{\rm LC}$ studied in the present paper, in particular in the deep purple region.
+Their energy is consumed by the stochastic heating and compression of the wind plasma, with radiative losses.
+Colored dots indicate the dominant damping channel: radiative losses ${\cal E}_{\rm rad}$ (red dots), energy kept by the plasma ${\cal E}_{\rm p}$ (cyan dots), and energy spent to compress the wind magnetic field ${\cal E}_{\rm mag}$ (yellow dots).
+Stochastic heating occurs at $r<r_{\rm stoch}$.
+In the light purple region, FRBs trap the wind plasma, carrying it in a leading part of the wave packet and avoiding significant energy losses.
+In the olive zone, $r_{\rm stoch}<r<r_{1}$, FRBs induce regular particle oscillations with $\tilde{\gamma}\gg 1$ and drive a compression wave through the wind, with a decreasing compression factor $C$.
+At radii $r>r_{\star}$, the compression wave is locally quasisteady and $C=C_{\star}\approx 1+a^{2}$.
+The FRB strength parameter $a$ drops below unity outside $r_{1}$ (grey region).
+[…]
 $\frac{\Delta{\cal E}(r,C_{\rm trap})}{{\cal E}}\approx 1.$ (121)
-It defines the boundary of the parameter space where FRBs are strongly damped.
-One can show that $\Delta{\cal E}(r,C_{\rm trap})/{\cal E}$ is a decreasing function of ${\cal E}$, so strong damping occurs for FRBs with energies ${\cal E}$ below a critical value ${\cal E}_{\rm esc}(r)$ defined by Equation (127) and the corresponding critical luminosity is $L_{\rm esc}\sim{\cal E}_{\rm esc}/T$.
-The damping region on the $r$-$L$ plane is shown in Figure 8.
-Figure 8: Regimes of FRB propagation through the wind ($r>R_{\rm LC}$), shown on the $r$-$L$ plane, for FRBs with emission frequency $\nu=1$ GHz and duration $T=1$ ms.
-The two plots assume the wind energy parameter $\eta=10^{4}$ (left) and $\eta=10^{3}$ (right).
-The magenta region $r<R_{\rm LC}$ indicates the closed magnetosphere.
-FRBs emitted inside the magnetosphere experience strong radiative damping (Beloborodov 2024).
-FRBs emitted at radii $R_{\rm LC}<r<r_{\rm stoch}$ induce fast stochastic heating and compression of the wind.
-In the light purple region, FRBs trap the wind plasma, carrying it in a leading part of the wave packet and avoiding significant energy losses.
-In the deep purple region, FRBs are damped: their energy is consumed by heating and compressing the wind plasma, with radiative losses.
-The colored dots indicate the dominant damping channel: radiative losses ${\cal E}_{\rm rad}$ (red dots), energy kept by the plasma ${\cal E}_{\rm p}$ (cyan dots) and energy spent to compress the wind magnetic field ${\cal E}_{\rm mag}$ (yellow dots).
-In the olive zone, $r_{\rm stoch}<r<r_{1}$, the wind is still heated to a relativistic enthalpy $w\gg 1$; however, here heating occurs in the regular regime (particles oscillate in the radio wave instead of stochastically gaining energy) with the moderate $w\approx(3/2)a$ and a weaker compression.
-At radii $r_{\star}<r<r_{1}$, the compression factor is $C=C_{\star}\approx 1+a_{0}^{2}$.
-The FRB strength parameter $a=a_{0}/\sqrt{2}$ drops below unity outside $r_{1}$ (grey region).
+[…]
 It is useful to examine more closely the loss $\Delta{\cal E}$ stated in Equation (120). The three terms in the square brackets correspond to three contributions,
@@
 where ${\cal E}_{\rm p}\propto\tilde{\gamma}$ is the energy kept by the plasma particles, ${\cal E}_{\rm mag}\propto\sigma_{\rm u}$ is the energy used to compress the wind magnetic field, and ${\cal E}_{\rm rad}\propto\tilde{\gamma}^{2}$ is the radiative losses.
-Let us compare radiative losses ${\cal E}_{\rm rad}$ with ${\cal E}_{\rm p}$ and ${\cal E}_{\rm mag}$.
+The largest of these three contributions defines the main damping channel for the FRB; it is indicated in Figure 10.
 It is easy to see that ${\cal E}_{\rm rad}>{\cal E}_{\rm p}$ if $\tilde{\gamma}_{\rm rad}<\tilde{\gamma}_{\rm stoch}$; in this regime, the plasma radiates more energy than it keeps.
-This occurs at radii that satisfy the condition given in Equation (111).
-With $C=C_{\rm trap}$, it can be stated as
-${\cal E}_{\rm rad}>{\cal E}_{\rm p}:\quad r<2\times 10^{11}\,\frac{{\cal E}_{39}^{20/47}T_{\rm ms}^{5/47}L_{\rm w,37}^{1/47}}{\gamma_{\rm u}^{4/47}\nu_{9}^{8/47}}\,{\rm cm}.$ (129)
-The ratio ${\cal E}_{\rm rad}/{\cal E}_{\rm mag}$ equals the ratio of the third and second terms in the brackets in Equation (126).
-Using $C=C_{\rm trap}$ and $\tilde{\gamma}=\tilde{\gamma}_{\rm rad}$ (i.e.
-assuming ${\cal E}_{\rm rad}>{\cal E}_{\rm p}$), we find
-${\cal E}_{\rm rad}>{\cal E}_{\rm mag}:\quad r<10^{11}\gamma_{\rm u}^{16/27}\frac{{\cal E}_{39}^{10/27}T_{\rm ms}^{5/27}L_{\rm w,37}^{1/27}}{\eta_{4}^{10/27}\nu_{9}^{8/27}}\,{\rm cm}.$ (130)
-${\cal E}_{\rm rad}$ dominates the FRB losses at radii that satisfy both conditions stated in Equations (129) and (130).
-11 Summary
-The basic physics problem studied in this paper is how a strong ($a_{0}\gg 1$) electromagnetic wave packet accelerates an MHD fluid.
-For weak electromagnetic waves ($a_{0}\ll 1$), the acceleration effect is well known as the radiation pressure force, used in astrophysics e.g.
-to define the Eddington limit.
-For strong waves, the acceleration effect was previously calculated in unmagnetized plasmas; the $e^{\pm}$ plasma then becomes compressed in the wave packet by the factor $C\approx 1+a_{0}^{2}$.
-A similar problem for a magnetized plasma was poorly understood.
-The present paper developed a method to solve this problem, obtained sample numerical solutions, and provided a physical picture of the wave-plasma interaction.
-Two key facts help solve the problem: (1) the wave strength parameter $a_{0}$ controls the plasma enthalpy as shown in section 3, and (2) quantity ${\cal Q}_{-}={\cal E}-c{\cal P}$ (where ${\cal E}$ is energy and ${\cal P}$ is momentum) is conserved in the wave-plasma interaction because the wave has ${\cal Q}_{-}=0$ and cannot transfer $\Delta{\cal Q}_{-}\neq 0$ to the MHD fluid.
-A first estimate would suggest that plasma with a sufficiently large magnetization $\sigma$ is not strongly compressed by the electromagnetic wave, since the deposited momentum is proportional to the small plasma density while compression requires work to be done against the large background magnetic pressure (Beloborodov 2021; see also section 6.1).
+This occurs at radii that satisfy the approximate condition stated in Equation (111) with $C=C_{\rm trap}$.
+XI Summary
+[…]
+A first estimate would suggest that plasma with a sufficiently large magnetization $\sigma$ is not strongly compressed by the radio wave, since the deposited momentum is proportional to the small plasma density while compression requires work to be done against the large background magnetic pressure (Beloborodov 2021; see also section VI).
 The actual solution demonstrates that this expectation holds initially, when the wave packet begins to interact with the plasma.
-Later, the propagating wave gradually develops a large compression factor $C$ due to a curious “memory” effect described in section 6.
-Given a sufficient time $t_{\star}\propto C\sigma$ the compression front relaxes to a steady state with $C=C_{\star}\approx 1+a_{0}^{2}$ independent of $\sigma$.
-However, large $C_{\star}$ (implied by large $a_{0}$) is not reached if $t_{\star}$ exceeds the wave propagation time.
-Besides the peculiar gradual evolution of the propagating wave packet, the background magnetic field brings another special feature that dramatically changes the wave-plasma interaction: compression increases the gyro-frequency and triggers strong stochastic heating of the plasma by the electromagnetic wave.
-The heating boosts momentum deposition, enhancing acceleration and compression of the MHD fluid in the wave packet.
-We have used the obtained results to examine FRB interaction with a magnetized wind flowing from the central object (a neutron star). Our conclusions are summarized below and illustrated in Figure 8.
-FRBs propagating at radii $R_{\rm LC}<r<r_{\rm stoch}$ have a huge effect on the wind medium.
-They heat the wind up to an average Lorentz factor $\tilde{\gamma}\sim 10^{3}a_{0}$ (measured in the fluid rest frame), and accelerate it to a bulk Lorentz factor exceeding $\gamma_{\rm trap}\sim(r/cT)^{1/2}$, which implies trapping of the wind plasma inside the radio wave packet.
-The energy per particle in the heated plasma $\gamma_{\rm trap}\tilde{\gamma}mc^{2}$ (measured in the static lab frame) is in the TeV range.
-Thus, the FRB-wind interaction converts the radio-wave energy into high energy particles, which experience radiative losses.
-A fraction of FRB energy is also spent to compress the wind magnetic field.
-Only sufficiently bright FRBs survive the energy losses (Figure 8).
-The FRB-wind interaction changes at larger radii: at $r=r_{\rm stoch}$ (Equation 104) there is a sharp transition from stochastic heating to regular heating (particle oscillation in the wave), which sustains a moderate $\tilde{\gamma}\sim a_{0}$.
-Energy losses of the radio wave are negligible at $r>r_{\rm stoch}$.
-FRBs emitted in this zone still induce strong bulk acceleration of the wind plasma.
-Here, instead of trapping the plasma, the FRB drives an MHD compression wave in the wind.
-At $r>r_{\star}$ (Equation 95), the compression wave is quasisteady ($r/c\gg t_{\star}$), and the compression factor equals $C_{\star}\approx 1+a_{0}^{2}$.
+If the propagating radio wave is given sufficient time, it gradually develops a much larger compression factor $C$.
+After time $t_{\star}\propto C\sigma$ the compression front relaxes to a steady state with $C=C_{\star}\approx 1+a^{2}$, which is independent of $\sigma$.
+Besides the peculiar gradual evolution of the propagating compression front, the background magnetic field brings another special feature: it enables powerful stochastic heating of the plasma by the radio wave, in particular when compression increases the local gyro-frequency.
+The heating boosts momentum deposition, further enhancing acceleration and compression of the MHD fluid in the wave packet.
+We have used the obtained results to examine FRB interaction with a magnetized wind flowing from the neutron star. Our conclusions are summarized below and illustrated in Figure 10.
 […]
-The results summarized in Figure 8 imply interesting constraints on FRB emission models.
-In particular, they exclude scenarios that place the radio source at radii $r<r_{\rm damp}\sim 10^{11}$ cm and rely on free FRB propagation through the magnetar wind.
-Thus, the previously found damping zone $r<R_{\rm LC}$ (Beloborodov, 2021, 2024; Golbraikh & Lyubarsky, 2023) is extended to $r_{\rm damp}\sim 10^{11}$ cm.
-The exact value of $r_{\rm damp}$ varies around $10^{11}$ cm for the relevant values of wind parameter $\eta=L_{\rm w}/\dot{\cal N}mc^{2}$ around $10^{4}$.
+The FRB-wind interaction changes at larger radii: at $r=r_{\rm stoch}\sim 10^{12}$ cm (Equation 105) there is a sharp transition from stochastic heating to regular particle oscillations in the radio wave, with a moderate Lorentz factor $\tilde{\gamma}\sim a$.
+FRB energy losses are negligible at $r>r_{\rm stoch}$.
+FRBs emitted in this zone still induce strong bulk acceleration of the wind plasma, but no longer trap it, and continue to drive an MHD compression wave in the wind.
+The compression wave becomes locally quasisteady at $r>r_{\star}\sim 10^{13}L_{42}^{1/3}$ cm (Equation 93), with compression factor following the simple expression $C_{\star}\approx 1+a^{2}$.
+In this region, our results agree with Sobacchi et al.
+(2024).
+It should be emphasized that we find the compression factor $C\ll C_{\star}$ at $r\ll r_{\star}$, where the timescale for the compression front relaxation toward $C_{\star}$ exceeds the FRB propagation time $r/c$.
+Therefore, our results do not support the suggestion of Sobacchi et al.
+(2024) that FRBs emitted at $r\lesssim 10^{12}$ cm lose energy to the huge compression $C_{\star}$ of the magnetized wind.
+[^14] Note also that energy consumption implied by $C_{\star}\approx 1+a^{2}$ is reversible: when the MHD fluid exits behind the wave packet (where $C_{\star}$ comes back to unity) it returns energy to the FRB.
+Our results still imply a significant constraint on the FRB emission radius (Figure 10).
+It is caused by the enormous stochastic heating of the wind exposed to the FRB at small $r$, which leads to catastrophic energy losses of the FRB.
+As a result, the previously found damping zone $r<R_{\rm LC}$ (Beloborodov, 2021, 2024; Golbraikh & Lyubarsky, 2023) is now extended to $r_{\rm damp}\sim 10^{11}$ cm.
+This will challenge any FRB scenarios that place radio emission at $r<10^{11}$ cm and rely on its escape through the magnetar wind.
+The exact $r_{\rm damp}$ depends on the wind parameter $\eta=L_{\rm w}/\dot{\cal N}mc^{2}$, which may vary around $\eta\sim 10^{4}$.
 We find that $r_{\rm damp}$ slowly decreases with increasing $\eta>10^{4}$.
-The compression and damping effects are particularly strong in winds with $\eta<10^{4}$.
+FRB damping is particularly strong in winds with $\eta<10^{4}$.
 […]
-In addition to plasma acceleration and compression studied in this paper, the ambient medium may be strongly affected by the nonlinear development of self-modulation instabilities of the radio wave (Sobacchi et al., 2023), investigated so far for waves with $a_{0}<1$.
-Instabilities in the wind zone with $a_{0}>1$ are poorly understood and will require future work.
-The author acknowledges support by NASA grants 21-ATP21-0056 and 80NSSC24K0282, NSF grant AST-2408199, and Simons Foundation grant No. 446228.
+In addition to plasma acceleration and compression by FRBs, the ambient medium may be strongly affected by self-modulation instabilities of the radio wave (Sobacchi et al., 2023), investigated so far for waves with $a<1$.
+Instabilities in the wind zone with $a>1$ are poorly understood and will require future work.
+After this work was finished, a paper by Sobacchi (2025) appeared on arxiv, discussing conditions for stochastic heating and its interpretation as synchrotron absorption calculated by Lyubarsky (2018).
+Within a numerical factor, their conditions are consistent with both Figure 3 of the present paper and the condition $\gamma\gtrsim\gamma_{\star}$ stated in B22.
+Links between stochastic heating (Equation 15) and the absorption rate calculated by Lyubarsky (2018) are explained in Appendix B.2.
+I thank the Kavli Institute for Theoretical Physics for hospitality during the plasma astrophysics program in summer 2025, where this work (including Figures 3 and 9) was presented and discussed with the program participants.
+I thank Yuri Lyubarsky for drawing my attention to his paper on synchrotron absorption of strong waves, which stimulated the discussion in Appendix B.2.
+The author acknowledges support by NASA grants 21-ATP21-0056 and 80NSSC24K0282, NSF grant AST-2408199, and Simons Foundation grant 446228.
+This work was also facilitated by Multimessenger Plasma Physics Center (MPPC) grant PHY-2206609.
+Appendix A Speed of the radio wave packet
+For FRBs propagating through the wind outside the magnetosphere (section VII), the relevant propagation regime is such that particle oscillation in the radio wave occurs faster than Larmor rotation in the local magnetic field: $\tilde{\omega}>\tilde{\omega}_{\rm L}$ (where tilde indicates that the quantity is measured in the local fluid frame $\tilde{\cal K}$).
+Then, the magnetic field weakly affects the FRB propagation speed.
+The plasma response to the radio wave (i.e.
+the excited electric current) depends on the local $e^{\pm}$ density $\tilde{n}$ and the effective particle mass $\tilde{\gamma}m$, where $\tilde{\gamma}$ is the average particle Lorentz factor in the fluid frame.
+The radio wave then satisfies the standard dispersion relation, which can be stated in the fluid frame as
+$\tilde{\omega}^{2}(\tilde{k})\approx c^{2}\tilde{k}^{2}+\tilde{\omega}_{\rm p}^{2}\qquad(\tilde{\omega}\gg\tilde{\omega}_{\rm L}),$ (A1)
+where
+$\tilde{\omega}_{\rm p}^{2}=\frac{4\pi e^{2}\tilde{n}}{\tilde{\gamma}m}.$ (A2)
+The FRB wave packet propagates relative to the plasma with group speed $\tilde{v}_{\rm gr}=d\tilde{\omega}/d\tilde{k}=c^{2}\tilde{k}/\tilde{\omega}$. It can also be characterized by Lorentz factor
+$\tilde{\gamma}_{\rm gr}=\left(1-\frac{\tilde{v}_{\rm gr}^{2}}{c^{2}}\right)^{-1/2}\!=\frac{\tilde{\omega}}{\tilde{\omega}_{\rm p}}\gg 1.$ (A3)
+All models considered below satisfy $\tilde{\omega}\gg\tilde{\omega}_{\rm p}$, so $\tilde{v}_{\rm gr}\approx c$.
+Throughout this paper, we set $\tilde{v}_{\rm gr}=c$ in the fluid frame, which also implies $v_{\rm gr}=c$ in the lab frame.
+In particular, bulk acceleration and compression of the plasma exposed to the FRB is calculated assuming that the FRB propagates with speed $c$, so the radio wave packet is static in the coordinate $\xi=t-z/c$.
+This approximation requires a more demanding condition than $\tilde{\gamma}_{\rm gr}\gg 1$.
+Note that the plasma interacting with the radio wave packet forms a smooth compressive MHD wave of width $cT$.
+In MHD, compression tends to be redistributed across the fluid with an effective sound speed $\tilde{v}_{\rm s}$, which approaches $c$ in a magnetically dominated plasma.
+The deviation of FRB speed from $c$ may be neglected in the MHD calculations of plasma motion in the radio wave packet if
+$\tilde{\gamma}_{\rm gr}\gg\tilde{\gamma}_{\rm s}=(1-\tilde{v}_{\rm s}^{2}/c^{2})^{-1/2}.$ (A4)
+Then, the approximation $\tilde{v}_{\rm gr}=c$ can be used.
+The value of $\tilde{\gamma}_{\rm s}$ for a relativistic, hot, magnetized fluid is derived in Appendix A in Beloborodov (2024). It may be approximately described as
+$\tilde{\gamma}_{\rm s}\approx\sqrt{1+\sigma_{\rm eff}},\qquad\sigma_{\rm eff}=\frac{\sigma}{\tilde{\gamma}}=\frac{\tilde{B}^{2}}{4\pi\tilde{\gamma}mc^{2}\tilde{n}}$. (A5)
+This expression for $\tilde{\gamma}_{\rm s}$ becomes exact for a cold plasma ($\tilde{\gamma}=1$).
+For a hot plasma, Equation (A5) approximates the accurate $\tilde{\gamma}_{\rm s}$ within a numerical factor, which is close to unity (and depends on the plasma equation of state).
+The condition $\tilde{\gamma}_{\rm gr}\gg\tilde{\gamma}_{\rm s}$ is significantly more demanding than $\tilde{\gamma}_{\rm gr}\gg 1$ if $\sigma_{\rm eff}\gg 1$.
+Then, it can be stated as
+$\frac{\tilde{\gamma}_{\rm gr}}{\tilde{\gamma}_{\rm s}}\approx\frac{\tilde{\omega}}{\tilde{\omega}_{\rm L}}\gg 1,\qquad\tilde{\omega}_{\rm L}=\frac{e\tilde{B}}{mc\tilde{\gamma}}$. (A6)
+Appendix B Particle motion in the radio wave
+Consider a linearly polarized radio wave packet of the form introduced in Equations (4) and (5):
+$E_{\rm wave}=E_{0}(\xi)\sin(\omega\xi),\quad\;a_{\rm wave}=a_{0}(\xi)\cos(\omega\xi),$ (B1)
+where $\xi=t-z/c$.
+We choose $x$ and $y$ axes so that $\boldsymbol{E}_{\rm wave}$ oscillates along $x$, and $\boldsymbol{B}_{\rm wave}$ oscillates along $y$.
+The packet propagates along $z$ in a background field $B_{\rm bg}\ll E_{0}$ with $\boldsymbol{B}_{\rm bg}\parallel\boldsymbol{B}_{\rm wave}$.
+The background magnetic field lines can drift along $z$ with some speed $\beta_{\rm D}=E_{\rm bg}/B_{\rm bg}$.
+This speed describes the plasma bulk motion in the MHD picture.
+One could choose a reference frame so that the plasma ahead of the wave packet is at rest, $\beta_{\rm D}=0$.
+However, $\beta_{\rm D}$ can slowly vary with $\xi$ inside the packet (slowly compared to the wave oscillation and the particle Larmor rotation).
+Then, there is no global inertial frame that would satisfy $\beta_{\rm D}(\xi)=0$ across the packet.
+At each $\xi$, one can define the local fluid frame $\tilde{\cal K}$ where $\tilde{\beta}_{\rm D}=0$ and $\tilde{E}_{\rm bg}=0$.
+B.1 Regular oscillation regime
+In the regular oscillation regime, the plasma exposed to the radio wave moves as a coherent stream, i.e.
+the plasma particles have a common four-velocity $u^{\mu}$ varying as a function of the oscillation phase in the wave $\varphi\equiv\omega\xi$.
+The four-velocity oscillates with frequency $\omega$ around the mean value $\langle u^{\mu}\rangle$, where the local averaging (over one oscillation) is defined as
+$\langle...\rangle\equiv\frac{1}{2\pi}\oint...\,d\varphi,$ (B2)
+with the integral taken over $\varphi=\omega\xi$ changing by $2\pi$.
+This averaging operation is equivalent to spatial averaging over $z$ at a fixed time $t$, since $d\varphi=-\omega dz/c$ at $t=const$.
+In the fluid frame $\tilde{\cal K}$, $\langle\tilde{u}^{i}\rangle=0$ ($i=x,y,z$), so
+$\langle\tilde{u}^{\mu}\rangle=(\langle\tilde{\gamma}\rangle,0,0,0).$ (B3)
+Here, we assumed that the wave amplitude in the packet rises on a timescale $\tilde{\xi}_{\rm rise}\gg a_{0}\tilde{\omega}_{B}^{-1}$; then, gyration of $\langle\tilde{u}^{\mu}\rangle$ in $\tilde{B}_{\rm bg}$ is not excited, as shown in B22.
+In an arbitrary frame $\cal K$} moving along $z$, we have
+$\langle u^{\mu}\rangle=(\langle\gamma\rangle,0,0,\langle u_{z}\rangle).$ (B4)
+It is related to $\langle\tilde{u}^{\mu}\rangle$ by Lorentz transformation, which gives (using $\langle\tilde{u}_{z}\rangle=0$):
+$\langle\gamma\rangle=\gamma_{\rm D}\langle\tilde{\gamma}\rangle,\qquad\langle u_{z}\rangle=u_{\rm D}\langle\tilde{\gamma}\rangle$, (B5)
+where $u_{\rm D}=\gamma_{\rm D}\beta_{\rm D}$ and $\gamma_{\rm D}=(1-\beta_{\rm D}^{2})^{-1/2}$.
+Note that in this appendix $\gamma$ denotes the instantaneous value of the particle Lorentz factor in the wave; in the main text, $\gamma$ stands for $\langle\gamma\rangle$, as we omit $\langle$$\rangle$ to simplify the notation.
+The oscillation form of $u^{\mu}(\varphi)$ can be found from Equation (18) and the corresponding energy equation,
+$mc\,\frac{d\gamma}{dt}=e\boldsymbol{E}\cdot\boldsymbol{\beta}=e(E_{\rm bg}+E_{\rm wave})\beta_{x}.$ (B6)
+Since the problem is translationally invariant along $x$, the generalized momentum $mcu_{x}+eA_{x}/c$ is conserved, where $A_{x}=A_{\rm wave}+A_{\rm bg}$ and $B_{\rm bg}=dA_{\rm bg}/dz$.
+The small $B_{\rm bg}\ll E_{0}$ implies a small variation of $A_{\rm bg}$ over one wavelength, so $u_{x}+a_{\rm wave}\approx const=\langle u_{x}\rangle=0$ and
+$u_{x}(\varphi)=-a_{0}\cos\varphi.$ (B7)
+Conservation of generalized momentum $mcu_{y}+eA_{y}/c$ (implied by $\partial_{y}=0$) gives $u_{y}(\varphi)=0$ since $A_{y}=0$.
+The remaining component $u_{z}(\varphi)$ and the Lorentz factor $\gamma(\varphi)$ can be found by considering $u_{\xi}\equiv d\xi/d\tau=\gamma-u_{z}$, where $\tau$ is the proper time along the particle worldline.
+Subtracting the $z$-component of Equation (18) from Equation (B6), one finds
+$\frac{du_{\xi}}{dt}=\frac{e(E\beta_{x}-B\beta_{x})}{mc}=\frac{e(E_{\rm bg}-B_{\rm bg})\beta_{x}}{mc}.$ (B8)
+Since $E_{\rm bg}\ll E_{0}$ and $B_{\rm bg}\ll E_{0}$, one can see that $u_{\xi}$ weakly varies during one wave oscillation, $u_{\xi}\approx const$.
+This fact implies that the proper density of the plasma stream $\hat{\rho}\approx const$ during each oscillation.
+Indeed, plasma density $\rho(\varphi)$ satisfies the continuity equation
+$\partial_{t}\rho+\partial_{z}(\rho v_{z})=0\quad\Rightarrow\quad\frac{d}{d\xi}\left[\rho(1-\beta_{z})\right]=0.$ (B9)
+Hence, the proper density of the oscillating stream $\hat{\rho}=\rho/\gamma$ satisfies
+$\hat{\rho}u_{\xi}=const,$ (B10)
+and so $\hat{\rho}\approx const$ during one oscillation. The weak variation of $u_{\xi}$ also implies the relation
+$u_{\xi}\approx\langle u_{\xi}\rangle=\langle\gamma-u_{z}\rangle=\frac{\langle\tilde{\gamma}\rangle}{\kappa},\qquad\kappa\equiv\gamma_{\rm D}+u_{\rm D},$ (B11)
+where we used Equation (B5).
+In particular, in the fluid rest frame, $\tilde{\kappa}=1$ and $\tilde{u}_{\xi}=\langle\tilde{\gamma}\rangle$.
+The value of $\langle\tilde{\gamma}\rangle$ can be determined using the identity $\gamma^{2}=1+u_{x}^{2}+u_{z}^{2}$, which gives
+$(u_{z}+u_{\xi})^{2}=1+u_{x}^{2}+u_{z}^{2}\;\;\Rightarrow\;\,2u_{\xi}\langle u_{z}\rangle+u_{\xi}^{2}=1+a^{2},$ (B12)
+where we used $\langle u_{x}^{2}\rangle=\langle a_{\rm wave}^{2}\rangle=a^{2}$.
+In the fluid frame $\tilde{\cal K}$, $\langle\tilde{u}_{z}\rangle=0$ and $\tilde{u}_{\xi}=\langle\tilde{\gamma}\rangle$, so the above identity gives
+$\langle\tilde{\gamma}\rangle^{2}=1+a^{2}.$ (B13)
+Explicit expressions for $\gamma(\varphi)$ and $u_{z}(\varphi)$ can be obtained as follows. Using $dt=(\gamma/u_{\xi})d\xi$, one finds from Equation (B6)
+$\frac{d\gamma}{d\xi}=\frac{eE_{\rm wave}u_{x}}{mcu_{\xi}}\;\;\Rightarrow\;\gamma=\frac{a_{0}^{2}\cos 2\varphi}{4u_{\xi}}+\langle\gamma\rangle,$ (B14)
+where $E_{\rm wave}u_{x}=-E_{0}a_{0}\sin\varphi\cos\varphi$ and we used $u_{\xi}\approx const$ during one oscillation. Equation (B14) also determines $u_{z}=\gamma-u_{\xi}$. Thus, one finds
+$\gamma(\varphi)$ $=\frac{\kappa a_{0}^{2}\cos 2\varphi}{4\langle\tilde{\gamma}\rangle}+\gamma_{\rm D}\langle\tilde{\gamma}\rangle$, (B15)
+$u_{z}(\varphi)$ $=\frac{\kappa a_{0}^{2}\cos 2\varphi}{4\langle\tilde{\gamma}\rangle}+u_{\rm D}\langle\tilde{\gamma}\rangle$. (B16)
+These expressions hold in any frame $\cal K$} moving along $z$.
+In particular, in the fluid frame, $\tilde{u}_{\rm D}=0$, $\tilde{\gamma}_{\rm D}=1$, and $\tilde{\kappa}=1$.
+In this frame, each particle executes a closed periodic orbit in the oscillating wave.
+Indeed, the net displacement in $\tilde{z}$ during one oscillation period is
+$\Delta\tilde{z}=\oint\tilde{v}_{z}d\tilde{t}=\oint\frac{\tilde{u}_{z}\,d\varphi}{\tilde{\omega}\tilde{u}_{\xi}}=\frac{2\pi\langle\tilde{u}_{z}\rangle}{\tilde{\omega}\tilde{u}_{\xi}}=0.$ (B17)
+Note that $\tilde{B}_{\rm bg}$ does not enter the solution for $\tilde{u}^{\mu}(\varphi)$; it is the same as at $\tilde{B}_{\rm bg}=0$.
+This solution is well known: the particle executes an “8”-shaped orbit (Landau & Lifshitz, 1975).
+The above derivation shows that the particles execute the same motion in a slowly varying $B_{\rm bg}$, as long as their motion is viewed in the local fluid frame $\tilde{\cal K}$.
+The role of $B_{\rm bg}$ and $E_{\rm bg}$ is that they control the fluid speed $\beta_{\rm D}$ relative to a fixed lab frame: $\beta_{\rm D}=E_{\rm bg}/B_{\rm bg}$.
+A slow (adiabatic) variation of $\beta_{\rm D}$ changes the frame $\tilde{\cal K}$ while the shape of the particle orbit viewed in $\tilde{\cal K}$ remains unchanged.
+The size of the orbit and the average Lorentz factor $\langle\tilde{\gamma}\rangle$ are determined by the local wave strength $a=\langle a_{\rm wave}^{2}\rangle^{1/2}$ (Equation B13).
+B.2 Stochastic heating regime
+In the regime of stochastic heating, the plasma particles gain very high energies. The particle motion viewed in the fluid frame $\tilde{\cal K}$ has two important features (B22):
+(1) The particle gyrates with a Lorentz factor $\tilde{\gamma}\gg a$ and Larmor frequency $\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}\ll\tilde{\omega}$.
+The gyration occurs in the magnetic field $B$ averaged over the fast wave oscillation: $\langle B_{\rm bg}+B_{\rm wave}\rangle=B_{\rm bg}$.
+The particle oscillation in the radio wave creates a small modulation $|\Delta u|\sim a$ superimposed on the circular gyration.
+(2) The particle experiences a resonance with the radio wave every Larmor rotation.
+It occurs at the gyro-phase $\psi$ where the particle velocity aligns with the wave propagation direction (i.e.
+with the $z$-axis); we define this gyro-phase as $\psi=0$.
+The resonance occurs in a narrow interval $\delta\psi\ll 1$ and suddenly changes $\tilde{\gamma}$ by a random $\Delta\tilde{\gamma}$ given in Equation (B21) below.
+Between the repeating resonances, the gyrating particle moves with $\tilde{\gamma}\approx const$, i.e.
+its trajectory in the fluid frame is a circle in the $\tilde{x}\tilde{z}$ plane.
+Its four-velocity satisfies $\tilde{u}_{x}^{2}+\tilde{u}_{z}^{2}=\tilde{\gamma}^{2}-1=const$ with
+$\tilde{u}_{x}=\sqrt{\tilde{\gamma}^{2}-1}\sin\psi,\qquad\tilde{u}_{z}=\sqrt{\tilde{\gamma}^{2}-1}\cos\psi.$ (B18)
+Quantity $\tilde{u}_{\xi}(\psi)=\tilde{\gamma}-\tilde{u}_{z}$ varies between $(2\tilde{\gamma})^{-1}$ and $2\tilde{\gamma}$; its value averaged over one full Larmor orbit between two consecutive resonances equals $\tilde{\gamma}$.
+In a different frame $\cal K$} moving along $z$, the fluid frame moves with speed $\beta_{\rm D}$, and the particle four-velocity is determined by Lorentz transformation:
+$u_{x}=\tilde{u}_{x},\qquad u_{z}=u_{\rm D}\tilde{\gamma}+\gamma_{\rm D}\sqrt{\tilde{\gamma}^{2}-1}\cos\psi.$ (B19)
+The stochastic pumping of $\tilde{\gamma}$ by the repeating resonance events implies a quickly developing chaos, destroying the coherent motion of the plasma stream in the wave (B22).
+As a result, at a given location $\xi$ in the radio wave packet, individual particles have random gyration phases $\psi$ and a broad distribution of $\tilde{\gamma}$.
+The ensemble average $\langle\tilde{\gamma}\rangle_{\rm ens}$ grows with time, which means heating of the plasma.
+It occurs when [^15] B22 defined the critical $\tilde{\gamma}_{\star}$ slightly differently, using $a_{0}$ instead of $a=a_{0}/\sqrt{2}$.
+All quantities in B22 were measured in frame $\tilde{\cal K}$ (where the background field is static) and denoted without tilde.
+$\tilde{\gamma}\gtrsim\tilde{\gamma}_{\star}=\left(\frac{a^{3}}{b}\right)^{1/2}\qquad(a\gg 1),$ (B20)
+where $b=\tilde{\omega}_{B}/\tilde{\omega}$.
+This condition implies that regular oscillations in the wave with $\tilde{\gamma}\sim a\gg 1$ switch to the stochastic heating regime if $b\gtrsim a$ (the more accurate condition is $b>a/3$, see Figure 3).
+The stochastic heating of gyrating particles consumes energy of the radio wave, i.e.
+it is an absorption process.
+This process is controlled by particle energy exchange with the radio wave at each resonance, $\Delta\tilde{\gamma}$. It may be written in the form
+$\Delta\tilde{\gamma}=H\left(\frac{a_{0}^{3}\tilde{\gamma}}{b}\right)^{1/3}\cos\varphi+\langle\Delta\tilde{\gamma}\rangle$, (B21)
+where $H\approx 2.6$ is a numerical factor, $\varphi$ is the (random) phase of the radio wave at the particle location when it has the gyro-phase $\psi=0$.
+The term $\propto\cos\varphi$ describes random kicks to $\tilde{\gamma}$; it is derived in B22.
+The term $\langle\Delta\tilde{\gamma}\rangle$ is a small systematic gain (seen in Figure 2 in B22 as a small shift of the measured energy exchange events above the $\cos\varphi$ curve).
+It was derived for particles with $\tilde{\gamma}\gg a$ by Lyubarsky (2018):
+$\langle\Delta\tilde{\gamma}\rangle=\frac{2^{7/3}\pi^{2}a_{0}^{2}}{3^{5/3}[\Gamma(1/3)]^{2}\,b^{2/3}\tilde{\gamma}^{1/3}}\approx 1.1\,\frac{a_{0}^{2}}{b^{2/3}\tilde{\gamma}^{1/3}},$ (B22)
+where $\Gamma(1/3)\approx 2.68$.
+This result may be interpreted as usual synchrotron absorption of a weak monochromatic wave by particles with effective rest mass $m_{\rm eff}=m\sqrt{1+a^{2}}$ gyrating with Lorentz factor $\tilde{\gamma}_{\rm eff}=\tilde{\gamma}/\sqrt{1+a^{2}}$ (Lyubarsky, 2018).
+One can see that $\langle\Delta\tilde{\gamma}\rangle$ is smaller than the typical random part of $\Delta\tilde{\gamma}$ by the factor of $\sim(\tilde{\gamma}_{\star}/\tilde{\gamma})^{2/3}$.
+[…]
+In section III.2, we calculate the heating rate $d\langle\tilde{\gamma}\rangle_{\rm ens}/dt$ using the method of B22: we follow the motion of a large ensemble of (initially static) particles that become exposed to the incoming radio wave, and find the evolution of $\langle\tilde{\gamma}\rangle_{\rm ens}$.
+The result is described by Equation (15) with $\chi\approx 0.8$.
+Appendix C Plasma stress-energy tensor
+C.1 Regular oscillation regime
+The coherently oscillating plasma stream has the proper density $\hat{\rho}=\rho/\gamma$ nearly constant during each oscillation (Appendix B.1).
+Therefore, the plasma density averaged over one oscillation may be expressed as
+$\langle\rho\rangle=\langle\gamma\rangle\,\hat{\rho}$, (C1)
+where the averaging operation $\langle$$\rangle$ is defined in Equation (B2).
+The fluid rest frame $\tilde{\cal K}$ satisfies $\langle\tilde{u}_{z}\rangle=0$ and $\langle\tilde{\gamma}\rangle=(1+a^{2})^{1/2}$ (Equation B13), so
+$\langle\tilde{\rho}\rangle=\sqrt{1+a^{2}}\,\hat{\rho}$. (C2)
+Lorentz transformation of four-velocity gives $\gamma=\gamma_{\rm D}\tilde{\gamma}+u_{\rm D}\tilde{u}_{z}$, which implies $\langle\gamma\rangle=\gamma_{\rm D}\langle\tilde{\gamma}\rangle$ and
+$\langle\rho\rangle=\gamma_{\rm D}\langle\tilde{\rho}\rangle$. (C3)
+The stress-energy tensor of the coherent oscillating stream is
+$T_{\rm stream}^{\mu\nu}(\varphi)=\hat{\rho}\,u^{\mu}u^{\nu}.$ (C4)
+In the plane-parallel problem, the relevant components of $T^{\mu\nu}$ are $T^{tt}$, $T^{tz}$, and $T^{zz}$.
+They are determined by $u^{t}=\gamma$ and $u^{z}$ (Equations (B15) and (B16)).
+The plasma stress-energy tensor averaged over one oscillation is
+$T_{\rm p}^{\mu\nu}=\langle T_{\rm stream}^{\mu\nu}\rangle=\frac{\hat{\rho}c^{2}}{2\pi}\oint u^{\mu}u^{\nu}\,d\varphi.$ (C5)
+The averaging operation is Lorentz invariant among all frames moving along $z$, since $\varphi=inv$. As a result, Equation (C5) yields a well-defined tensor
+$T_{\rm p}^{\mu\nu}=\frac{c^{2}\langle\tilde{\rho}\rangle\kappa^{2}a^{4}}{8(1+a^{2})^{3/2}}+c^{2}\langle\tilde{\rho}\rangle(1+a^{2})^{1/2}u_{\rm D}^{\mu}u_{\rm D}^{\nu},$ (C6)
+where $u_{\rm D}^{\mu}=(\gamma_{\rm D},u_{\rm D})$.
+It is straightforward to verify that the components $T_{\rm p}^{\mu\nu}$ with $\mu,\nu\in\{t,z\}$ satisfy the tensor transformation law under Lorentz boosts along zz.
+[^16] In the fluid frame, the components have the form $T_{\rm p}^{\tilde{\mu}\tilde{\nu}}=K+G\delta^{\tilde{\mu}}_{0}\delta^{\tilde{\nu}}_{0}$ where $K$ and $G$ are scalars.
+Tensor transformation with Lorentz matrix $\Lambda^{\mu}_{\tilde{\mu}}$ then gives $T_{\rm p}^{\mu\nu}=\Lambda^{\mu}_{\tilde{\mu}}\Lambda^{\nu}_{\tilde{\nu}}T_{\rm p}^{\tilde{\mu}\tilde{\nu}}=\kappa^{2}K+Gu_{\rm D}^{\mu}u_{\rm D}^{\nu}$, consistent with Equation (C6).
+In the main text, we drop $\langle$$\rangle$ to simplify notation:
+$\langle\tilde{\gamma}\rangle\rightarrow\tilde{\gamma},\qquad\langle\tilde{\rho}\rangle\rightarrow\tilde{\rho}$.
+C.2 Stochastic heating regime
+In the stochastic heating regime, the particle motion is dominated by gyration about $\boldsymbol{B}_{\rm bg}$ with $\tilde{\gamma}\gg a$.
+Fast oscillations in the wave occur with a smaller four-velocity $|\Delta\boldsymbol{u}|\sim a$ and are superimposed on the circular Larmor orbit.
+In this regime, the gyrating plasma may be approximated as a two-dimensional relativistic gas.
+Its internal energy density $U_{\rm p}$ and pressure $P_{\rm p}$ (measured in frame $\tilde{\cal K}$) are given by
+$U_{\rm p}=2P_{\rm p}=(\tilde{\gamma}-1)\tilde{\rho}c^{2}.$ (C7)
+The plasma stress-energy tensor is
+$T^{\mu\nu}_{\rm p}=(\tilde{\rho}c^{2}+U_{\rm p}+P_{\rm p})u_{\rm D}^{\mu}u_{\rm D}^{\nu}+\eta^{\mu\nu}P_{\rm p},$ (C8)
+where $\eta^{\mu\nu}$ is the Minkowski metric.
+Appendix D Radiative losses
+A relativistic electron (or positron) with an instantaneous four-velocity $u^{\mu}=(\gamma,\boldsymbol{u})$ exposed to fields $\boldsymbol{E}$ and $\boldsymbol{B}$ radiates energy with rate (Landau & Lifshitz, 1975)
+$\dot{{\cal E}}_{e}=\frac{c\sigma_{\rm T}}{4\pi}\left[(\gamma\boldsymbol{E}+\boldsymbol{u}\times\boldsymbol{B})^{2}-(\boldsymbol{u}\cdot\boldsymbol{E})^{2}\right],$ (D1)
+where $\sigma_{\rm T}$ is the Thomson cross section.
+$\dot{{\cal E}}_{e}$ is a Lorentz-invariant quantity.
+The particle also radiates momentum with rate $\dot{\boldsymbol{\cal P}}_{e}=\boldsymbol{\beta}\dot{{\cal E}}_{e}/c$.
+In our case, $\boldsymbol{E}=\boldsymbol{E}_{\rm bg}+\boldsymbol{E}_{\rm wave}$ is along $x$, $\boldsymbol{B}=\boldsymbol{B}_{\rm bg}+\boldsymbol{B}_{\rm wave}$ is along $y$, and $u_{y}=0$.
+Then, Equation (D1) gives
+$\dot{{\cal E}}_{e}=\frac{c\sigma_{\rm T}}{4\pi}\left\{\left[u_{\xi}E+\gamma(B_{\rm bg}-E_{\rm bg})\right]^{2}-B^{2}+E^{2}\right\}.$ (D2)
+In both regimes, with regular oscillations or stochastic heating, emission in the fluid frame $\tilde{\cal K}$ occurs with $\tilde{u}_{\xi}\sim\tilde{\gamma}\gg 1$.
+Then, for radio waves with amplitudes $\tilde{E}_{0}\gg\tilde{B}_{\rm bg}$, Equation (D2) simplifies to
+$\dot{{\cal E}}_{e}=\frac{c\sigma_{\rm T}}{4\pi}u_{\xi}^{2}E_{\rm wave}^{2}\left[1+{\cal O}\left(\frac{\tilde{B}_{\rm bg}}{\tilde{E}_{\rm wave}}\right)\right].$ (D3)
+Let $\delta t$ be the time it takes the particle to move one oscillation period $\delta\xi=2\pi/\omega$ through the wave:
+$\delta t=\int dt=\oint\frac{\gamma\,d\varphi}{u_{\xi}\omega}\approx\frac{2\pi}{\omega u_{\xi}}\langle\gamma\rangle,$ (D4)
+where we used the fact that quantity $u_{\xi}=\gamma-u_{z}$ weakly varies during one oscillation of the wave: $u_{\xi}\approx const$ (Appendix B.1).
+The energy and momentum emitted by the particle during time $\delta t$ are
+$\delta{\cal E}_{e}=\!\frac{c\sigma_{\rm T}}{4\pi}\!\int E_{\rm wave}^{2}\,u_{\xi}^{2}\,dt=\frac{c\sigma_{\rm T}u_{\xi}}{4\pi\omega}\oint E_{\rm wave}^{2}\gamma\,d\varphi,$ (D5)
+$\delta\boldsymbol{{\cal P}}_{e}=\frac{\sigma_{\rm T}}{4\pi}\int E_{\rm wave}^{2}u_{\xi}^{2}\boldsymbol{\beta}dt=\frac{\sigma_{\rm T}u_{\xi}}{4\pi\omega}\!\oint E_{\rm wave}^{2}\boldsymbol{u}\,d\varphi.$ (D6)
+Thus, the emitted four-momentum $\delta{\cal P}_{e}^{\mu}=(\delta{\cal E}_{e}/c,\delta\boldsymbol{{\cal P}}_{e})$ is
+$\delta{\cal P}^{\mu}_{e}=\frac{2\sigma_{\rm T}Uu_{\xi}}{\omega}\oint u^{\mu}\sin^{2}\!\varphi\,d\varphi,$ (D7)
+where $U=E_{0}^{2}/8\pi$ is the energy density of the radio wave. The average emission rate during $\delta t$ is
+$\frac{\delta{\cal P}^{\mu}_{e}}{\delta t}=\frac{2\sigma_{\rm T}Uu_{\xi}^{2}}{\langle\gamma\rangle}\langle u^{\mu}\sin^{2}\!\varphi\rangle.$ (D8)
+D.1 Regular oscillation regime
+In the regular oscillation regime, the components of $u^{\mu}(\varphi)$ are given by Equations (B7), (B15), and (B16), with $\langle\gamma\rangle=\gamma_{\rm D}\sqrt{1+a^{2}}$ and $u_{\xi}=\kappa^{-1}\sqrt{1+a^{2}}$.
+Then, Equation (D8) yields two non-zero components of $\delta{\cal P}^{\mu}_{e}/\delta t$:
+$\frac{\delta{\cal E}_{e}}{\delta t}$ $=\frac{c\sigma_{\rm T}U}{\kappa^{2}}\left[1+a^{2}-\frac{\kappa a^{2}}{4\gamma_{\rm D}}\right],$ (D9)
+$\frac{\delta{\cal P}^{z}_{e}}{\delta t}$ $=\frac{\sigma_{\rm T}U}{\kappa^{2}}\left[\beta_{\rm D}(1+a^{2})-\frac{\kappa a^{2}}{4\gamma_{\rm D}}\right].$ (D10)
+These expressions hold in any frame moving along $z$.
+While the instantaneous losses $\dot{{\cal E}}_{e}$ are Lorentz-invariant, their average value is not — $\delta{\cal E}_{e}/\delta t$ depends on the frame (because of the term $\kappa a^{2}/4\gamma_{\rm D}$).
+Note that in the fluid frame $\tilde{\cal K}$ the particle radiates $z$-momentum with rate $\delta\tilde{\cal P}^{z}_{e}/\delta\tilde{t}=-\sigma_{\rm T}\tilde{U}a^{2}/4\neq 0$, even though its orbit has the symmetric “8” shape.
+D.2 Stochastic heating regime
+In the regime of stochastic heating, the particle executes gyration in $B_{\rm bg}$ with four-velocity $u^{\mu}$ weakly modulated by the fast oscillations in the wave.
+Therefore, in Equation (D7) one can take $u^{\mu}\approx const$, and the emitted four-momentum averaged over one wave oscillation becomes
+$\frac{\delta{\cal P}^{\mu}_{e}}{\delta t}=\sigma_{\rm T}U\,\frac{u_{\xi}^{2}u^{\mu}}{\gamma}.$ (D11)
+In the MHD description of plasma dynamics, the radiative losses averaged over the fast oscillation in the wave should be further averaged over Larmor rotation.
+This is convenient to do in the fluid frame $\tilde{\cal K}$ because in this frame the particle motion between the successive resonances has $\tilde{\gamma}\approx const$ (section B.2), so its orbit is a Larmor circle with the orbital time
+$\Delta\tilde{t}=\frac{2\pi\tilde{\gamma}}{\tilde{\omega}_{B}}\gg\delta\tilde{t}$. (D12)
+The gyro-phase $\psi$ is defined as the angle between the particle velocity $\tilde{\boldsymbol{\beta}}$ and the $z$-axis, so
+$\tilde{\beta}_{z}=\tilde{\beta}\cos\psi\approx\cos\psi.$ (D13)
+Integration over $\Delta\tilde{t}$ can be reduced to the integration over $0<\psi<2\pi$ using $d\psi/d\tilde{t}=\tilde{\omega}_{\rm L}=\tilde{\omega}_{B}/\tilde{\gamma}$.
+Then, the energy and $z$-momentum emitted by the particle during $\Delta\tilde{t}$ can be expressed as
+$\Delta\tilde{\cal E}_{e}=\frac{c\sigma_{\rm T}\tilde{U}}{\tilde{\omega}_{\rm L}}\!\int_{0}^{2\pi}\!\!\tilde{u}_{\xi}^{2}\,d\psi,$ (D14)
+$\Delta\tilde{\cal P}_{e}=\frac{\sigma_{\rm T}\tilde{U}}{\tilde{\omega}_{\rm L}}\!\int_{0}^{2\pi}\!\!\tilde{u}_{\xi}^{2}\tilde{\beta}_{z}\,d\psi.$ (D15)
+Substituting $\tilde{u}_{\xi}=\tilde{\gamma}(1-\tilde{\beta}_{z})$ and $\tilde{\beta}_{z}\approx\cos\psi$, we find
+$\frac{\Delta\tilde{\cal E}_{e}}{\Delta\tilde{t}}=\frac{3}{2}c\sigma_{\rm T}\tilde{U}\tilde{\gamma}^{2},\qquad\frac{\Delta\tilde{\cal P}_{e}}{\Delta\tilde{t}}=-\sigma_{\rm T}\tilde{U}\tilde{\gamma}^{2}.$ (D16)
+Although the particle orbit is nearly circular, it radiates a large net momentum along z-z.
+This occurs because the instantaneous particle acceleration is dominated by the fast-oscillating electromagnetic field of the wave (even though the modulation effect of the wave on the particle velocity is small, so that the particle continues to move on the circular Larmor orbit with $\tilde{\gamma}\approx const$).
+The fact that the wave field propagates along $\tilde{z}$ breaks the emission symmetry $\tilde{z}\leftrightarrow\tilde{z}$, as seen in Equation (D15): $\tilde{u}_{\xi}^{2}=(\tilde{\gamma}-\tilde{u}_{z})^{2}$ decreases at $\tilde{u}_{z}>0$ and increases at $\tilde{u}_{z}<0$.
+The average loss rate in another frame $\cal K$} moving along $z$ is determined by the Lorentz transformation,
+$\Delta{\cal E}_{e}$ $=\gamma_{\rm D}\Delta\tilde{\cal E}_{e}+u_{\rm D}c\Delta\tilde{\cal P}_{e},$ (D17)
+$c\Delta{\cal P}_{e}$ $=u_{\rm D}\Delta\tilde{\cal E}_{e}+\gamma_{\rm D}c\Delta\tilde{\cal P}_{e},$ (D18)
+and $\Delta t=\gamma_{\rm D}\Delta\tilde{t}$. Thus, we find
+$\frac{\Delta{\cal E}_{e}}{\Delta t}$ $=\frac{c\sigma_{\rm T}U}{\kappa^{2}}\left(\frac{3}{2}-\beta_{\rm D}\right)\tilde{\gamma}^{2},$ (D19)
+$\frac{\Delta{\cal P}_{e}}{\Delta t}$ $=\frac{\sigma_{\rm T}U}{\kappa^{2}}\left(\frac{3}{2}\beta_{\rm D}-1\right)\tilde{\gamma}^{2},$ (D20)
+where we used $\tilde{U}=U/\kappa^{2}$, which expresses transformation of the wave energy density from lab frame $\cal K$} to the fluid frame $\tilde{\cal K}$ (Equation 11).